### v1

# Regularized KL-Divergence for Well-Defined Function-Space Variational Inference in Bayesian neural networks

Abstract
Bayesian neural networks (BNN) promise to combine the predictive performance of neural networks with principled uncertainty modeling important for safety-critical systems and decision making. However, posterior uncertainty estimates depend on the choice of prior, and finding informative priors in weight-space has proven difficult. This has motivated variational inference (VI) methods that pose priors directly on the function generated by the BNN rather than on weights. In this paper, we address a fundamental issue with such function-space VI approaches pointed out by Burt et al., (2020), who showed that the objective function (ELBO) is negative infinite for most priors of interest. Our solution builds on generalized VI (Knoblauch et al.,, 2019) with the regularized KL divergence (Quang,, 2019) and is, to the best of our knowledge, the first well-defined variational objective for function-space inference in BNNs with Gaussian process (GP) priors. Experiments show that our method incorporates the properties specified by the GP prior on synthetic and small real-world data sets, and provides competitive uncertainty estimates for regression, classification and out-of-distribution detection compared to BNN baselines with both function and weight-space priors.
1 Introduction
Neural networks have shown impressive results in many fields but fail to provide well-calibrated uncertainty estimates, which are essential in applications associated with risk, such as healthcare (Kompa et al.,, 2021; Abdullah et al.,, 2022) or finance (Bew et al.,, 2019; Wong,, 2023). Bayesian neural networks (BNNs) offer to combine the scalability and predictive performance of neural networks with principled uncertainty modeling by explicitly capturing epistemic uncertainty, i.e., uncertainty resulting from finite training data. While the choice of prior in the Bayesian framework strongly affects the uncertainty estimates later obtained from the posterior, specifying informative priors on BNN weights has proven difficult and is hypothesized to have limited their practical applicability (Knoblauch et al.,, 2019; Cinquin et al.,, 2021; Tran et al.,, 2022). For instance, the default isotropic Gaussian prior, which is often chosen for tractability rather than for the beliefs it carries (Knoblauch et al.,, 2019), is known to have pathological behavior in some cases (Tran et al.,, 2022). A promising approach to solve this issue is to place priors directly on the function represented by the BNN instead of the weights. While being technically more challenging, function-space priors allow incorporating interpretable knowledge into the inference, for instance enabling the use of the extensive Gaussian Process (GP) literature to improve prior selection and design (Williams and Rasmussen,, 2006).
A recent line of work has focused on using function-space priors in BNNs with variational inference (VI) (Sun et al.,, 2019). VI is an appealing method because of its successful application to BNNs, its flexibility in terms of approximate posterior parameterization, and its scalability to large datasets and models (Hoffman et al.,, 2013; Blundell et al.,, 2015; Tomczak et al.,, 2020). Unfortunately, in the context of BNNs with function-space priors, the Kullbach-Leibler (KL) divergence term in the VI objective (ELBO) involves two intractabilities: (i) a supremum over infinitely many subsets and (ii) access to the density of the distribution of function represented by the BNN, which has no closed-form expression. Sun et al., (2019) propose to address problem (i) by approximating the supremum in the KL divergence by an expectation, and problem (ii) by using implicit score function estimators (which make this method difficult to use in practice (Ma and Hernández-Lobato,, 2021)). However, the problem is actually more severe. Not only is the KL divergence intractable, it is infinite in many cases of interest (Burt et al.,, 2020), such as when the prior is a non-degenerate GP or a BNN with a different architecture. Thus, in these (and many more) situations, the KL divergence cannot even be approximated. As a consequence, more recent work abandons using BNNs and instead uses deterministic neural networks to parameterize basis functions (Ma and Hernández-Lobato,, 2021) or a GP mean (Wild et al.,, 2022). The only prior work (Rudner et al., 2022b, ) on function-space VI in BNNs that overcomes the issue pointed out by Burt et al., (2020) does so by deliberately limiting itself to cases where the KL divergence is known to be finite (by defining the prior as the pushforward of a weight-space distribution). Therefore, the method by Rudner et al., 2022b suffers from the same issues regarding prior specification as any other weight-space inference method for BNNs.
In this paper, we address the argument by Burt et al., (2020) that VI does not provide a valid objective for inference in BNNs with genuine function-space priors, and we propose to apply the framework of generalized variational inference (Knoblauch et al.,, 2019). We present a simple method for function-space inference with GP priors that builds on the so-called regularized KL divergence (Quang,, 2019), which generalizes the conventional KL divergence and is finite for any pair of Gaussian measures. We obtain a Gaussian measure for the variational posterior by considering the linearized BNN from Rudner et al., 2022b , and we are free to choose a prior from a large set of GPs which have an associated Gaussian measure on the considered function space. While the regularized KL divergence is still intractable, it can be consistently estimated from samples with a known error bound. We find that our method effectively incorporates the beliefs specified by GP priors into the inference process (see figure 1) and provides competitive performance compared to BNN baselines. To the best of our knowledge, our method is the first to provide a well-defined objective for function-space inference in BNNs with informative GP priors. Our contributions are summarized below:
  1. 1.
    We propose to use generalized VI with the so-called regularized KL divergence to mitigate the issue of an infinite KL divergence when using VI in BNNs with function-space priors.
  2. 2.
    We present a new, well-defined objective for function-space inference in the linearized BNN with GP priors, which results in a simple algorithm.
  3. 3.
    We analyze our method empirically on synthetic and real-world data sets, and find that it accurately captures structural properties specified by the GP prior and provides competitive uncertainty estimates for regression, classification, and out-of-distribution detection compared to BNN baselines with both function-space and weight-space priors.
The paper is structured as follows. section 2 introduces function-space VI in BNNs and the regularized KL divergence. section 3 presents our proposed method for generalized function-space VI (GFSVI) in BNNs. section 4 presents experimental results. We discuss related work in section 5 and conclude in section 6.
Figure 1: BNN inference on synthetic data (gray circles) with Gaussian process priors encoding different properties such as smoothness (increasing from Matern-1/2 to RBF) and periodicity (periodic).
2 Background
In this section, we provide background on function-space variational inference (VI) in BNNs and discuss the fundamental issue of an infinite KL-divergence. We then introduce the regularized KL divergence, which is the basis for our solution presented in Section 3.
2.1 Function-space VI in BNNs
We consider a neural network $f({\,\cdot\,;{\bm{w}}})$ with weights ${\bm{w}}\in^{p}$, and a data set $\mathcal{D}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}$ with features ${\bm{x}}_{i}\in\mathcal{X}\subset^{d}$ and values $y_{i}\in\mathcal{Y}$. Bayesian Neural Networks (BNNs) are specified further by a likelihood function $p\mathopen{}\left(\mathcal{D}\,|\,{\bm{w}}\right)\mathclose{}=\prod_{i=1}^{N}p\mathopen{}\left(y_{i}\,|\,f({\bm{x}}_{i};{\bm{w}})\right)\mathclose{}$ and—traditionally—a prior $p\mathopen{}\left({\bm{w}}\right)\mathclose{}$ on the weights, and one seeks the posterior distribution $p\mathopen{}\left({\bm{w}}\,|\,\mathcal{D}\right)\mathclose{}\propto p\mathopen{}\left(\mathcal{D}\,|\,{\bm{w}}\right)\mathclose{}\,p\mathopen{}\left({\bm{w}}\right)\mathclose{}$. The method proposed in this paper builds on variational inference (VI), which approximates $p\mathopen{}\left({\bm{w}}\,|\,\mathcal{D}\right)\mathclose{}$ with a variational distribution $q_{\phi}\mathopen{}\left({\bm{w}}\right)\mathclose{}$, whose variational parameters $\phi$ maximize the evidence lower bound (ELBO),
$\mathcal{L}(\phi):={\mathop{{}\mathbb{E}}}_{q_{\phi}\mathopen{}\left({\bm{w}}\right)\mathclose{}}\mathopen{}\left[\log p\mathopen{}\left(\mathcal{D}\,|\,{\bm{w}}\right)\mathclose{}\right]\mathclose{}-D_{\mathrm{KL}}\mathopen{}\left(q_{\phi}\,\|\,p\right)\mathclose{}$ (2.1)
where $D_{\text{KL}}$ is the Kullback-Leibler (KL) divergence
$D_{\mathrm{KL}}\mathopen{}\left(q_{\phi}\,\|\,p\right)\mathclose{}:={\mathop{{}\mathbb{E}}}_{q_{\phi}\mathopen{}\left({\bm{w}}\right)\mathclose{}}\mathopen{}\left[\log\mathopen{}\left(q_{\phi}\mathopen{}\left({\bm{w}}\right)\mathclose{}\big{/}p\mathopen{}\left({\bm{w}}\right)\mathclose{}\right)\mathclose{}\right]\mathclose{}.$ (2.2)
At test time, we approximate the predictive distribution for given features ${\bm{x}}^{*}$ as $p\mathopen{}\left(y^{*}\,|\,{\bm{x}}^{*}\right)\mathclose{}={\mathop{{}\mathbb{E}}}_{p\mathopen{}\left({\bm{w}}\,|\,\mathcal{D}\right)\mathclose{}}\mathopen{}\left[\big{.}p\mathopen{}\left(y^{*}\,|\,f({\bm{x}}^{*};{\bm{w}})\right)\mathclose{}\right]\mathclose{}\approx{\mathop{{}\mathbb{E}}}_{q_{\phi}({\bm{w}})}\mathopen{}\left[\big{.}p\mathopen{}\left(y^{*}\,|\,f({\bm{x}}^{*};{\bm{w}})\right)\mathclose{}\right]\mathclose{}$.
Function-space variational inference
Since weights of neural networks are not interpretable, we abandon the weight-space prior $p({\bm{w}})$ and instead pose a prior $\mathbb{P}$ directly on the function $f\mathopen{}\left(\,\cdot\,;{\bm{w}}\right)\mathclose{}$, which we denote simply as $f$ when there is no ambiguity. Here, the symbol $\mathbb{P}$ denotes a probability measure that does not admit a density since the function space is infinite-dimensional. We compute the expected log-likelihood as in the first term of Eq. 2.1, but we calculate the KL-term (Eq. 2.2) using the pushforward of $q_{\phi}({\bm{w}})$ along the mapping ${\bm{w}}\mapsto f\mathopen{}\left(\,\cdot\,;{\bm{w}}\right)\mathclose{}$, which defines the variational measure $\mathbb{Q}_{\phi}$, resulting in the ELBO in function space,
$\mathcal{L}(\phi):={\mathop{{}\mathbb{E}}}_{q_{\phi}\mathopen{}\left({\bm{w}}\right)\mathclose{}}\mathopen{}\left[\log p\mathopen{}\left(\mathcal{D}\,|\,{\bm{w}}\right)\mathclose{}\right]\mathclose{}-D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}$ (2.3)
with $D_{\text{KL}}$ the KL divergence between measures given by
[…]
Here, the Raydon-Nikodym derivative $d\mathbb{Q}_{\phi}/d\mathbb{P}$ generalizes the density ratio from Eq. 2.2. Like Eq. 2.1, the ELBO in Eq. 2.3 is a lower bound on the evidence (Burt et al.,, 2020). In fact, if $\mathbb{P}$ is the push-forward of $p({\bm{w}})$ then Eq. 2.3 is a tighter bound than Eq. 2.1 by the data processing inequality, $D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}\leq D_{\mathrm{KL}}\mathopen{}\left(q_{\phi}\,\|\,p\right)\mathclose{}$. However, we motivated function-space VI to avoid weight-space priors, and in this case the bound in Eq. 2.3 can be looser. We will indeed see below that the bound becomes infinitely loose in practice, and we thus propose a different objective in Section 3.
Two intractabilities prevent directly maximizing the ELBO in function space (Eq 2.3). First, it is not obvious how to practically evaluate or estimate the KL divergence between two measures in function space in Eq 2.4. Sun et al., (2019) showed that it can be expressed as a supremum of KL divergences between finite-dimensional distributions,
$D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}=\sup_{{\bm{x}}\in\mathcal{X}^{M}}D_{\mathrm{KL}}\mathopen{}\left(q_{\phi}(f({\bm{x}}))\,\|\,p(f({\bm{x}}))\right)\mathclose{}.$ (2.5)
Here, ${\bm{x}}=\{{\bm{x}}^{(i)}\}_{i=1}^{M}\in\mathcal{X}^{M}$ is a set of $M$ points in feature space $\mathcal{X}$, and $q_{\phi}(f({\bm{x}}))$ and $p(f({\bm{x}}))$ are densities of the marginals of $\mathbb{Q}_{\phi}$ and $\mathbb{P}$ on $\{f({\bm{x}}^{(i)})\}_{i=1}^{M}$ respectively. To obtain a tractable approximation of the supremum over infinitely many sets, Sun et al., (2019) replace it by an expectation and Rudner et al., 2022b estimate it from samples.
Second, we cannot express the pushforward measure $\mathbb{Q}_{\phi}$ of $q_{\phi}({\bm{w}})$ in a closed form because the neural network is nonlinear. Previous work has proposed to mitigate this issue using implicit score function estimators (Sun et al.,, 2019) or a linearized BNN $f_{L}$ to obtain a closed-form Gaussian variational measure (Rudner et al., 2022a, ; Rudner et al., 2022b, ). Our proposal in Section 3 follows the linearized BNN approach as it only minimally modifies the BNN, preserving most of its inductive bias (Maddox et al.,, 2021) while considerably simplifying the problem by turning the pushforward of $q_{\phi}({\bm{w}})$ into a Gaussian process (GP). More specifically, we consider a Gaussian variational distribution $q_{\phi}\mathopen{}\left({\bm{w}}\right)\mathclose{}=\mathcal{N}\mathopen{}\left({\bm{m}},{\bm{S}}\right)\mathclose{}$ with parameters $\phi=\left\{{\bm{m}},{\bm{S}}\right\}$, and we obtain $f_{L}$ by linearizing $f$ in the weights around ${\bm{w}}={\bm{m}}$,
$f_{L}({\bm{x}};{\bm{w}}):=f({\bm{x}};{\bm{m}})+J({\bm{x}};{\bm{m}})({\bm{w}}-{\bm{m}})\qquad\text{where}\qquad J({\bm{x}};{\bm{m}})=\nabla_{\!{\bm{w}}}f({{\bm{x}};{\bm{w}}})|_{{\bm{w}}={\bm{m}}}.$ (2.6)
Thus, for ${\bm{w}}\sim q_{\phi}({\bm{w}})$ we have, $f_{L}({\bm{x}};{\bm{w}})\sim\mathcal{N}\mathopen{}\left(f({\bm{x}};{\bm{m}}),J({\bm{x}};{\bm{m}}){\bm{S}}J({\bm{x}};{\bm{m}})^{\top}\right)\mathclose{}\;\forall{\bm{x}}$, and the function $f_{L}(\,\cdot\,;{\bm{w}})$ is a degenerate GP (as $\operatorname{rank}(J(\,\cdot\,;{\bm{m}}){\bm{S}}J(\,\cdot\,;{\bm{m}})^{\top})\leq\text{number\leavevmode\nobreak\ of\leavevmode\nobreak\ weights}<\infty$),
$f_{L}\sim\mathcal{GP}\mathopen{}\left(f(\,\cdot\,;{\bm{m}}),J(\,\cdot\,;{\bm{m}}){\bm{S}}J(\,\cdot\,;{\bm{m}})^{\top}\right)\mathclose{}.$ (2.7)
$D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}$ is infinite in most practically relevant cases.
Burt et al., (2020) point out an even more severe issue of function-space VI in BNNs: $D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}$ is in fact infinite in many relevant cases, in particular for non-degenerate GP-priors with BNNs. Thus, approximations in these cases are futile. The proof in Burt et al., (2020) is somewhat involved, but the fundamental reason for $D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}=\infty$ is that $\mathbb{Q}_{\phi}$ has support on a finite-dimensional submanifold of the infinite-dimensional function space, while the measure $\mathbb{P}$ induced by a (non-degenerate) GP prior has support on the entire function space. That such a dimensionality mismatch can lead to infinite KL divergence can be seen in a finite-dimensional example: consider the KL-divergence between two Gaussians in $\mathbb{R}^{n}$ for some $n\geq 2$, one of which has support on the entire $\mathbb{R}^{n}$ (i.e., its covariance matrix ${\bm{\Sigma}}_{1}$ has full rank) while the other one has support only on a proper subspace of $\mathbb{R}^{n}$ (i.e., its covariance matrix ${\bm{\Sigma}}_{2}$ is singular). The KL divergence between two multivariate Gaussians has a closed form expression (Eq 2.9 with ${\gamma=0}$) that contains $\log\det\mathopen{}\left({\bm{\Sigma}}_{2}^{-1}{\bm{\Sigma}}_{1}\right)\mathclose{}$, which is infinite for singular ${\bm{\Sigma}}_{2}$.
We find that the fact that $D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}=\infty$ has severe practical consequences even when the KL divergence is only estimated from samples. It naturally explains the stability issues discussed in Appendix D.1 of Sun et al., (2019) (we compare the authors’ solution to this stability issue to our method in Section 3.2). Surprisingly, similar complications arise even in the setup by Rudner et al., 2022b , which performs VI in function space with the pushforward of a weight-space prior. While the KL divergence is technically finite because prior and variational posterior have the same support, numerical errors lead to mismatching supports and thus to stability issues even there.
In summary, the ELBO for VI in BNNs is not well-defined for a large class of interesting function-space priors. In section 3, we propose a solution by using the so-called regularized KL divergence, which we introduce next.
2.2 Regularized KL divergence for Gaussian measures
Our solution to the negative infinite function-space ELBO builds on a regularized KL divergence, which is expressed in terms of Gaussian measures for the variational posterior and prior. We obtain these Gaussian measures from corresponding Gaussian processes (GP). We first discuss under which conditions a GP induces a Gaussian measure, and then present the regularized KL divergence.
Gaussian measures and Gaussian processes
[…]
Definition 2.1 (Gaussian measure, Kerrigan et al., (2023), Definition 1).
Let $(\Omega,\mathcal{B},\mathbb{P})$ be a probability space. A measurable function $F:\Omega\mapsto L^{2}(\mathcal{X},\rho)$ is called a Gaussian random element (GRE) if for any $g\in L^{2}(\mathcal{X},\rho)$ the random variable $\langle g,F\rangle$ has a Gaussian distribution on . For every GRE $F$, there exists a unique mean element $m\in L^{2}(\mathcal{X},\rho)$ and a finite trace linear covariance operator $C:L^{2}(\mathcal{X},\rho)\mapsto L^{2}(\mathcal{X},\rho)$ such that $\langle g,F\rangle\sim\mathcal{N}\mathopen{}\left(\langle g,m\rangle,\langle Cg,g\rangle\right)\mathclose{}$ for all $g\in L^{2}(\mathcal{X},\rho)$. The pushforward of $\mathbb{P}$ along $F$ denoted $\mathbb{P}^{F}:=F_{\#}\mathbb{P}$ is a Gaussian (probability) measure on $L^{2}(\mathcal{X},\rho)$.
Gaussian measures generalize Gaussian distributions to infinite-dimensional spaces where measures do not have associated densities. Following Wild et al., (2022), we notate the Gaussian measure obtained from the GRE $F$ with mean element $m$ and covariance operator $C$ as $\mathbb{P}^{F}:=\mathcal{N}\mathopen{}\left(m,C\right)\mathclose{}$.
[…]
Definition 2.2 (Regularized KL divergence, Quang, (2022) Definition 5).
Let $\nu_{1}=\mathcal{N}\mathopen{}\left(m_{1},C_{1}\right)\mathclose{}$ and $\nu_{2}=\mathcal{N}\mathopen{}\left(m_{2},C_{2}\right)\mathclose{}$ be two Gaussian measures with $m_{1},m_{2}\in L^{2}(\mathcal{X},\rho)$ and $C_{1},C_{2}$ bounded, self-adjoint, positive and trace-class linear operators on $L^{2}(\mathcal{X},\rho)$. Let $\gamma\in$, $\gamma>0$ be fixed. The regularized KL divergence is defined as follows,
$D^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\nu_{1}\,\|\,\nu_{2}\right)\mathclose{}:=$ $\,\frac{1}{2}\langle m_{1}-m_{2},(C_{2}+\gamma\mathbb{I})^{-1}\mathopen{}\left(m_{1}-m_{2}\right)\mathclose{}\rangle+\frac{1}{2}\operatorname{Tr}_{X}\left[\mathopen{}\left(C_{2}+\gamma\mathbb{I}\right)\mathclose{}^{-1}\mathopen{}\left(C_{1}+\gamma\mathbb{I}\right)\mathclose{}-\mathbb{I}\right]$
$-\frac{1}{2}\log\operatorname{det}_{X}\left[\mathopen{}\left(C_{2}+\gamma\mathbb{I}\right)\mathclose{}^{-1}\mathopen{}\left(C_{1}+\gamma\mathbb{I}\right)\mathclose{}\right].$ (2.8)
For any $\gamma>0$, the regularized KL divergence is well-defined and finite, even if the Gaussian measures are singular (Quang,, 2019). It converges to the conventional KL divergence (if it is well-defined) for $\gamma\to 0$ (Quang,, 2022, Theorem 6). Furthermore, and importantly for our application to VI, if two Gaussian measures $\nu_{1}$ and $\nu_{2}$ are induced by GPs $f_{i}\sim\mathcal{GP}(\mu_{i},K_{i})$ for $i=1,2$, then $D^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\nu_{1}\,\|\,\nu_{2}\right)\mathclose{}$ can be consistently estimated (Quang,, 2022) from a finite number $M$ of samples using the estimator
$\hat{D}^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\nu_{1}\,\|\,\nu_{2}\right)\mathclose{}:=$ $\,\frac{1}{2}\mathopen{}\left({\bm{m}}_{1}-{\bm{m}}_{2}\right)\mathclose{}^{\top}({\bm{\Sigma}}^{(\gamma)}_{2})^{-1}\mathopen{}\left({\bm{m}}_{1}-{\bm{m}}_{2}\right)\mathclose{}+\frac{1}{2}\operatorname{Tr}\big{[}({\bm{\Sigma}}_{2}^{(\gamma)})^{-1}{\bm{\Sigma}}^{(\gamma)}_{1}-\mathbb{I}_{M}\big{]}$
$-\frac{1}{2}\log\det\big{[}({\bm{\Sigma}}_{2}^{(\gamma)})^{-1}{\bm{\Sigma}}^{(\gamma)}_{1}\big{]}$ (2.9)
[…]
3 Generalized function-space VI with the regularized KL divergence
This section presents our proposed generalized function-space variational inference (GFSVI) method, which addresses the problem of the infinite KL divergence discussed in section 2.1, which we take for an indication that VI is too restrictive if one wants to use genuine function-space priors. We instead consider generalized variational inference (Knoblauch et al.,, 2019), which reinterprets the ELBO in Eq. 2.1 as a regularized expected log-likelihood and explores alternative divergences for the regularizer. Specifically, we propose to use the regularized KL divergence. The presentation in this section builds heavily on tools introduced in section 2, which turn out to fit together perfectly: the pushforward of a Gaussian variational distribution in weight-space through the linearized neural network (Eq. 2.6) induces a GP variational posterior (Eq. 2.7) that admits a Gaussian measure on $L^{2}(\mathcal{X},\rho)$ (definition 2.1). Further selecting a GP prior which has an associated Gaussian measure on $L^{2}(\mathcal{X},\rho)$ allows us to use the regularized KL divergence (Eq. 2.8). We present the GFSVI method in section 3.1 and discuss connections to prior work in section 3.2.
3.1 Generalized function-space variational inference
We present a simple and well-defined objective for function-space inference, and an algorithm for its optimization.
Objective function
We start from the ELBO in Eq. 2.3, where we use the Gaussian variational measure $\mathbb{Q}_{\phi}^{F}$ induced by the pushforward of a Gaussian variational distribution $q_{\phi}({\bm{w}})=\mathcal{N}\mathopen{}\left({\bm{w}}\,|\,{\bm{m}},{\bm{S}}\right)\mathclose{}$ along the linearized network $f_{L}$ (Eq. 2.6). The function-space prior may be any GP that has an associated Gaussian measure $\mathbb{P}^{F}$ on $L^{2}(\mathcal{X},\rho)$ (see definition 2.1 and discussion below it). We now replace the KL divergence in the ELBO with the regularized KL divergence $D_{\text{KL}}^{\gamma}$ (Eq. 2.8), which is well-defined and finite for any pair of Gaussian measures. For a likelihood $p\mathopen{}\left(\mathcal{D}\,|\,{\bm{w}}\right)\mathclose{}=\prod_{i=1}^{N}p\mathopen{}\left(y_{i}\,|\,f_{L}({\bm{x}}_{i};{\bm{w}})\right)\mathclose{}$, we thus obtain
$\mathcal{L}(\phi)\!:=\!\!\sum_{i=1}^{N}{\mathop{{}\mathbb{E}}}_{q_{\phi}({\bm{w}})}\mathopen{}\left[\log p\mathopen{}\left(y_{i}|f_{L}({\bm{x}}_{i};{\bm{w}})\right)\mathclose{}\right]\mathclose{}\!-\!D^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}^{F}\,\|\,\mathbb{P}^{F}\right)\mathclose{}.$ (3.1)
Input: linearized BNN $f_{L}$ with measure $\mathbb{Q}_{\phi}^{F}$, Gaussian process prior $\mathcal{GP}(\mu,K)$ with measure $\mathbb{P}^{F}$, measurement point distribution $\rho({\bm{x}})$, data $\mathcal{D}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}$, batch size $B$, $\gamma>0$.
for each minibatch $({\bm{x}}_{\mathcal{B}},y_{\mathcal{B}})\sim\mathcal{D}$ do
       Calculate ^$\hat{\ell}_{1}=\frac{N}{B}{\mathop{{}\mathbb{E}}}_{q_{\phi}({\bm{w}})}\mathopen{}\left[\log p\mathopen{}\left(y_{\mathcal{B}}\,|\,f_{L}({\bm{x}}_{\mathcal{B}},{\bm{w}})\right)\mathclose{}\right]\mathclose{}$;
      [-1.5ex] Draw measurement points ${\bm{x}}=\{{\bm{x}}^{(i)}\}_{i=1}^{M}\stackrel{{\scriptstyle\!\!\text{i.i.d.}\!\!}}{{\sim}}\rho({\bm{x}})$;
      [-1.5ex] Calculate ^$\hat{\ell}_{2}=\hat{D}_{\text{KL}}^{\gamma}\big{(}\mathbb{Q}_{\phi}^{F}\,\big{|}\!\big{|}\,{\mathbb{P}^{F}}\big{)}$ using ${\bm{x}}$ (Eq. 2.9);
      [-1.5ex] Calculate ^$\hat{\mathcal{L}}(\phi)=\hat{\ell}_{1}-\hat{\ell}_{2}$;
       Update $\phi$ using a step in the direction $\nabla_{\!\phi}\hat{\mathcal{L}}(\phi)$;
end for
Algorithm 1 Generalized function-space variational inference (GFSVI)
Estimation and optimization
[…]
Technical details ($\gamma$ and $\rho$)
It turns out that increasing $\gamma$ reduces the influence of the prior on inference (see figure 17 in Appendix). At the same time, $\gamma$ acts as jitter that prevents numerical errors (see section 3.2). We recommend setting $\gamma$ large enough to avoid numerical errors but sufficiently small to strongly regularize the objective in Eq. 3.1 (see figure 15 in Appendix). We found that GFSVI is robust to a wide range of values and fixed $\gamma=10^{-10}$. Furthermore, the probability measure $\rho$ defined with $L^{2}(\mathcal{X},\rho)$ has to assign non-zero probability to any open set of $\mathcal{X}$ to regularize the BNN on all of its support. Following Rudner et al., 2022b , we draw measurement points from a uniform distribution over $\mathcal{X}$ when using tabular data and explore different configurations (such as samples from other data sets) when using high-dimensional image data (see section A.3.4).
3.2 Connections to prior work
TFSVI (Rudner et al., 2022b, ) and FVI (Sun et al.,, 2019) solve stability issues by introducing jitter/noise, which has a similar effect as the regularization in Eq. 2.8. However, as mentioned in Section 2.1, TFSVI introduces jitter only to overcome numerical issues and is fundamentally restricted to prior specification in weight space since its function-space prior is the pushforward of a weight-space prior. FVI does not linearize the BNN, and therefore does not have access to an explicit variational measure in function space. This severely complicates the estimation of (gradients of) the KL divergence, and the authors resort to implicit score function estimators, which make their method difficult to use in practice (Ma and Hernández-Lobato,, 2021). Our proposed GFSVI does not suffer from these difficulties as the variational posterior is an explicit (Gaussian) measure. This allows us to estimate the (regularized) KL divergence without sampling any noise or using implicit score function estimators.
4 Experiments
Figure 2: Inference on synthetic data (gray circles) using an RBF prior for function-space methods GFSVI and FVI. The proposed GFSVI provides the best approximation of the exact GP posterior.
In this section, we evaluate our generalized function-space variational inference (GFSVI) method qualitatively on synthetic data (Section 4.1) and quantitatively on real-world data (Section 4.2). We find that GFSVI accurately captures structural properties specified by the GP prior, that it approximates the exact Gaussian process (GP) posterior more faithfully than BNN baselines, and that it performs competitively on regression, classification and out-of-distribution detection tasks. We also discuss the influence of the BNN’s inductive biases.
Baselines
We compare the proposed GFSVI method to two weight-space inference methods (mean-field variational inference (Blundell et al.,, 2015) and linearized Laplace (Immer et al.,, 2021)) and to three function-space inference methods (FVI (Sun et al.,, 2019), TFVSI (Rudner et al., 2022b, ) and VIP (Ma et al.,, 2019), where TFSVI performs inference in function space but with the pushforward of a weight-space prior and VIP using a BNN prior). All BNNs have the same architecture and fully-factorized Gaussian approximate posterior. We also include results for a sparse GP with a posterior mean parameterized by a neural network (GWI) (Wild et al.,, 2022), as well as for a Gaussian Process (GP) (Williams and Rasmussen,, 2006) (when the size of the dataset allows it), and for a sparse GP (Hensman et al.,, 2013) for regression tasks. Ẇe consider the GP, sparse GP and GWI as gold standards as they represent the exact (or near exact) posterior for models with GP priors.
4.1 Qualitative results on synthetic data
Regression
We consider synthetic regression data with a one-dimensional feature-space $\mathcal{X}\subset$, where the values $y_{i}$ are randomly sampled around $\sin(2\pi x_{i})$ (circles in Figures 1-11, see Appendix A.3.1). The green lines in Figures 1-8 and 11 show functions sampled from the (approximate) posteriors, and the red lines are the inferred mean functions. Figure 2 compares GFSVI with an RBF GP prior to baselines and to the exact posterior. We find that GFSVI visually matches the true posterior best (the same holds for a Matern-1/2 prior, see Figure 5 in the Appendix). Figure 1 in the Introduction and Figure 6 in the Appendix show that GFSVI notably adapts to varying prior assumptions (varying smoothness and length scale, respectively). In addition, Figures 4 and 8 in the Appendix show that GFSVI provides strong regularization when the data generative process is noisy, and that it can be trained with fewer measurement points $M$ than FVI without significant degradation.
Classification
We further consider a binary classification task derived from the 2-dimensional two moons data set (Pedregosa et al.,, 2011), see red and blue dots in Figures 9 and 10 in the Appendix. The first and second row in both figures show the inferred mean probability of class 1 (blue dots) and its 2-standard deviations with respect to posterior samples. Just like for regression, we find that GFSVI captures the beliefs of the RBF and Matern-1/2 GP priors better than all BNN-baselines, and shows greater uncertainty outside of the support of the data than the BNN baselines.
Inductive biases
figure 11 in the Appendix compares GFSVI to the exact GP-posterior across two different priors and three model architectures (details in Appendix A.3.1). We find that, with ReLU activations, small models are prone to underfitting for smooth priors (RBF), and to collapsing uncertainty for rough priors (Matern-1/2). By contrast, with smooth activations (Tanh), smaller models suffice, and they are compatible with most standard GP priors (the results shown in figure 11 extend to RBF, Matern, and Rational Quadratic in our experiments). We also analyzed how the number $M$ of measurement points affects performance. figures 7 and 14 in Appendix show that capturing the beliefs of rough GP priors and estimating $D_{\text{KL}}^{\gamma}$ with these priors requires larger $M$.
Table 1: Test expected log-likelihood (higher is better) of evaluated methods on regression datasets. GFSVI performs competitively compared to all BNN baselines and obtains the best mean rank.
Dataset Function-space priors Weight-space priors Gaussian Processes (Gold Standards)
GFSVI (ours) FVI TFSVI MFVI VIP Laplace GWI Sparse GP GP
Boston -0.733 $\pm$ 0.144 -0.571 $\pm$ 0.113 -1.416 $\pm$ 0.046 -1.308 $\pm$ 0.052 -0.722 $\pm$ 0.196 -0.812 $\pm$ 0.205 -0.940 $\pm$ 0.145 -0.884 $\pm$ 0.182 -1.594 $\pm$ 0.556
Concrete -0.457 $\pm$ 0.041 -0.390 $\pm$ 0.017 -0.983 $\pm$ 0.012 -1.353 $\pm$ 0.018 -0.427 $\pm$ 0.050 -0.715 $\pm$ 0.025 -0.744 $\pm$ 0.079 -0.966 $\pm$ 0.025 -2.099 $\pm$ 0.421
Energy -1.319 $\pm$ 0.052 -1.377 $\pm$ 0.042 -0.797 $\pm$ 0.098 -0.926 $\pm$ 0.197 -1.046 $\pm$ 0.378 -1.304 $\pm$ 0.043 -0.461 $\pm$ 0.093 -0.206 $\pm$ 0.027 -0.205 $\pm$ 0.022
Kin8nm -0.136 $\pm$ 0.013 -0.141 $\pm$ 0.023 -0.182 $\pm$ 0.011 -0.641 $\pm$ 0.225 -0.102 $\pm$ 0.013 -0.285 $\pm$ 0.014 -0.708 $\pm$ 0.054 -0.443 $\pm$ 0.014 (infeasible)
Naval -3.637 $\pm$ 0.132 -2.165 $\pm$ 0.194 -2.758 $\pm$ 0.044 -1.034 $\pm$ 0.160 -1.502 $\pm$ 0.061 -3.404 $\pm$ 0.084 -0.301 $\pm$ 0.254 -4.951 $\pm$ 0.014 (infeasible)
Power -0.044 $\pm$ 0.011 -0.031 $\pm$ 0.021 -0.007 $\pm$ 0.013 -0.003 $\pm$ 0.015 -0.036 $\pm$ 0.018 -0.002 $\pm$ 0.019 -0.043 $\pm$ 0.009 -0.100 $\pm$ 0.010 (infeasible)
Protein -1.036 $\pm$ 0.005 -1.045 $\pm$ 0.005 -1.010 $\pm$ 0.004 -1.112 $\pm$ 0.007 -0.994 $\pm$ 0.007 -1.037 $\pm$ 0.006 -1.050 $\pm$ 0.009 -1.035 $\pm$ 0.002 (infeasible)
Wine -1.289 $\pm$ 0.040 -1.215 $\pm$ 0.007 -2.138 $\pm$ 0.221 -1.248 $\pm$ 0.018 -1.262 $\pm$ 0.025 -1.249 $\pm$ 0.025 -1.232 $\pm$ 0.038 -1.240 $\pm$ 0.037 -1.219 $\pm$ 0.035
Yacht -1.058 $\pm$ 0.080 -0.545 $\pm$ 0.735 -1.187 $\pm$ 0.064 -1.638 $\pm$ 0.030 -0.062 $\pm$ 1.378 -0.680 $\pm$ 0.171 -0.441 $\pm$ 0.138 -0.976 $\pm$ 0.092 -0.914 $\pm$ 0.045
Wave -5.521 $\pm$ 0.036 -6.612 $\pm$ 0.008 -5.148 $\pm$ 0.117 -6.883 $\pm$ 0.008 -4.043 $\pm$ 0.093 -4.658 $\pm$ 0.027 -1.566 $\pm$ 0.123 -4.909 $\pm$ 0.001 (infeasible)
Denmark -0.487 $\pm$ 0.013 -0.801 $\pm$ 0.005 -0.513 $\pm$ 0.013 -0.675 $\pm$ 0.007 -0.583 $\pm$ 0.021 -0.600 $\pm$ 0.008 -0.841 $\pm$ 0.026 -0.768 $\pm$ 0.001 (infeasible)
Mean rank 1.545 2.000 2.727 3.455 2.091 2.455 - - -
Table 2: Test expected log-likelihood, accuracy, expected calibration error and OOD detection accuracy on MNIST and Fashion MNIST.
Data Metric Function-space priors Weight-space priors GP-based
GFSVI (rnd) GFSVI (kmnist) FVI (rnd) FVI (kmnist) TFSVI (rnd) TFSVI (kmnist) MFVI VIP Laplace GWI
MNIST Log-like. ($\uparrow$) -0.033 $\pm$ $0$0.041 $\pm$ $0$0.145 $\pm$ 0.005 -0.238 $\pm$ 0.006 -0.047 $\pm$ 0.003 -0.041 $\pm$ 0.001 -0.078 $\pm$ 0.001 -0.033 $\pm$ 0.001 -0.108 $\pm$ 0.002 -0.090 $\pm$ 0.003
Acc. ($\uparrow$) -0.992 $\pm$ $0$0.991 $\pm$ $0$0.976 $\pm$ 0.001 -0.943 $\pm$ 0.001 -0.989 $\pm$ $0$0.989 $\pm$ $0$0.990 $\pm$ $0$0.989 $\pm$ $0$0.984 $\pm$ $0$0.971 $\pm$ 0.001
ECE ($\downarrow$) -0.002 $\pm$ $0$0.006 $\pm$ $0$0.064 $\pm$ 0.001 -0.073 $\pm$ 0.003 -0.007 $\pm$ $0$0.006 $\pm$ $0$0.021 $\pm$ $0$0.002 $\pm$ 0.001 -0.048 $\pm$ 0.001 -0.003 $\pm$ 0.000
OOD acc. ($\uparrow$) -0.921 $\pm$ 0.008 -0.980 $\pm$ 0.004 -0.894 $\pm$ 0.010 -0.891 $\pm$ 0.006 -0.887 $\pm$ 0.011 -0.893 $\pm$ 0.005 -0.928 $\pm$ 0.002 -0.871 $\pm$ 0.012 -0.903 $\pm$ 0.007 -0.829 $\pm$ 0.007
FMNIST Log-like. ($\uparrow$) -0.260 $\pm$ 0.003 -0.294 $\pm$ 0.006 -0.300 $\pm$ 0.002 -0.311 $\pm$ 0.005 -0.261 $\pm$ 0.001 -0.261 $\pm$ 0.002 -0.290 $\pm$ 0.002 -0.252 $\pm$ 0.001 -0.426 $\pm$ 0.009 -0.260 $\pm$ 0.001
Acc. ($\uparrow$) -0.910 $\pm$ 0.001 -0.909 $\pm$ 0.001 -0.910 $\pm$ 0.002 -0.906 $\pm$ 0.002 -0.909 $\pm$ 0.001 -0.907 $\pm$ 0.001 -0.913 $\pm$ 0.001 -0.911 $\pm$ 0.001 -0.886 $\pm$ 0.001 -0.906 $\pm$ 0.000
ECE ($\downarrow$) -0.020 $\pm$ 0.003 -0.042 $\pm$ 0.002 -0.027 $\pm$ 0.005 -0.024 $\pm$ 0.002 -0.022 $\pm$ 0.002 -0.021 $\pm$ 0.002 -0.010 $\pm$ 0.001 -0.024 $\pm$ 0.001 -0.060 $\pm$ 0.004 -0.016 $\pm$ 0.001
OOD acc. ($\uparrow$) -0.853 $\pm$ 0.005 -0.997 $\pm$ 0.001 -0.925 $\pm$ 0.005 -0.975 $\pm$ 0.002 -0.802 $\pm$ 0.006 -0.779 $\pm$ 0.010 -0.805 $\pm$ 0.010 -0.790 $\pm$ 0.010 -0.826 $\pm$ 0.006 -0.792 $\pm$ 0.005
4.2 Quantitative results on real-world data
We study the fidelity of GFSVI compared to the true GP-posterior on real data, and evaluate GFSVI on regression, classification, and out-of-distribution detection tasks. In all tables, we bold the highest score and any score whose error bar (standard error) overlaps with the highest score’s error bar.
Ocean current modeling
Figure 3: Results for the ocean current modeling experiment. We report the mean velocity vectors and the squared errors of compared methods. Unlike TFSVI, we find that GFSVI accurately captures ocean current dynamics.
We test the extend to which GFSVI can incorporate prior knowledge specified via a GP prior by considering the problem of modeling ocean currents in the Gulf of Mexico. We are given $20$ two-dimensional velocity vectors collected from ocean drifter buoys, and we are interested in estimating ocean currents. We use the GulfDrifters dataset (Lilly and Pérez-Brunius,, 2021) and we follow the setup by Shalashilin, (2024). We compare our GFSVI to a GP and to TFSVI. We incorporate known physical properties of fluid motions by applying the Helmholtz decomposition to the GP prior (Berlinghieri et al.,, 2023) as well as to the neural networks. More details can be found in section A.3.2.
Table 3: Results for the ocean current modeling task. We find that incorporating prior knowledge via a Gaussian process prior in GFSVI improves score over weight-space priors in TFSVI.
Dataset GFSVI (ours) TFSVI GP
Log-like. -6.627 $\pm$ 0.753 -22.651 $\pm$ 2.947 -0.507 $\pm$ 0.000
MSE -0.021 $\pm$ 0.002 -0.034 $\pm$ 0.003 -0.013 $\pm$ 0.000
Results are shown in table 3 and in figure 3. We find that incorporating prior knowledge via an informative GP prior in GFSVI improves expected log-likelihood (Log-like.) and mean squared-error (MSE) over weight-space priors in TFSVI. However, for this relatively small dataset, exact GP inference is still possible, and it outperforms both BNN-based predictions. This suggests that the physically motivated GP describes the true fluid dynamics well enough that the additional inductive bias introduced by a neural network hurts performance rather than helping it. In the following, we consider experiments with larger datasets (making exact GP inference computationally infeasible in many cases), and where structural prior knowledge in function space exists but is not derived from strict laws of nature.
Regression
We assess the predictive performance of GFSVI on data sets from the UCI repository (Dua and Graff,, 2017) (described in table 6 in the Appendix). Table 1, and Table 7 in the Appendix, show expected log-likelihood and mean squared error, respectively. We perform 5-fold cross validation and report means and standard errors across the test folds. We also rank the methods for each dataset and report the mean rank of each method across all datasets. See section A.3.3 for more details. We find that GFSVI performs competitively compared to baselines and obtains the best mean rank for both metrics, matching the top performing methods on nearly all datasets. In particular, we find that using GP priors in the linearized BNN with GFSVI yields improvements over the weight-space priors used in TFSVI and that GFSVI performs slightly better than FVI.
Table 4: Out-of-distribution accuracy (higher is better) of evaluated methods on regression datasets. GFSVI (ours) performs competitively on OOD detection and obtains the highest mean rank.
Dataset Function-space priors Weight-space priors Gaussian Processes (Gold Standards)
GFSVI (ours) FVI TFSVI MFVI VIP Laplace GWI Sparse GP GP
Boston 0.893 $\pm$ 0.011 0.594 $\pm$ 0.024 0.705 $\pm$ 0.107 0.563 $\pm$ 0.013 0.628 $\pm$ 0.010 0.557 $\pm$ 0.009 0.817 $\pm$ 0.017 0.947 $\pm$ 0.011 0.952 $\pm$ 0.003
Concrete 0.656 $\pm$ 0.016 0.583 $\pm$ 0.022 0.511 $\pm$ 0.003 0.605 $\pm$ 0.012 0.601 $\pm$ 0.024 0.578 $\pm$ 0.015 0.730 $\pm$ 0.020 0.776 $\pm$ 0.006 0.933 $\pm$ 0.004
Energy 0.997 $\pm$ 0.002 0.696 $\pm$ 0.017 0.997 $\pm$ 0.001 0.678 $\pm$ 0.014 0.682 $\pm$ 0.037 0.782 $\pm$ 0.020 0.998 $\pm$ 0.001 0.998 $\pm$ 0.001 0.998 $\pm$ 0.001
Kin8nm 0.588 $\pm$ 0.007 0.604 $\pm$ 0.023 0.576 $\pm$ 0.008 0.570 $\pm$ 0.009 0.563 $\pm$ 0.015 0.606 $\pm$ 0.009 0.602 $\pm$ 0.011 0.608 $\pm$ 0.014 (infeasible)
Naval 1.000 $\pm$ 0.000 1.000 $\pm$ 0.000 1.000 $\pm$ 0.000 0.919 $\pm$ 0.017 0.621 $\pm$ 0.059 1.000 $\pm$ 0.000 1.000 $\pm$ 0.000 1.000 $\pm$ 0.000 (infeasible)
Power 0.698 $\pm$ 0.006 0.663 $\pm$ 0.021 0.676 $\pm$ 0.008 0.636 $\pm$ 0.019 0.514 $\pm$ 0.004 0.654 $\pm$ 0.013 0.754 $\pm$ 0.004 0.717 $\pm$ 0.004 (infeasible)
Protein 0.860 $\pm$ 0.011 0.810 $\pm$ 0.022 0.841 $\pm$ 0.018 0.693 $\pm$ 0.020 0.549 $\pm$ 0.020 0.629 $\pm$ 0.013 0.942 $\pm$ 0.002 0.967 $\pm$ 0.001 (infeasible)
Wine 0.665 $\pm$ 0.013 0.517 $\pm$ 0.004 0.549 $\pm$ 0.015 0.542 $\pm$ 0.009 0.706 $\pm$ 0.028 0.531 $\pm$ 0.007 0.810 $\pm$ 0.008 0.781 $\pm$ 0.014 0.787 $\pm$ 0.007
Yacht 0.616 $\pm$ 0.030 0.604 $\pm$ 0.025 0.659 $\pm$ 0.043 0.642 $\pm$ 0.035 0.688 $\pm$ 0.040 0.612 $\pm$ 0.024 0.563 $\pm$ 0.014 0.762 $\pm$ 0.018 0.787 $\pm$ 0.011
Wave 0.975 $\pm$ 0.005 0.642 $\pm$ 0.004 0.835 $\pm$ 0.034 0.658 $\pm$ 0.026 0.500 $\pm$ 0.000 0.529 $\pm$ 0.005 0.903 $\pm$ 0.001 0.513 $\pm$ 0.001 (infeasible)
Denmark 0.521 $\pm$ 0.006 0.612 $\pm$ 0.008 0.519 $\pm$ 0.006 0.513 $\pm$ 0.003 0.500 $\pm$ 0.000 0.529 $\pm$ 0.008 0.688 $\pm$ 0.003 0.626 $\pm$ 0.002 (infeasible)
Mean rank 1.455 2.364 1.909 2.909 3.364 2.909 - - -
Table 5: Average point-wise Wasserstein-2 distance (lower is better) between exact and approximate posterior of reported methods. GFSVI (ours) provides a more accurate approximation than FVI.
Dataset Boston Concrete Energy Wine Yacht Mean rank
GFSVI (ours) 0.0259 $\pm$ 0.0040 0.0499 $\pm$ 0.0029 0.0035 $\pm$ 0.0004 0.0571 $\pm$ 0.0097 0.0036 $\pm$ 0.0006 1.0
FVI 0.0469 $\pm$ 0.0044 0.0652 $\pm$ 0.0037 0.0037 $\pm$ 0.0004 0.1224 $\pm$ 0.0167 0.0052 $\pm$ 0.0013 1.6
GP sparse 0.0074 $\pm$ 0.0022 0.0125 $\pm$ 0.0016 0.0042 $\pm$ 0.0003 0.0170 $\pm$ 0.0035 0.0035 $\pm$ 0.0008 -
Classification
We further evaluate the classification performance of our method on the MNIST (LeCun et al.,, 2010) and FashionMNIST (Xiao et al.,, 2017) image data sets. We fit the models on a random subset of 90% of the training set, use the remaining 10% as validation data, and evaluate on the provided test split. We repeat with 5 different random seeds and report the mean and standard error of the expected log-likelihood, accuracy and expected calibration error (ECE) of the predictive distribution’s mean in table 2. For GFSVI, FVI, and TFSVI, we tested measurement points from both a random (rnd) distribution $\rho({\bm{x}})$ and from kmnist (see Algorithm 1 and section A.3.4). We find that GFSVI performs competitively on MNIST, exceeding the expected log-likelihood and accuracy of top-scoring baselines (see table 2) and similarly to best baselines on FashionMNIST. Additionally, GFSVI yields well-calibrated models with low ECE.
Out-of-distribution detection
We further evaluate our method by testing if its epistemic uncertainty is predictive of out-of-distribution (OOD) data. We consider two settings: (i) with tabular data and a Gaussian likelihood following Malinin et al., (2020), and (ii) with image data and a categorical likelihood following Osawa et al., (2019). We report the accuracy of classifying OOD from in-distribution (ID) data using a (learned) threshold on the predictive uncertainty. Additional details are provided in section A.3.5. In setting (i) (table 4), we find that GFSVI performs competitively and obtains the highest mean rank. Likewise, in setting (ii) (table 8 in the Appendix), GFSVI strongly outperforms all baselines when using the kmnist measurement point distribution $\rho({\bm{x}})$. We find that with high-dimensional image data, the choice of measurement point distribution highly influences OOD detection accuracy. We provide more details on this point in Appendix A.4.4. In both settings, we find that using GP priors with GFSVI rather than weight-space priors with TFSVI is beneficial, and that GFSVI improves over FVI.
Variational measure evaluation
table 5 evaluates our inference method by comparing samples drawn from the exact posterior (where computationally feasible) with the approximate posterior obtained with our method (GFSVI). More details are provided in section A.3.6. We find that GFSVI approximates the exact posterior more accurately that FVI, obtaining a higher mean rank, but worse than the gold standard sparse GP, which demonstrates to be most accurate.
5 Related work
In this section, we review related work on function-space variational inference with neural networks, and on approximating functions-space measures with weight-space priors.
Function-space inference with neural networks
Prior work on function-space VI in BNNs has addressed issues (i) and (ii) mentioned in Section 2.1. Sun et al., (2019) address (i) (intractable variational posterior in function space) by using implicit score function estimators, and (ii) (intractable KL divergence) by replacing the supremum with an expectation. Rudner et al., 2022a ; Rudner et al., 2022b address (i) by using a linearized BNN (Khan et al.,, 2020; Immer et al.,, 2021; Maddox et al.,, 2021), and (ii) by replacing the supremum with a maximum over a finite set of samples. Other work abandons approximating the posterior over neural network weights altogether and instead uses a BNN only to specify a prior (Ma et al.,, 2019), or deterministic neural networks to fit basis functions for Bayesian linear regression (Ma and Hernández-Lobato,, 2021) or the mean of a generalized sparse GP with Wasserstein-2 metric (Wild et al.,, 2022). Our work combines linearized BNNs with generalized variational inference, but we use the regularized KL divergence (Quang,, 2019), which naturally generalizes the conventional KL divergence and allows for informative GP priors.
Approximating function-space measures with weight-space priors in BNNs
Flam-Shepherd et al., (2017); Tran et al., (2022) minimize a divergence between the BNN’s prior predictive and a GP before performing inference on weights, while Wu et al., (2023) directly incorporate the bridging divergence inside the inference objective. Alternatively, Pearce et al., (2020) derive BNN architectures mirroring GPs, and Matsubara et al., (2022) use the Ridgelet transform to design weight-spaces priors approximating a GP in function space. Another line of work considers weight-space priors which regularize in function space by comparing the model’s predictions to those of a reference model (Nalisnick et al.,, 2020) and using an empirical prior (Rudner et al.,, 2023).
6 Discussion
We proposed a simple inference method with a well-defined variational objective function for Bayesian neural networks with Gaussian process (GP) priors in function-space. As standard VI with functions-space priors suffers from an infinite KL divergence problem, we propose to follow the generalized VI framework. Specifically, we substitute the conventional KL divergence in the ELBO by the regularized KL divergence, which is always finite, and which can be estimated consistently within the linearized BNN approximation. We demonstrated that our method allows to incorporate interpretable structural properties via a GP prior, accurately approximates the true GP posterior on synthetic and small real-world data sets, and provides competitive uncertainty estimates for regression, classification and out-of-distribution detection compared to BNN baselines with both function-space and weight-space priors.
Future work should investigate the use of more expressive variational distributions, such as Gaussian with low-rank plus diagonal covariance proposed by Tomczak et al., (2020), which is compatible with our proposed method.
Acknowledgments
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645. Additional support was provided by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039A. Robert Bamler acknowledges funding by the German Research Foundation (DFG) for project 448588364 of the Emmy Noether Programme. The authors extend their gratitude to the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Tristan Cinquin. Finally, we thank Vincent Fortuin and Marvin Pförtner for useful discussions and feedback.
Appendix A Appendix
A.1 Divergences between Gaussian measures
A.1.1 The KL divergence is infinite
In this section, we show that the Kullbach-Liebler (KL) divergence between the Gaussian measures $\mathbb{Q}_{\phi}^{F}\sim\mathcal{N}\mathopen{}\left(m_{Q},C_{Q}\right)\mathclose{}$ and $\mathbb{P}^{F}\sim\mathcal{N}\mathopen{}\left(m_{P},C_{P}\right)\mathclose{}$, respectively induced by the linearized BNN in Eq 2.7 and by a non-degenerate Gaussian process satisfying conditions given in section 2.2, is infinite. While this has already been shown by Burt et al., (2020), the proof is easier for Gaussian measures. We first need the Feldman-Hàjek theorem which tells us when the KL divergence between two Gaussian measures is well-defined.
Theorem A.1 (Feldman-Hàjek, Quang, (2022) Theorem 2, Simpson, (2022) Theorem 7).
Consider two Gaussian measures $\nu_{1}=\mathcal{N}\mathopen{}\left(m_{1},C_{1}\right)\mathclose{}$ and $\nu_{2}=\mathcal{N}\mathopen{}\left(m_{2},C_{2}\right)\mathclose{}$ on $L^{2}(\mathcal{X},\rho)$. Then $\nu_{1}$ and $\nu_{2}$ are called equivalent if and only if the following holds:
  1. 1.
    $m_{1}-m_{2}\in\text{Im}(C_{2}^{1/2})$
  2. 2.
    The operator $T$ such that $C_{1}=C_{2}^{1/2}(I-T)C_{2}^{1/2}$ is Hilbert-Schmidt, that is $T$ has a countable set of eigenvalues $\lambda_{i}$ that satisfy $\lambda_{i}<1$ and $\sum_{i=1}^{\infty}\lambda_{i}^{2}<\infty$.
otherwise $\nu_{1}$ and $\nu_{2}$ are singular. If $\nu_{1}$ and $\nu_{2}$ are equivalent, then the Radon-Nikodym dervative exists and $D_{\mathrm{KL}}\mathopen{}\left(\nu_{1}\,\|\,\nu_{2}\right)\mathclose{}$ admits an explicit formula. Otherwise, $D_{\mathrm{KL}}\mathopen{}\left(\nu_{1}\,\|\,\nu_{2}\right)\mathclose{}=\infty$.
Let us now show that the KL divergence between $\mathbb{Q}_{\phi}^{F}$ and $\mathbb{P}^{F}$ is indeed infinite.
Proposition 1.
The Gaussian measures $\mathbb{Q}_{\phi}^{F}$ and $\mathbb{P}^{F}$ are mutually singular and $D_{\text{KL}}(\mathbb{Q}_{\phi}^{F}||\mathbb{P}^{F})=\infty$.
Proof.
The proof follows from the Feldman-Hàjek theorem (theorem A.1). In our case, $C_{Q}$ has at most $p$ non-zero eigenvalues as the covariance function of the GP induced by the BNN is degenerate, while $C_{P}$ has a set of (countably) infinite non-zeros eigenvalues (prior is non-degenerate as per assumption). Hence, for the equality in condition (2) to hold, $T$ must have eigenvalue $1$ which violates the requirement that $T$ is Hilbert-Schmidt i.e. that its eigenvalues $\left\{\lambda_{i}\right\}_{i=1}^{\infty}$ satisfy $\lambda_{i}<1$ and $\sum_{i=1}^{\infty}\lambda_{i}^{2}<\infty$. Therefore, $\mathbb{Q}_{\phi}^{F}$ and $\mathbb{P}^{F}$ are mutually singular and $D_{\text{KL}}(\mathbb{Q}_{\phi}^{F}||\mathbb{P}^{F})=\infty$. ∎
A.1.2 The regularized KL divergence
In this section, we provide the bound describing the asymptotic convergence of the regularized KL divergence estimator.
Theorem A.2 (Convergence of estimator, Quang, (2022) Theorem 45).
Assume the following:
  1. 1.
    Let T be a $\sigma-\text{compact}$ metric space, that is $T=\cup_{i=1}^{\infty}T_{i}$, where $T_{1}\subset T_{2}\subset\cdots$ with each $T_{i}$ being compact.
  2. 2.
    $\rho$ is a non-degenerate Borel probability measure on T, that is $\rho\mathopen{}\left(B\right)\mathclose{}>0$ for each open set $B\subset T$.
  3. 3.
[…]
  4. 4.
    $\operatorname{sup}_{x\in T}K_{i}(x,x)\leq\kappa_{i}^{2}\;$ for $i=1,2$.
  5. 5.
    $f_{i}\sim GP\mathopen{}\left(\mu_{i},K_{i}\right)\mathclose{}$, where $\mu_{i}\in L^{2}(T,\rho)$ for $i=1,2$.
  6. 6.
    $\exists B_{i}>0$ such that $\left\lVert\mu_{i}\right\rVert_{\infty}\leq B_{i}$ for $i=1,2$.
[…]
$|D_{\mathrm{KL}}\mathopen{}\left(\mathcal{N}\mathopen{}\left(\mu_{1}(\mathbf{x}),K_{1}(\mathbf{x},\mathbf{x})+M\gamma\mathbb{I}_{M}\right)\mathclose{}\,\|\,{\mathcal{N}\mathopen{}\left(\mu_{2}(\mathbf{x}),K_{2}(\mathbf{x},\mathbf{x})+M\gamma\mathbb{I}_{M}\right)\mathclose{}}\right)\mathclose{}\\ -D^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\mathcal{N}\mathopen{}\left(m_{1},C_{1}\right)\mathclose{}\,\|\,\mathcal{N}\mathopen{}\left(m_{2},C_{2}\right)\mathclose{}\right)\mathclose{}|\\ \leq\frac{1}{2\gamma}(B_{1}+B_{2})^{2}[1+\kappa_{2}^{2}/\gamma]\mathopen{}\left(\frac{2\log\frac{48}{\gamma}}{M}+\sqrt{\frac{2\log\frac{48}{\gamma}}{M}}\right)\mathclose{}\\ +\frac{1}{2\gamma^{2}}[\kappa_{1}^{4}+\kappa_{2}^{4}+\kappa_{1}^{2}\kappa_{2}^{2}(2+\kappa_{2}^{2}/\gamma)]\mathopen{}\left(\frac{2\log\frac{12}{\gamma}}{M}+\sqrt{\frac{2\log\frac{12}{\gamma}}{M}}\right)\mathclose{}$ (A.1)
A.2 Additional details on the GFSVI objective estimator
In this section, we present details on the estimation of the generalized function-space variational inference (GFSVI) objective. Let $f_{L}({\,\cdot\,;{\bm{w}}})$ be the linearized BNN (Eq 2.6) with weights ${\bm{w}}\in^{p}$, and $\mathcal{D}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}$ a data set with features ${\bm{x}}_{i}\in\mathcal{X}\subset^{d}$ and associated values $y_{i}\in\mathcal{Y}$. Assuming a likelihood $p\mathopen{}\left(\mathcal{D}\,|\,{\bm{w}}\right)\mathclose{}=\prod_{i=1}^{N}p\mathopen{}\left(y_{i}\,|\,f({\bm{x}}_{i};{\bm{w}})\right)\mathclose{}$ and a Gaussian variational distribution on model weights $q_{\phi}({\bm{w}})=\mathcal{N}\mathopen{}\left({\bm{w}}\,|\,{\bm{m}},{\bm{S}}\right)\mathclose{}$, the GFSVI objective function is
$\mathcal{L}(\phi)=\sum_{i=1}^{N}{\mathop{{}\mathbb{E}}}_{q_{\phi}({\bm{w}})}\mathopen{}\left[\log p\mathopen{}\left(y_{i}\,|\,f_{L}({\bm{x}}_{i};{\bm{w}})\right)\mathclose{}\right]\mathclose{}-D^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}^{F}\,\|\,\mathbb{P}^{F}\right)\mathclose{}$ (A.2)
where $\mathbb{Q}_{\phi}^{F}$ and $\mathbb{P}^{F}$ are the Gaussian measures induced by the linearized BNN and a Gaussian process prior respectively.
Expected log-likelihood
When considering a Gaussian likelihood, we use the closed form expression available due to the Gaussian variational measure over functions induced by the linearized BNN
${\mathop{{}\mathbb{E}}}_{q_{\phi}({\bm{w}})}\mathopen{}\left[\log\mathcal{N}\mathopen{}\left(y_{i}\,|\,f_{L}({\bm{x}}_{i};{\bm{w}}),\sigma_{y}^{2}\right)\mathclose{}\right]\mathclose{}=-\frac{1}{2}\log\mathopen{}\left(2\pi\sigma_{y}^{2}\right)\mathclose{}-\frac{\mathopen{}\left(y_{i}-f\mathopen{}\left({\bm{x}}_{i};{\bm{m}}\right)\mathclose{}\right)\mathclose{}^{2}+J\mathopen{}\left({\bm{x}}_{i};{\bm{m}}\right)\mathclose{}{\bm{S}}J\mathopen{}\left({\bm{x}}_{i};{\bm{m}}\right)\mathclose{}^{\top}}{2\sigma_{y}^{2}}.$ (A.3)
When considering a Categorical likelihood with $C$ different classes, we estimate the expected log-likelihood term using Monte-Carlo integration as
$\begin{split}{\mathop{{}\mathbb{E}}}_{q_{\phi}({\bm{w}})}\mathopen{}\left[\log\text{Cat}(y_{i}\,|\,\sigma(f_{L}({\bm{x}}_{i};{\bm{w}})))\right]\mathclose{}&=\frac{1}{K}\sum_{k=1}^{K}\sum_{c=1}^{C}\mathbb{I}[y_{i}=c]\Bigl{[}f^{c}_{L}({\bm{x}}_{i};{\bm{w}}^{(k)})\\ &\qquad-\log\left[\sum_{c^{\prime}=1}^{C}\exp\left(f^{c^{\prime}}_{L}({\bm{x}}_{i};{\bm{w}}^{(k)})\right)\right]\Bigr{]}\end{split}$ (A.4)
where ${\bm{w}}^{(k)}\sim q_{\phi}({\bm{w}})$ for $k=1,\dots,K$, $\mathbb{I}[\cdot]$ is the indicator function, $\sigma(\cdot)$ is the softmax function and $f^{c}_{L}(\,\cdot\,;{\bm{w}})$ is the logit for class $c$ obtained from $f_{L}$.
Regularized KL divergence
We estimate the regularized KL divergence using its consistent estimator (Eq. 2.9)
$\hat{D}^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}^{F}\,\|\,\mathbb{P}^{F}\right)\mathclose{}=\frac{1}{2}\mathopen{}\left(f({\bm{x}};{\bm{m}})-\mu({\bm{x}})\right)\mathclose{}^{\top}(K({\bm{x}},{\bm{x}})+\gamma M\mathbb{I}_{M})^{-1}\mathopen{}\left(f({\bm{x}};{\bm{m}})-\mu({\bm{x}})\right)\mathclose{}\\ +\frac{1}{2}\operatorname{Tr}\big{[}(K({\bm{x}},{\bm{x}})+\gamma M\mathbb{I}_{M})^{-1}(J({\bm{x}};{\bm{m}})SJ({\bm{x}};{\bm{m}})^{\top}+\gamma M\mathbb{I}_{M})-\mathbb{I}_{M}\big{]}\\ -\frac{1}{2}\log\det\big{[}(K({\bm{x}},{\bm{x}})+\gamma M\mathbb{I}_{M})^{-1}(J({\bm{x}};{\bm{m}})SJ({\bm{x}};{\bm{m}})^{\top}+\gamma M\mathbb{I}_{M})\big{]}$ (A.5)
[…]
A.3 Additional details on the experimental setup
A.3.1 Experiments on synthetic data
Regression
We consider the following generative model for the toy data
$y_{i}=\operatorname{sin}(2\pi x_{i})+\epsilon\quad\text{with}\quad\epsilon\sim\mathcal{N}\mathopen{}\left(0,\sigma_{n}^{2}\right)\mathclose{}$ (A.6)
and draw $x_{i}\sim\mathcal{U}([-1,-0.5]\cup[0.5,1])$. When not otherwise specified, we use $\sigma_{n}=0.1$. On the plots, the data points are shown as gray circles, inferred mean functions as red lines, their 2-standard-deviations interval around the mean in light green, and functions sampled from the approximate posterior as green lines. In general, we consider two hidden-layer BNNs with 30 neurons per layer and hyperbolic tangent activation (Tanh) functions. Specifically in figure 11, the small BNN has the same architecture as above while the large BNN has 100 neurons per layer. All the BNN baselines have the same architecture and fully-factorized Gaussian approximate posterior. The prior scale of TFSVI (Rudner et al., 2022b, ) is set to $\sigma_{p}=0.2$ and $\sigma_{p}=0.75$ for MFVI (Blundell et al.,, 2015) and Laplace (Immer et al.,, 2021). For the Gaussian process posterior baseline, we fit the prior parameters by maximizing the log-marginal likelihood (Williams and Rasmussen,, 2006). Apart from the cases where the parameters of the GP prior used for GFSVI (our method) and FVI (Sun et al.,, 2019) are explicitly stated, we consider a constant zero-mean function and find the parameters of the covariance function by maximizing the log-marginal likelihood from mini-batches (Chen et al.,, 2021). Except where otherwise stated, we estimate the functional KL divergences with 500 measurement points and use the regularized KL divergence with $\gamma=10^{-10}$.
Classification
We sample 100 data points perturbed by Gaussian noise with $\sigma_{n}=0.1$ from the two moons data (Pedregosa et al.,, 2011). On the plots, the data points are shown as red (class 0) and blue (class 1) dots. We plot the mean and 2-standard-deviations of the probability that ${\bm{x}}$ belongs to class 1 with respect to the posterior (i.e. $p(y=1\,|\,{\bm{w}}^{(k)},{\bm{x}})$) which we estimate from $K=100$ samples ${\bm{w}}^{(k)}\sim q_{\phi}({\bm{w}})$ for $k=1,\dots,K$. We consider two hidden-layer BNNs with 100 neurons per layer and hyperbolic tangent activation (Tanh) functions. All the BNN baselines have the same architecture and fully-factorized Gaussian approximate posterior. The prior scale of MFVI (Blundell et al.,, 2015) is set to $\sigma_{p}=0.8$ and $\sigma_{p}=1.0$ for TFSVI (Rudner et al., 2022b, ) and Laplace (Immer et al.,, 2021). For the Gaussian process posterior baseline, we approximate the intractable posterior using the Laplace approximation and find the prior parameters by maximizing the log-marginal likelihood (Williams and Rasmussen,, 2006). The GP prior for GFSVI (our method) and FVI (Sun et al.,, 2019) has a constant zero-mean function and we find the parameters of the covariance function by maximizing the log-marginal likelihood from mini-batches (Chen et al.,, 2021) using the method to transform classifications labels into regression targets from Milios et al., (2018). We estimate the functional KL divergences with 500 measurement points and use the regularized KL divergence with $\gamma=10^{-10}$.
A.3.2 Ocean current modeling experiment
We use $20$ two-dimensional velocity vectors from the GulfDrifters dataset (Lilly and Pérez-Brunius,, 2021) and we follow the setup by Shalashilin, (2024). We apply the Helmholtz decomposition to the neural network $f$ using the following parameterization
$f(\cdot,{\bm{w}})=\operatorname{grad}\Phi(\cdot,{\bm{w}}_{1})+\operatorname{rot}\Psi(\cdot,{\bm{w}}_{2})$ (A.7)
where ${\bm{w}}=\left\{{\bm{w}}_{1},{\bm{w}}_{2}\right\}$ and, $\Phi(\cdot,{\bm{w}}_{1})$ and $\Psi(\cdot,{\bm{w}}_{2})$ are two 2-layered fully-connected neural networks with hyperbolic tangent activation functions and $50$ hidden units per layer. GFSVI and TFSVI both use $160$ fixed context points. The prior scale of TFSVI is set to $\sigma_{p}=0.5$. We fit the neural networks on the entire dataset and average the scores with respect to five different random seeds.
A.3.3 Regression experiments with tabular data
Datasets and pre-processing
We evaluate the predictive performance of our model on regression datasets from the UCI repository (Dua and Graff,, 2017) described in table 6. These datasets are also considered in Sun et al., (2019); Wild et al., (2022) but we include two additional larger ones (Wave and Denmark). We perform 5-fold cross validation, leave out one fold for testing, consider 10% of the remaining 4 folds as validation data and the rest as training data. We report mean and standard-deviation of the average expected log-likelihood and average mean square error on the test fold. We also report the mean rank of the methods across all datasets by assigning rank 1 to the best scoring method as well as any method who’s error bars overlap with the highest score’s error bars, and recursively apply this procedure to the methods not having yet been assigned a rank. The expected log-likelihood is estimated by Monte Carlo integration when it is not available in closed form (MFVI, TFSVI and FVI) with 100 posterior samples. We preprocess the dataset by encoding categorical features as one-hot vectors and standardizing the features and labels.
Baseline specification
We compare our GFSVI method to two weight-space inference methods (mean-field variational inference (Blundell et al.,, 2015) and linearized Laplace (Immer et al.,, 2021)) and two function-space inference methods (FVI (Sun et al.,, 2019) and TFSVI (Rudner et al., 2022b, )). While FVI uses GP priors, TFSVI performs inference in function space but with the pushforward to function space of the variational distribution and prior on the weights. We compute the function-space (regularized) KL divergence using a set of 500 measurement points sampled from a uniform distribution for GFSVI and TFSVI, and 50 points drawn from a uniform distribution along with 450 samples from the training batch for FVI as specified in Sun et al., (2019). All the BNN baselines have the same architecture and fully-factorized Gaussian approximate posterior. We also provide results with a GP (Williams and Rasmussen,, 2006) when the size of the dataset allows it, and a sparse GP (Hensman et al.,, 2013). As we restrict our comparison to BNNs, we do not consider the GP and sparse GP as baselines but rather as gold-standards. All models have a Gaussian heteroskedastic noise model with a learned scale parameter. All the BNNs are fit using the Adam optimizer (Kingma and Ba,, 2017) using a mini-batch size of 2000 samples. We also perform early stopping when the validation loss stops decreasing.
Model selection
Hyper-parameter optimization is conducted using the Bayesian optimization tool provided by Wandb (Biewald,, 2020). BNN parameters are selected to maximize the average validation expected log-likelihood across the 5 cross-validation folds. We optimize over prior parameters (kernel and prior scale), learning-rate and activation function. We select priors for GFSVI, FVI, sparse GP and GP among the RBF, Matern-1/2, Matern-3/2, Matern-5/2, Linear and Rational Quadratic covariance functions. The GP prior parameters used with GFSVI and FVI are selected by maximizing the log-marginal likelihood from batches as proposed by Chen et al., (2021) and done in Sun et al., (2019). Hyper-parameters for GPs and sparse GPs (kernel parameters and learning-rate) are selected to maximize the mean log-marginal likelihood of the validation data across the 5 cross-validation folds.
Table 6: UCI regression dataset description
Dataset Boston Naval Power Protein Yacht Concrete Energy Kin8nm Wine Wave Denmark
Number samples 506 11 934 9 568 45 730 308 1 030 768 8 192 1 599 288 000 434 874
Number features 13 16 4 9 6 8 8 8 11 49 2
A.3.4 Classification experiments with image data
Datasets and pre-processing
We further evaluate the predictive performance of our model on classification tasks with the MNIST (LeCun et al.,, 2010) and Fashion MNIST (Xiao et al.,, 2017) image data sets. We fit the models on a random subset of 90% of the provided training split, consider the remaining 10% as validation data and evaluate on the provided test split. We repeat this procedure 5 times with different random seeds and report the mean and standard-deviation of the average expected log-likelihood, accuracy and expected calibration error (ECE) of the mean of the predictive distribution on the test set. The expected log-likelihood is estimated by Monte Carlo integration with 100 posterior samples when it is not available in closed form (MFVI, TFSVI and FVI). We estimate the mean of the predictive distribution to compute the accuracy and the ECE with 100 posterior samples. We preprocess the dataset by standardizing the images.
Baseline specification
We compare our GFSVI method to the same baselines as for the regression experiments (see A.3.3). All the BNN baselines have the same architecture and fully-factorized Gaussian approximate posterior. More specifically, we consider a CNN with three convolutional layers (with output channels 16, 32 and 64) before two fully connected layers (with output size 128 and 10). The convolutional layers use $3\times 3$ shaped kernels. Each pair of convolutional layers is interleaved with a max-pooling layer. We consider three different measurement point distributions $\rho$ to estimate the (regularized) KL divergence in GFSVI, FVI and TFSVI: random, random pixel and kmnist. The random measurement point distribution is sampled from by drawing 50% of the samples from the training data batch and 50% of the samples from a uniform distribution over $[p_{min},p_{max}]^{H\times W\times C}$, where $H$, $W$ and $C$ are respectively the height, width and number of channels of the images, and $p_{min}=v_{min}-0.5\times\Delta$ and $p_{max}=v_{max}+0.5\times\Delta$ where $\Delta=v_{max}-v_{min}$ is the difference between the minimal ($v_{min}$) and maximal ($v_{max}$) pixel values of the data set. The random pixel measurement point distribution is taken from Rudner et al., 2022b and is sampled from by randomly choosing each pixel value among the ones available from the training data batch at the same position in the $28\times 28$ pixel grid. Finally, the kmnist measurement point distribution is also taken from Rudner et al., 2022b and is drawn from by randomly sampling data points from the Kuzushiji-MNIST (KMNIST) dataset (Clanuwat et al.,, 2018). The KMNIST dataset is a collection of 70’000 gray-scale images of size $28\times 28$ which we preprocess by standardizing the images. We sample 10 measurement points when using random, 25 measurement points when using random pixel and 20 when using kmnist. All the BNNs are trained using the Adam optimizer (Kingma and Ba,, 2017) using a mini-batch size of 100. We also perform early stopping when the validation loss stops decreasing.
Model selection
Hyper-parameter optimization is conducted just like for the regression tasks (see A.3.3). The Gaussian process prior parameters used with GFSVI and FVI are selected by maximizing the log-marginal likelihood from batches (Chen et al.,, 2021) using the method to transform classifications labels into regression targets from Milios et al., (2018). We optimize the same hyper-parameters as for the regression experiments with the exception of the additional $\alpha_{\epsilon}$ parameter introduced by Milios et al., (2018) for the function-space VI methods with GP priors (FVI and GFSVI).
A.3.5 OOD detection
Tabular data with a Gaussian likelihood
Following the setup from Malinin et al., (2020) we take epistemic uncertainty to be the variance of the mean prediction with respect to samples from the posterior. We consider the test data to be in-distribution (ID) data and a subset of the song dataset (Bertin-Mahieux et al.,, 2011) of equal length and with an equal number of features as out-of-distribution (OOD) data. We use the same preprocessing as for regression as well as the same baselines with the same hyper-parameters (see section A.3.3). We first fit a model, then evaluate the extend by which the epistemic uncertainty under the model is predictive of the ID and OOD data using a single threshold obtained by a depth-1 decision tree fit to minimize the classification loss. We report the mean and standard error of the accuracy of the threshold to classify OOD from ID data based on epistemic uncertainty across the 5 folds of cross-validation. We also provide results obtained using a GP and sparse GP as gold standard.
Image data with a Categorical likelihood
Following the setup by Osawa et al., (2019), we take the epistemic uncertainty to be the entropy of the mean of the predictive distribution with respect to samples from the posterior. We evaluate models trained on MNIST using MNIST’s test split as ID data and a subset of the training set of Fashion MNIST as OOD data. Likewise, we evaluate models trained on Fashion MNIST using Fashion MNIST’s test split as ID data and a subset of the training set of MNIST as OOD data. We use the same preprocessing as for classification, as well as the same baselines with the same hyper-parameters (see section A.3.4). We first fit a model, then evaluate the extend by which the epistemic uncertainty under the model is predictive of the ID and OOD data using a single threshold obtained by a depth-1 decision tree fit to minimize the classification loss. We estimate mean of the predictive distribution by Monte-Carlo integration using 100 posterior samples. We report the mean and standard error of the accuracy of the threshold to classify OOD from ID data based on epistemic uncertainty for the 5 models trained on different random seeds (see section A.3.4).
A.3.6 Variational measure evaluation
We evaluate our inference method by comparing the samples drawn from the exact posterior over functions with the approximate posterior obtained with our method (GFSVI). We follow the setup by Wilson et al., (2022) and we compute the average Wasserstein-2 metric between 1000 samples drawn from a GP posterior with a RBF kernel evaluated at the test points, and samples from the approximate posterior of GFSVI, sparse GP and FVI evaluated at the same points and with the same prior. We consider the Boston, Concrete, Energy, Wine and Yacht datasets for which the exact GP posterior can be computed and use the same preprocessing as for regression (see section A.3.3). We report the mean and standard error of the average Wasserstein-2 metric across the 5 folds of cross-validation. The Wasserstein-2 metric is computed using the Python Optimal Transport library (Flamary et al.,, 2021).
Baseline specification
FVI and GFSVI have the same two hidden layer neural network architecture with 100 neurons each and hyperbolic tangent activation. These models are fit with the same learning rate and set of 500 measurement points jointly sampled from a uniform distribution over the feature-space and mini-batch of training samples. We use $\gamma=10^{-15}$ for the regularized KL divergence. We further consider a sparse GP with 100 inducing points.
A.3.7 Software
We use the JAX (Bradbury et al.,, 2018) and DM-Haiku (Hennigan et al.,, 2020) Python libraries to implement our Bayesian neural networks. MFVI, linearized Laplace and TFSVI were implemented based on the information in the papers, and code for FVI was adapted to the JAX library from the implementation provided by the authors. We further use the GPJAX Python library for experiments involving Gaussian processes (Pinder and Dodd,, 2022).
A.3.8 Hardware
All models were fit using a single NVIDIA RTX 2080Ti GPU with 11GB of memory.
A.4 Additional experimental results
In this section, we present additional figures for our qualitative uncertainty evaluation as well as further experimental results on regression, out-of-distribution detection and robustness under input distribution shift tasks. We also provide plots illustrating the eigenvalue decay of different kernels, and figures showing the influence of $\gamma$ in the regularized KL divergence.
A.4.1 Qualitative uncertainty evaluation
Figure 4: Our method (GFSVI) effectively regularizes functions generated by the Bayesian neural network (BNN) both in settings where the generative process is very noisy ($\sigma_{n}=1$) or not ($\sigma_{n}=0.1$).
Regression
We further find that our method (GFSVI) provides strong regularization when the data generative process is noisy (see figure 4) and is more robust than FVI to situations where ones computational budget constrains the number of measurement points $M$ to be small (figure 8). In contrast to FVI, GFSVI accurately approximates the exact GP posterior under rough (Matern-1/2) GP priors effectively incorporating prior knowledge defined by the GP prior to the inference process (see figure 5). Likewise, GFSVI adapts to the variability of the functions specified by the kernel (see figure 6). We also find that GFSVI requires a larger number of measurement points to capture the behavior of a rougher prior (see figure 7).
Figure 5: Our method (GFSVI) with a Matern-1/2 Gaussian process (GP) prior accurately approximates the exact GP posterior unlike the function-space prior baseline (FVI). Weight-space prior baselines do not provide a straight-forward mechanism to incorporate prior assumptions regarding the functions generated by BNNs and underestimate the epistemic uncertainty (MFVI, Laplace). The lower row is identical to the one in figure 2 in the main text and is reproduced here to make comparison easier.
Figure 6: Our method (GFSVI) allows to incorporate prior beliefs in terms of function variability using the characteristic length-scale parameter of the Gaussian process (GP) prior. GFSVI was fit using a GP prior with RBF covariance function.
Figure 7: Our method (GFSVI) captures the smooth behavior of a Gaussian process (GP) prior with RBF covariance function even if the number of measurement points is small (M=10). However, in that setting GFSVI fails to reproduce the rough effect of a GP prior with a Matern-1/2 covariance function, and requires a larger amount of measurement points to do so (M=100).
Figure 8: Our method (GFSVI) already provides a reasonable approximation to the exact posterior with small numbers of measurement points (M=10) while function-space baseline FVI requires many more (M=100).
Classification
We find that GFSVI better captures the beliefs induced by the smooth RBF and rough Matern-1/2 Gaussian process priors compared to FVI (see Figures 9 and 10). Moreover, GFVSI both accurately fits the training data and shows greater uncertainty outside of its support relative to BNNs baselines with weight-space and function-space priors. Unlike for the toy data regression experiments where the GP posterior was the ground truth, the Laplace (approximate) GP posterior in Figures 9 and 10 only represents a possible approximation to the now in-tractable posterior (due to the softmax inverse link function). Thus the GP should not be considered as the ground truth nor as the optimal approximation in the classification setting, but is nevertheless useful to give a idea of the level of uncertainty a BNN with a GP prior should provide outside of the support of the data.
Figure 9: Our method (GFSVI) with a RBF Gaussian process (GP) prior accurately captures the smooth decision boundary induced by the prior and shows high uncertainty outside of the data support. Weight-space baselines do not provide a straight-forward mechanism to incorporate prior assumptions regarding the functions generated by BNNs and underestimate the epistemic uncertainty (TFSVI, Laplace) or underfit the data (MFVI).
Figure 10: Our method (GFSVI) with a Matern-1/2 Gaussian process (GP) prior accurately captures the rough decision boundary unlike the function-space baseline (FVI). Weight-space baselines do not provide a straight-forward mechanism to incorporate prior assumptions regarding the functions generated by BNNs and underestimate the epistemic uncertainty (TFSVI, Laplace) or underfit the data (MFVI).
Inductive biases
figure 11 compares GFSVI to the exact posterior across two different priors and three model architectures (details in A.3.1). We find that the BNN’s ability to incorporate the beliefs introduced by the GP prior depends on its size and activation function. When using piece-wise linear activations (ReLU), small models are prone to underfitting for smooth priors (RBF), and to collapsing uncertainty for rough priors (Matern-1/2). By contrast, when using smooth activations (Tanh), smaller models suffice, and they are compatible with most standard GP priors (the results shown in figure 11 extend to RBF, Matern family, and Rational Quadratic in our experiments). We also analyzed how the number $M$ of measurement points affects performance. figures 7 and 14 show that capturing the properties of rough GP priors and estimating $D_{\text{KL}}^{\gamma}$ with these priors requires larger $M$.
Figure 11: Our method requires that the Bayesian neural network (BNN) and Gaussian process (GP) prior share similar inductive biases to provide an accurate approximation to the exact posterior.
A.4.2 Regression on tabular data
Table 7: Test mean square error (MSE) of evaluated methods on regression datasets. We find that GFSVI (ours) also performs competitively in terms of MSE compared to baselines and obtains the best mean rank, matching best the performing methods on nearly all datasets.
Dataset Function-space priors Weight-space priors Gaussian Processes (Gold Standards)
GFSVI (ours) FVI TFSVI MFVI VIP Laplace GWI Sparse GP GP
Boston 0.123 $\pm$ 0.021 0.136 $\pm$ 0.022 0.995 $\pm$ 0.092 0.532 $\pm$ 0.072 0.201 $\pm$ 0.056 0.203 $\pm$ 0.047 0.273 $\pm$ 0.069 0.122 $\pm$ 0.014 0.115 $\pm$ 0.020
Concrete 0.114 $\pm$ 0.008 0.116 $\pm$ 0.004 0.389 $\pm$ 0.015 0.698 $\pm$ 0.046 0.109 $\pm$ 0.008 0.116 $\pm$ 0.007 0.145 $\pm$ 0.017 0.399 $\pm$ 0.020 0.116 $\pm$ 0.007
Energy 0.003 $\pm$ 0.000 0.003 $\pm$ 0.000 0.003 $\pm$ 0.000 0.152 $\pm$ 0.024 0.043 $\pm$ 0.036 0.002 $\pm$ 0.000 0.003 $\pm$ 0.001 0.087 $\pm$ 0.005 0.087 $\pm$ 0.004
Kin8nm 0.071 $\pm$ 0.001 0.075 $\pm$ 0.003 0.073 $\pm$ 0.001 0.290 $\pm$ 0.111 0.068 $\pm$ 0.002 0.083 $\pm$ 0.001 0.071 $\pm$ 0.001 0.088 $\pm$ 0.002 (infeasible)
Naval $0$$\pm$ 0.000 0.001 $\pm$ 0.001 $0$$\pm$ 0.000 0.007 $\pm$ 0.003 0.002 $\pm$ 0.000 $0$$\pm$ 0.000 0.197 $\pm$ 0.174 $0$$\pm$ 0.000 (infeasible)
Power 0.052 $\pm$ 0.001 0.054 $\pm$ 0.002 0.054 $\pm$ 0.001 0.058 $\pm$ 0.002 0.054 $\pm$ 0.002 0.054 $\pm$ 0.002 0.052 $\pm$ 0.001 0.071 $\pm$ 0.001 (infeasible)
Protein 0.459 $\pm$ 0.005 0.466 $\pm$ 0.004 0.429 $\pm$ 0.004 0.537 $\pm$ 0.008 0.421 $\pm$ 0.005 0.446 $\pm$ 0.006 0.425 $\pm$ 0.003 0.408 $\pm$ 0.002 (infeasible)
Wine 0.652 $\pm$ 0.022 0.663 $\pm$ 0.009 1.297 $\pm$ 0.093 0.655 $\pm$ 0.023 0.627 $\pm$ 0.013 0.637 $\pm$ 0.031 0.682 $\pm$ 0.048 0.607 $\pm$ 0.033 0.585 $\pm$ 0.032
Yacht 0.003 $\pm$ 0.001 0.004 $\pm$ 0.001 0.221 $\pm$ 0.037 0.682 $\pm$ 0.140 0.004 $\pm$ 0.001 0.002 $\pm$ 0.001 0.008 $\pm$ 0.003 0.399 $\pm$ 0.064 0.355 $\pm$ 0.030
Wave $0$$\pm$ 0.000 $0$$\pm$ 0.000 $0$$\pm$ 0.000 $0$$\pm$ 0.000 $0$$\pm$ 0.000 $0$$\pm$ 0.000 0.001 $\pm$ 0.001 $0$$\pm$ 0.000 (infeasible)
Denmark 0.155 $\pm$ 0.004 0.287 $\pm$ 0.003 0.163 $\pm$ 0.004 0.225 $\pm$ 0.003 0.189 $\pm$ 0.008 0.194 $\pm$ 0.003 0.197 $\pm$ 0.004 0.260 $\pm$ 0.001 (infeasible)
Mean rank 1.364 2.000 2.182 3.182 1.636 1.727 - - -
We present additional regression results reporting the mean square error (MSE) of evaluated methods across the considered baselines, see table 7. We find that GFSVI also performs competitively in terms of MSE compared to baselines and obtains the best mean rank, matching best the performing methods on nearly all datasets. In particular, we find that using GP priors in the linearized BNN setup with GFSVI yields improvements over the weight-space priors used in TFSVI and that GFSVI performs slightly better than FVI. Function-space VI methods (TFSVI, GFSVI, FVI) significantly improves over weight-space VI mostly performing similarly to the linearized Laplace approximation. Further improvement over baselines are obtained when considering GP priors with GFSVI and FVI. Finally, GFSVI compares favorably to the GP and sparse GP.
A.4.3 Out-of-distribution detection with image data
We here show an additional plot from our out-of-distribution detection experiment with image data (details in A.3.5). figure 12 shows the (normalized) histograms of the entropy of the mean prediction produced by each model on the in-distribution (blue) and out-of-distribution (red) data sets considered in our OOD detection experiment. Methods which estimate the (regularized) KL-divergence in function-space (GFSVI, FVI and TFSVI) use the kmnist measurement distribution. We find that the entropy produced by GFSVI on in-distribution data highly peaks around 0 while the entropy produced from out-of-distribution data strongly concentrates around its maximum $\ln(10)$. GFSVI best partitions ID and OOD data based on predictive entropy improving over the function-space prior (FVI) and weight-space prior (TFSVI, MFVI, Laplace) BNN baselines (see table 2).
Figure 12: Histograms of the entropy of the mean predictive distribution produced by evaluated methods in the out-of-distribution detection with image data experiment. GFSVI (ours) best partitions in-distribution and out-of-distribution data based on the entropy of its mean predictive distribution.
A.4.4 Influence of measurement point distribution for image data.
We present additional results evaluating the influence of the measurement point distribution $\rho$ on the the performance of function-space inference methods when using high-dimensional image data. The measurement point distribution are described in section A.3.4. Just like in Rudner et al., 2022b , we find that the choice of measurement point distribution may highly influence the OOD detection accuracy. While the expected log-likelihood, accuracy and expected calibration error (ECE) of a model generally remains comparable across measurement point distributions, the OOD accuracy of GFSVI is greatly improved by using samples from kmnist to evaluate the (regularized) KL divergence. The measurement point distribution determines where the BNN is regularized and thus should be carefully selected especially for high dimensional data.
Table 8: Influence of the measurement point distribution $\rho$ on expected log-likelihood (log-like.), accuracy (acc.), expected calibration error (ECE) and out-of-distribution detection accuracy (OOD acc.). $\rho$ determines where the BNN will be regularized and strongly influences the out-of-distribution performance of the BNN.
Data Metric GFSVI FVI TFSVI
Random Random Pixel KMNIST Random Random Pixel KMNIST Random Random Pixel KMNIST
MNIST Log-like. ($\uparrow$) -0.033 $\pm$ $0$0.034 $\pm$ $0$0.041 $\pm$ $0$0.145 $\pm$ 0.005 -0.038 $\pm$ $0$0.238 $\pm$ 0.006 -0.047 $\pm$ 0.003 -0.032 $\pm$ 0.001 -0.041 $\pm$ 0.001
Acc. ($\uparrow$) -0.992 $\pm$ $0$0.989 $\pm$ $0$0.991 $\pm$ $0$0.976 $\pm$ 0.001 -0.988 $\pm$ $0$0.943 $\pm$ 0.001 -0.989 $\pm$ $0$0.989 $\pm$ $0$0.989 $\pm$ 0.000
ECE ($\downarrow$) -0.002 $\pm$ $0$0.004 $\pm$ $0$0.006 $\pm$ $0$0.064 $\pm$ 0.001 -0.003 $\pm$ $0$0.073 $\pm$ 0.003 -0.007 $\pm$ $0$0.003 $\pm$ $0$0.006 $\pm$ 0.000
OOD acc. ($\uparrow$) -0.921 $\pm$ 0.008 -0.868 $\pm$ 0.010 -0.980 $\pm$ 0.004 -0.894 $\pm$ 0.010 -0.863 $\pm$ 0.003 -0.891 $\pm$ 0.006 -0.887 $\pm$ 0.011 -0.861 $\pm$ 0.008 -0.893 $\pm$ 0.005
FMNIST Log-like. ($\uparrow$) -0.260 $\pm$ 0.003 -0.258 $\pm$ 0.002 -0.294 $\pm$ 0.006 -0.300 $\pm$ 0.002 -0.293 $\pm$ 0.003 -0.311 $\pm$ 0.005 -0.261 $\pm$ 0.001 -0.258 $\pm$ 0.001 -0.261 $\pm$ 0.002
Acc. ($\uparrow$) -0.910 $\pm$ 0.001 -0.908 $\pm$ 0.001 -0.909 $\pm$ 0.001 -0.910 $\pm$ 0.002 -0.900 $\pm$ 0.001 -0.906 $\pm$ 0.002 -0.909 $\pm$ 0.001 -0.908 $\pm$ 0.001 -0.907 $\pm$ 0.001
ECE ($\downarrow$) -0.020 $\pm$ 0.003 -0.022 $\pm$ 0.001 -0.042 $\pm$ 0.002 -0.027 $\pm$ 0.005 -0.018 $\pm$ 0.002 -0.024 $\pm$ 0.002 -0.022 $\pm$ 0.002 -0.018 $\pm$ 0.001 -0.021 $\pm$ 0.002
OOD acc. ($\uparrow$) -0.853 $\pm$ 0.005 -0.867 $\pm$ 0.005 -0.997 $\pm$ 0.001 -0.925 $\pm$ 0.005 -0.842 $\pm$ 0.006 -0.975 $\pm$ 0.002 -0.802 $\pm$ 0.006 -0.800 $\pm$ 0.007 -0.779 $\pm$ 0.010
A.4.5 Input distribution shift with rotated image data
We here provide an experiment evaluating our method’s (GFSVI) robustness in detecting input distribution shift. We expect the predictive uncertainty of a well-calibrated Bayesian model to be low for in-distribution data and to gradually increase as the input distribution shifts further away from the training data distribution. To test this property, we follow the setup by Sensoy et al., (2018); Rudner et al., 2022b and assume like the related work that increasing the rotation angle of images gradually increases the level of input "distribution shift". We report the mean and standard-deviation of the average mean predictive entropy of models fit on MNIST (LeCun et al.,, 2010) and Fashion MNIST (Xiao et al.,, 2017) for increasingly large angles of rotation of their respective test data partition. We find that GFSVI is confident (low predictive entropy) for images with small rotation angles, and that its predictive entropy increases with the angle. GFSVI therefore exhibits the expected behavior of a well-calibrated Bayesian model. We note that FVI, Laplace and MFVI tend to be under-confident (high predictive entropy) for small rotation angles, which might be a symptom of underfitting further supported by the results in table 2. Also, with the exception of TFSVI, the predictive entropy of baselines across different rotation angles is generally higher than the one produced by GFSVI.
Figure 13: Average predictive entropy of models trained on MNIST and Fashion MNIST and evaluated for different rotation angles of their respective test data partitions. We see that our method (GFSVI) exhibits the behavior of a well-calibrated Bayesian model.
A.4.6 Additional plots for kernel eigenvalue decay
figure 14 shows a plot demonstrating the decay rate of the eigenvalues of RBF and Matern-1/2 kernels evaluated at points sampled uniformly over $\mathcal{X}$ The rate of decay of covariance operator’s eigenvalues gives important information about the smoothness of stationary kernels (Williams and Rasmussen,, 2006) and that increased smoothness of the kernel leads to faster decay of eigenvalues Santin and Schaback, (2016). For instance, RBF covariance operator eigenvalues decay at near exponential rate independent of the underlying measure (Belkin,, 2018) and Matern kernels eigenvalues decay polynomialy (Chen et al.,, 2021). We find that the kernel evaluated at points sampled from a uniform distribution over $\mathcal{X}$ share this same behavior (see figure 14).
(a) $D=1$
(b) $D=100$
Figure 14: Mean eigenvalues of the Gram matrix obtained for different kernels and for varying length-scales over $10$ draws from a uniform distribution on $[-2,2]^{D}$. The mean eigenvalues are arranged in increasing order. The eigenvalues of the Gram matrix associated with the smooth RBF kernel decays much faster than those of the Matern-1/2. Furthermore, the eigenvalues decay at a slower rate in high dimensions (D=100).
A.4.7 Additional plots for choosing $\gamma$ in $D_{\text{KL}}^{\gamma}$
The $\gamma$ parameter controls the magnitude of the regularized KL divergence (see figure 17) and adjusts the relative weight of the regularized KL divergence and expected log-likelihood term in the training objective (see figure 15). Furthermore, $\gamma$ also acts as "jitter" preventing numerical errors. We recommend choosing $\gamma$ large enough to avoid numerical errors while remaining small enough to provide strong regularization.
Figure 15: The $\gamma$ parameter of the regularized KL divergence controls the magnitude of the regularizer in the objective and should be small enough to provide strong regularization.
Figure 16: The BNN’s covariance adaptation to the prior’s covariance rank depends on its activation function. BNNs fit with a RBF prior (full) show lower rank than with a Matern-1/2 (dotted).
Figure 17: $\gamma$ explicitly controls the magnitude of the regularized KL-divergence $D_{\text{KL}}^{\gamma}$. Rougher priors (Matern-1/2) require more measurement points to accurately estimate $D_{\text{KL}}^{\gamma}$ than smooth priors (RBF).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The well-definiteness of our variational objective follows from the fact that the regularized KL divergence is finite for any pair of Gaussian measures Quang, (2019). tables 1, 2, 4 and 5 as well as Figures 2 to 10 support our claim about the performance of our method.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The inductive biases paragraph in section 4.1 discusses limitations.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Our only theoretical result is proved in Appendix and is the direct application of the Feldman-Hàjek theorem. We specify all assumptions.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The experimental setup is rigorously described in Appendix A.3.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The pseudo code, experiment descriptions in appendix as well as the code provides all the relevant information to reproduce experiments.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The experimental setup is rigorously described in Appendix A.3.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report mean and standard-error of our method across five cross-validation folds (regression) or random partitions of the data (classification).
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We specify computational resources in Appendix A.3.7.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: To the best of our knowledge, we have fully conformed with the code of ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [No]
  49. Justification: This paper presents work whose goal is to advance the field of Bayesian deep learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The methods developed in the paper have no risk for misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Datasets used for the experiments are cited in the main text and their license and terms of use are properly respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: the paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: that the paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v2

@@
 Burt et al., (2020) point out an even more severe issue of function-space VI in BNNs: $D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}$ is in fact infinite in many relevant cases, in particular for non-degenerate GP-priors with BNNs.
-Thus, approximations in these cases are futile.
 The proof in Burt et al., (2020) is somewhat involved, but the fundamental reason for $D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}=\infty$ is that $\mathbb{Q}_{\phi}$ has support on a finite-dimensional submanifold of the infinite-dimensional function space, while the measure $\mathbb{P}$ induced by a (non-degenerate) GP prior has support on the entire function space.
@@
 Unlike TFSVI, we find that GFSVI accurately captures ocean current dynamics.
-We test the extend to which GFSVI can incorporate prior knowledge specified via a GP prior by considering the problem of modeling ocean currents in the Gulf of Mexico.
-We are given $20$ two-dimensional velocity vectors collected from ocean drifter buoys, and we are interested in estimating ocean currents.
-We use the GulfDrifters dataset (Lilly and Pérez-Brunius,, 2021) and we follow the setup by Shalashilin, (2024).
+We consider the problem of modeling ocean currents in the Gulf of Mexico to measure the extend to which GFSVI can incorporate knowledge specified via a GP prior on real-world data.
+Specifically, we follow the setup by Shalashilin, (2024) and Cinquin et al., (2024), and use the GulfDrifters dataset (Lilly and Pérez-Brunius,, 2021) to estimate ocean currents from $20$ two-dimensional velocity vectors collected from drifter buoys.
+We embed physical properties of fluid motions into the GP prior and to the neural networks by applying the Helmholtz decomposition (Berlinghieri et al.,, 2023; Cinquin et al.,, 2024).
 We compare our GFSVI to a GP and to TFSVI.
-We incorporate known physical properties of fluid motions by applying the Helmholtz decomposition to the GP prior (Berlinghieri et al.,, 2023) as well as to the neural networks.
 More details can be found in section A.3.2.
@@
 The authors extend their gratitude to the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Tristan Cinquin.
-Finally, we thank Vincent Fortuin and Marvin Pförtner for useful discussions and feedback.
+Finally, we thank Marvin Pförtner for very useful discussions on the theory and experiments, and Vincent Fortuin for his feedback.
 Appendix A Appendix
@@
 A.3.2 Ocean current modeling experiment
-We use $20$ two-dimensional velocity vectors from the GulfDrifters dataset (Lilly and Pérez-Brunius,, 2021) and we follow the setup by Shalashilin, (2024).
-We apply the Helmholtz decomposition to the neural network $f$ using the following parameterization
+Following Cinquin et al., (2024), we apply the Helmholtz decomposition to the neural network $f$ as
 $f(\cdot,{\bm{w}})=\operatorname{grad}\Phi(\cdot,{\bm{w}}_{1})+\operatorname{rot}\Psi(\cdot,{\bm{w}}_{2})$ (A.7)
-where ${\bm{w}}=\left\{{\bm{w}}_{1},{\bm{w}}_{2}\right\}$ and, $\Phi(\cdot,{\bm{w}}_{1})$ and $\Psi(\cdot,{\bm{w}}_{2})$ are two 2-layered fully-connected neural networks with hyperbolic tangent activation functions and $50$ hidden units per layer.
+where ${\bm{w}}=\left\{{\bm{w}}_{1},{\bm{w}}_{2}\right\}$ and, $\Phi(\cdot,{\bm{w}}_{1})$ and $\Psi(\cdot,{\bm{w}}_{2})$ are 2-layer fully-connected neural networks with $50$ hidden units per layer and hyperbolic tangent activation functions.
 GFSVI and TFSVI both use $160$ fixed context points.
@@
 Rougher priors (Matern-1/2) require more measurement points to accurately estimate $D_{\text{KL}}^{\gamma}$ than smooth priors (RBF).
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4.
-Justification: The well-definiteness of our variational objective follows from the fact that the regularized KL divergence is finite for any pair of Gaussian measures Quang, (2019).
-tables 1, 2, 4 and 5 as well as Figures 2 to 10 support our claim about the performance of our method.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: The inductive biases paragraph in section 4.1 discusses limitations.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory Assumptions and Proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [Yes]
-  14. Justification: Our only theoretical result is proved in Appendix and is the direct application of the Feldman-Hàjek theorem. We specify all assumptions.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental Result Reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: The experimental setup is rigorously described in Appendix A.3.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: The pseudo code, experiment descriptions in appendix as well as the code provides all the relevant information to reproduce experiments.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental Setting/Details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: The experimental setup is rigorously described in Appendix A.3.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment Statistical Significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: We report mean and standard-error of our method across five cross-validation folds (regression) or random partitions of the data (classification).
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments Compute Resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: We specify computational resources in Appendix A.3.7.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code Of Ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: To the best of our knowledge, we have fully conformed with the code of ethics.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader Impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [No]
-  49.
-Justification: This paper presents work whose goal is to advance the field of Bayesian deep learning.
-There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: The methods developed in the paper have no risk for misuse.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: Datasets used for the experiments are cited in the main text and their license and terms of use are properly respected.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New Assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [N/A]
-  64. Justification: The paper does not release new assets.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and Research with Human Subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: the paper does not involve crowdsourcing nor research with human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: that the paper does not involve crowdsourcing nor research with human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v3

@@
-Regularized KL-Divergence for Well-Defined Function-Space Variational Inference in Bayesian Neural Networks
+Well-Defined Function-Space Variational Inference in Bayesian Neural Networks via Regularized KL-Divergence
 Abstract
-Bayesian neural networks (BNN) promise to combine the predictive performance of neural networks with principled uncertainty modeling important for safety-critical systems and decision making.
-However, posterior uncertainty estimates depend on the choice of prior, and finding informative priors in weight-space has proven difficult.
-This has motivated variational inference (VI) methods that pose priors directly on the function generated by the BNN rather than on weights.
-In this paper, we address a fundamental issue with such function-space VI approaches pointed out by Burt et al., (2020), who showed that the objective function (ELBO) is negative infinite for most priors of interest.
-Our solution builds on generalized VI (Knoblauch et al.,, 2019) with the regularized KL divergence (Quang,, 2019) and is, to the best of our knowledge, the first well-defined variational objective for function-space inference in BNNs with Gaussian process (GP) priors.
-Experiments show that our method incorporates the properties specified by the GP prior on synthetic and small real-world data sets, and provides competitive uncertainty estimates for regression, classification and out-of-distribution detection compared to BNN baselines with both function and weight-space priors.
+Bayesian neural networks (BNN) promise to combine the predictive performance of neural networks with principled uncertainty modeling crucial for safety-critical systems and decision making.
+However, posterior uncertainties depend on the choice of prior, and finding informative priors in weight-space has proven difficult.
+This has motivated variational inference (VI) methods that pose priors directly on the function represented by the BNN rather than on weights.
+In this paper, we address a fundamental issue with such function-space VI approaches pointed out by Wild et al.
+[2022a], who showed that the objective function (ELBO) is negative infinite for most priors of interest.
+Our solution builds on generalized VI with the regularized KL divergence and is, to the best of our knowledge, the first well-defined variational objective for inference in BNNs with Gaussian process (GP) priors.
+Experiments show that our method successfully incorporates the properties specified by the GP prior, and that it provides competitive uncertainty estimates for regression, classification and out-of-distribution detection compared to BNN baselines with both function and weight-space priors.
 1 Introduction
-Neural networks have shown impressive results in many fields but fail to provide well-calibrated uncertainty estimates, which are essential in applications associated with risk, such as healthcare (Kompa et al.,, 2021; Abdullah et al.,, 2022) or finance (Bew et al.,, 2019; Wong,, 2023).
-Bayesian neural networks (BNNs) offer to combine the scalability and predictive performance of neural networks with principled uncertainty modeling by explicitly capturing epistemic uncertainty, i.e., uncertainty resulting from finite training data.
-While the choice of prior in the Bayesian framework strongly affects the uncertainty estimates later obtained from the posterior, specifying informative priors on BNN weights has proven difficult and is hypothesized to have limited their practical applicability (Knoblauch et al.,, 2019; Cinquin et al.,, 2021; Tran et al.,, 2022).
-For instance, the default isotropic Gaussian prior, which is often chosen for tractability rather than for the beliefs it carries (Knoblauch et al.,, 2019), is known to have pathological behavior in some cases (Tran et al.,, 2022).
+Figure 1: Inference with our GFSVI on synthetic data (gray circles) with Gaussian process priors encoding different properties such as smoothness (increasing from Matérn-1/2 to RBF) and periodicity (last panel).
+Neural networks have shown impressive results in many fields but fail to provide well-calibrated uncertainty estimates, which are essential in applications associated with risk, such as healthcare [Abdullah et al., 2022] or finance [Bew et al., 2019].
+Bayesian neural networks (BNNs) offer to combine the scalability and predictive performance of neural networks with principled uncertainty modeling by explicitly capturing epistemic uncertainty, which results from finite training data.
+While the choice of prior strongly affects posterior uncertainties, specifying informative priors on BNN weights has proven difficult and is hypothesized to have limited their practical applicability [Knoblauch et al., 2022, Tran et al., 2022].
+For instance, the default isotropic Gaussian prior, which is often chosen for tractability rather than for the beliefs it carries [Knoblauch et al., 2022], is known to have pathological behavior in some cases [Cinquin et al., 2021, Tran et al., 2022].
 A promising approach to solve this issue is to place priors directly on the function represented by the BNN instead of the weights.
-While being technically more challenging, function-space priors allow incorporating interpretable knowledge into the inference, for instance enabling the use of the extensive Gaussian Process (GP) literature to improve prior selection and design (Williams and Rasmussen,, 2006).
-A recent line of work has focused on using function-space priors in BNNs with variational inference (VI) (Sun et al.,, 2019).
-VI is an appealing method because of its successful application to BNNs, its flexibility in terms of approximate posterior parameterization, and its scalability to large datasets and models (Hoffman et al.,, 2013; Blundell et al.,, 2015; Tomczak et al.,, 2020).
-Unfortunately, in the context of BNNs with function-space priors, the Kullbach-Leibler (KL) divergence term in the VI objective (ELBO) involves two intractabilities: (i) a supremum over infinitely many subsets and (ii) access to the density of the distribution of function represented by the BNN, which has no closed-form expression.
-Sun et al., (2019) propose to address problem (i) by approximating the supremum in the KL divergence by an expectation, and problem (ii) by using implicit score function estimators (which make this method difficult to use in practice (Ma and Hernández-Lobato,, 2021)).
+Function-space priors allow incorporating interpretable knowledge, for instance using the Gaussian Process (GP) literature to improve prior design and selection [Williams and Rasmussen, 2006].
+A recent line of work has focused on using function-space priors in BNNs with variational inference (VI) [Sun et al., 2019].
+VI is appealing because of its successful application to BNNs, its flexibility in terms of approximate posterior parameterization, and its scalability to large datasets and models [Hoffman et al., 2013, Blundell et al., 2015].
+Unfortunately, for BNNs with function-space priors, the Kullbach-Leibler (KL) divergence term in the VI objective (ELBO) involves two intractabilities: (i) a supremum over infinitely many subsets and (ii) access to the density of the distribution of the BNN’s function, which has no closed-form expression.
+Sun et al.
+[2019] propose to address problem (i) by approximating the supremum in the KL divergence by an expectation, and problem (ii) by using implicit score function estimators (which make this method difficult to use in practice [Ma and Hernández-Lobato, 2021]).
 However, the problem is actually more severe.
-Not only is the KL divergence intractable, it is infinite in many cases of interest (Burt et al.,, 2020), such as when the prior is a non-degenerate GP or a BNN with a different architecture.
+Not only is the KL divergence intractable, it is infinite in most cases of interest [Wild et al., 2022a], such as when the prior is a non-degenerate GP or a BNN with a different architecture.
 Thus, in these (and many more) situations, the KL divergence cannot even be approximated.
-As a consequence, more recent work abandons using BNNs and instead uses deterministic neural networks to parameterize basis functions (Ma and Hernández-Lobato,, 2021) or a GP mean (Wild et al.,, 2022).
-The only prior work (Rudner et al., 2022b, ) on function-space VI in BNNs that overcomes the issue pointed out by Burt et al., (2020) does so by deliberately limiting itself to cases where the KL divergence is known to be finite (by defining the prior as the pushforward of a weight-space distribution).
-Therefore, the method by Rudner et al., 2022b suffers from the same issues regarding prior specification as any other weight-space inference method for BNNs.
-In this paper, we address the argument by Burt et al., (2020) that VI does not provide a valid objective for inference in BNNs with genuine function-space priors, and we propose to apply the framework of generalized variational inference (Knoblauch et al.,, 2019).
-We present a simple method for function-space inference with GP priors that builds on the so-called regularized KL divergence (Quang,, 2019), which generalizes the conventional KL divergence and is finite for any pair of Gaussian measures.
-We obtain a Gaussian measure for the variational posterior by considering the linearized BNN from Rudner et al., 2022b , and we are free to choose a prior from a large set of GPs which have an associated Gaussian measure on the considered function space.
+As a consequence, more recent work abandons using BNNs and instead uses deterministic neural networks to parameterize basis functions [Ma and Hernández-Lobato, 2021] or a GP mean [Wild et al., 2022b].
+The only prior work [Rudner et al., 2022b] that overcomes the issue pointed out by Wild et al.
+[2022a] does so by deliberately limiting itself to cases where the KL divergence is known to be finite (by defining the prior as the pushforward of a weight-space distribution).
+Therefore, the method by Rudner et al.
+[2022b] suffers from the same issues regarding prior specification as other weight-space inference method.
+In this paper, we address the argument by Wild et al.
+[2022a] that VI does not provide a valid objective for inference in BNNs with genuine function-space priors, and we propose to apply the framework of generalized VI [Knoblauch et al., 2022].
+We present a simple method for function-space inference with GP priors that builds on the regularized KL divergence [Quang, 2019], which generalizes the conventional KL divergence and is finite for any pair of Gaussian measures.
+We obtain a Gaussian measure for the variational posterior by considering the linearized BNN from Rudner et al.
+[2022b], and we are free to choose a function-space prior from a large set of GPs which have an associated Gaussian measure on the considered function space.
 While the regularized KL divergence is still intractable, it can be consistently estimated from samples with a known error bound.
-We find that our method effectively incorporates the beliefs specified by GP priors into the inference process (see figure 1) and provides competitive performance compared to BNN baselines.
+We find that our method effectively incorporates the beliefs specified by GP priors (see figure 1, discussed further in section 4) and that it yields competitive performance compared to BNN baselines.
 To the best of our knowledge, our method is the first to provide a well-defined objective for function-space inference in BNNs with informative GP priors.
@@
   1. 1.
-    We propose to use generalized VI with the so-called regularized KL divergence to mitigate the issue of an infinite KL divergence when using VI in BNNs with function-space priors.
+    We use generalized VI with the regularized KL divergence to mitigate the issue of an infinite KL divergence when using VI in BNNs with function-space priors.
   2. 2.
-    We present a new, well-defined objective for function-space inference in the linearized BNN with GP priors, which results in a simple algorithm.
+    We present a new and well-defined objective for function-space inference in the linearized BNN with GP priors, resulting in a simple algorithm.
   3. 3.
-    We analyze our method empirically on synthetic and real-world data sets, and find that it accurately captures structural properties specified by the GP prior and provides competitive uncertainty estimates for regression, classification, and out-of-distribution detection compared to BNN baselines with both function-space and weight-space priors.
-The paper is structured as follows.
-section 2 introduces function-space VI in BNNs and the regularized KL divergence.
-section 3 presents our proposed method for generalized function-space VI (GFSVI) in BNNs.
-section 4 presents experimental results.
-We discuss related work in section 5 and conclude in section 6.
-Figure 1: BNN inference on synthetic data (gray circles) with Gaussian process priors encoding different properties such as smoothness (increasing from Matern-1/2 to RBF) and periodicity (periodic).
+    We show that our method accurately captures structural properties specified by the GP prior and provides competitive uncertainty estimates for regression, classification, and out-of-distribution detection compared to baselines with both function- and weight-space priors.
+The paper is structured as follows: section 2 introduces function-space VI and the regularized KL divergence; section 3 presents our method for generalized function-space VI (GFSVI) in BNNs; section 4 reports experimental results; section 5 discusses related work; and section 6 concludes.
 2 Background
-In this section, we provide background on function-space variational inference (VI) in BNNs and discuss the fundamental issue of an infinite KL-divergence.
+In this section, we provide background on function-space variational inference in BNNs and discuss the fundamental issue of an infinite KL divergence.
 We then introduce the regularized KL divergence, which is the basis for our solution presented in Section 3.
@@
 We consider a neural network $f({\,\cdot\,;{\bm{w}}})$ with weights ${\bm{w}}\in{}^{p}$, and a data set $\mathcal{D}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}$ with features ${\bm{x}}_{i}\in\mathcal{X}\subset{}^{d}$ and values $y_{i}\in\mathcal{Y}$.
-Bayesian Neural Networks (BNNs) are specified further by a likelihood function $p\mathopen{}\left(\mathcal{D}\,|\,{\bm{w}}\right)\mathclose{}=\prod_{i=1}^{N}p\mathopen{}\left(y_{i}\,|\,f({\bm{x}}_{i};{\bm{w}})\right)\mathclose{}$ and—traditionally—a prior $p\mathopen{}\left({\bm{w}}\right)\mathclose{}$ on the weights, and one seeks the posterior distribution $p\mathopen{}\left({\bm{w}}\,|\,\mathcal{D}\right)\mathclose{}\propto p\mathopen{}\left(\mathcal{D}\,|\,{\bm{w}}\right)\mathclose{}\,p\mathopen{}\left({\bm{w}}\right)\mathclose{}$.
-The method proposed in this paper builds on variational inference (VI), which approximates $p\mathopen{}\left({\bm{w}}\,|\,\mathcal{D}\right)\mathclose{}$ with a variational distribution $q_{\phi}\mathopen{}\left({\bm{w}}\right)\mathclose{}$, whose variational parameters $\phi$ maximize the evidence lower bound (ELBO),
+Bayesian Neural Networks are specified further by a likelihood function $p\mathopen{}\left(\mathcal{D}\,|\,{\bm{w}}\right)\mathclose{}=\prod_{i=1}^{N}p\mathopen{}\left(y_{i}\,|\,f({\bm{x}}_{i};{\bm{w}})\right)\mathclose{}$ and—traditionally—a prior $p\mathopen{}\left({\bm{w}}\right)\mathclose{}$ on the weights, and one seeks the posterior distribution $p\mathopen{}\left({\bm{w}}\,|\,\mathcal{D}\right)\mathclose{}\propto p\mathopen{}\left(\mathcal{D}\,|\,{\bm{w}}\right)\mathclose{}\,p\mathopen{}\left({\bm{w}}\right)\mathclose{}$.
+The method proposed in this paper builds on variational inference, which approximates $p\mathopen{}\left({\bm{w}}\,|\,\mathcal{D}\right)\mathclose{}$ with a variational distribution $q_{\phi}\mathopen{}\left({\bm{w}}\right)\mathclose{}$, whose variational parameters $\phi$ maximize the evidence lower bound (ELBO),
 $\mathcal{L}(\phi):={\mathop{{}\mathbb{E}}}_{q_{\phi}\mathopen{}\left({\bm{w}}\right)\mathclose{}}\mathopen{}\left[\log p\mathopen{}\left(\mathcal{D}\,|\,{\bm{w}}\right)\mathclose{}\right]\mathclose{}-D_{\mathrm{KL}}\mathopen{}\left(q_{\phi}\,\|\,p\right)\mathclose{}$ (2.1)
-where $D_{\text{KL}}$ is the Kullback-Leibler (KL) divergence
+where $D_{\text{KL}}$ is the Kullback-Leibler divergence,
 $D_{\mathrm{KL}}\mathopen{}\left(q_{\phi}\,\|\,p\right)\mathclose{}:={\mathop{{}\mathbb{E}}}_{q_{\phi}\mathopen{}\left({\bm{w}}\right)\mathclose{}}\mathopen{}\left[\log\mathopen{}\left(q_{\phi}\mathopen{}\left({\bm{w}}\right)\mathclose{}\big{/}p\mathopen{}\left({\bm{w}}\right)\mathclose{}\right)\mathclose{}\right]\mathclose{}.$ (2.2)
 At test time, we approximate the predictive distribution for given features ${\bm{x}}^{*}$ as $p\mathopen{}\left(y^{*}\,|\,{\bm{x}}^{*}\right)\mathclose{}={\mathop{{}\mathbb{E}}}_{p\mathopen{}\left({\bm{w}}\,|\,\mathcal{D}\right)\mathclose{}}\mathopen{}\left[\big{.}p\mathopen{}\left(y^{*}\,|\,f({\bm{x}}^{*};{\bm{w}})\right)\mathclose{}\right]\mathclose{}\approx{\mathop{{}\mathbb{E}}}_{q_{\phi}({\bm{w}})}\mathopen{}\left[\big{.}p\mathopen{}\left(y^{*}\,|\,f({\bm{x}}^{*};{\bm{w}})\right)\mathclose{}\right]\mathclose{}$.
-Function-space variational inference
-Since weights of neural networks are not interpretable, we abandon the weight-space prior $p({\bm{w}})$ and instead pose a prior $\mathbb{P}$ directly on the function $f\mathopen{}\left(\,\cdot\,;{\bm{w}}\right)\mathclose{}$, which we denote simply as $f$ when there is no ambiguity.
+Function-space variational inference.
+Since neural network weights are not interpretable, we replace the weight-space prior $p({\bm{w}})$ with a prior $\mathbb{P}$ directly on the function $f\mathopen{}\left(\,\cdot\,;{\bm{w}}\right)\mathclose{}$, which we denote simply as $f$ when there is no ambiguity.
 Here, the symbol $\mathbb{P}$ denotes a probability measure that does not admit a density since the function space is infinite-dimensional.
 We compute the expected log-likelihood as in the first term of Eq.
-2.1, but we calculate the KL-term (Eq.
-2.2) using the pushforward of $q_{\phi}({\bm{w}})$ along the mapping ${\bm{w}}\mapsto f\mathopen{}\left(\,\cdot\,;{\bm{w}}\right)\mathclose{}$, which defines the variational measure $\mathbb{Q}_{\phi}$, resulting in the ELBO in function space,
+2.1.
+For the KL-term (Eq.
+2.2), a naive VI-method would use the pushforward of $q_{\phi}({\bm{w}})$ along ${\bm{w}}\mapsto f\mathopen{}\left(\,\cdot\,;{\bm{w}}\right)\mathclose{}$, which defines the variational measure $\mathbb{Q}_{\phi}$, resulting in the ELBO in function space,
 $\mathcal{L}(\phi):={\mathop{{}\mathbb{E}}}_{q_{\phi}\mathopen{}\left({\bm{w}}\right)\mathclose{}}\mathopen{}\left[\log p\mathopen{}\left(\mathcal{D}\,|\,{\bm{w}}\right)\mathclose{}\right]\mathclose{}-D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}$ (2.3)
-with $D_{\text{KL}}$ the KL divergence between measures given by
+with $D_{\text{KL}}$ the KL divergence between measures
 […]
@@
 2.1, the ELBO in Eq.
-2.3 is a lower bound on the evidence (Burt et al.,, 2020).
+2.3 is a lower bound on the evidence [Wild et al., 2022a].
 In fact, if $\mathbb{P}$ is the push-forward of $p({\bm{w}})$ then Eq.
@@
 Two intractabilities prevent directly maximizing the ELBO in function space (Eq 2.3).
-First, it is not obvious how to practically evaluate or estimate the KL divergence between two measures in function space in Eq 2.4.
-Sun et al., (2019) showed that it can be expressed as a supremum of KL divergences between finite-dimensional distributions,
-$D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}=\sup_{{\bm{x}}\in\mathcal{X}^{M}}D_{\mathrm{KL}}\mathopen{}\left(q_{\phi}(f({\bm{x}}))\,\|\,p(f({\bm{x}}))\right)\mathclose{}.$ (2.5)
+First, it is not obvious how to evaluate or estimate the KL divergence between two measures in Eq 2.4.
+Sun et al.
+[2019] showed that it can be expressed as a supremum of KL divergences between finite-dimensional distributions,
+$D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}=\sup_{{\bm{x}}\in\mathcal{X}^{M},M\in\mathbb{N}}D_{\mathrm{KL}}\mathopen{}\left(q_{\phi}(f({\bm{x}}))\,\|\,p(f({\bm{x}}))\right)\mathclose{}.$ (2.5)
 Here, ${\bm{x}}=\{{\bm{x}}^{(i)}\}_{i=1}^{M}\in\mathcal{X}^{M}$ is a set of $M$ points in feature space $\mathcal{X}$, and $q_{\phi}(f({\bm{x}}))$ and $p(f({\bm{x}}))$ are densities of the marginals of $\mathbb{Q}_{\phi}$ and $\mathbb{P}$ on $\{f({\bm{x}}^{(i)})\}_{i=1}^{M}$ respectively.
-To obtain a tractable approximation of the supremum over infinitely many sets, Sun et al., (2019) replace it by an expectation and Rudner et al., 2022b estimate it from samples.
-Second, we cannot express the pushforward measure $\mathbb{Q}_{\phi}$ of $q_{\phi}({\bm{w}})$ in a closed form because the neural network is nonlinear.
-Previous work has proposed to mitigate this issue using implicit score function estimators (Sun et al.,, 2019) or a linearized BNN $f_{L}$ to obtain a closed-form Gaussian variational measure (Rudner et al., 2022a, ; Rudner et al., 2022b, ).
-Our proposal in Section 3 follows the linearized BNN approach as it only minimally modifies the BNN, preserving most of its inductive bias (Maddox et al.,, 2021) while considerably simplifying the problem by turning the pushforward of $q_{\phi}({\bm{w}})$ into a Gaussian process (GP).
-More specifically, we consider a Gaussian variational distribution $q_{\phi}\mathopen{}\left({\bm{w}}\right)\mathclose{}=\mathcal{N}\mathopen{}\left({\bm{m}},{\bm{S}}\right)\mathclose{}$ with parameters $\phi=\left\{{\bm{m}},{\bm{S}}\right\}$, and we obtain $f_{L}$ by linearizing $f$ in the weights around ${\bm{w}}={\bm{m}}$,
-$f_{L}({\bm{x}};{\bm{w}}):=f({\bm{x}};{\bm{m}})+J({\bm{x}};{\bm{m}})({\bm{w}}-{\bm{m}})\qquad\text{where}\qquad J({\bm{x}};{\bm{m}})=\nabla_{\!{\bm{w}}}f({{\bm{x}};{\bm{w}}})|_{{\bm{w}}={\bm{m}}}.$ (2.6)
-Thus, for ${\bm{w}}\sim q_{\phi}({\bm{w}})$ we have, $f_{L}({\bm{x}};{\bm{w}})\sim\mathcal{N}\mathopen{}\left(f({\bm{x}};{\bm{m}}),J({\bm{x}};{\bm{m}}){\bm{S}}J({\bm{x}};{\bm{m}})^{\top}\right)\mathclose{}\;\forall{\bm{x}}$, and the function $f_{L}(\,\cdot\,;{\bm{w}})$ is a degenerate GP (as $\operatorname{rank}(J(\,\cdot\,;{\bm{m}}){\bm{S}}J(\,\cdot\,;{\bm{m}})^{\top})\leq\text{number\leavevmode\nobreak\ of\leavevmode\nobreak\ weights}<\infty$),
+Sun et al.
+[2019] approximates the supremum over infinitely many sets by an expectation, and Rudner et al.
+[2022b] estimates it from samples.
+Second, we cannot express the pushforward measure $\mathbb{Q}_{\phi}$ in closed form because the neural network is nonlinear.
+Previous work has proposed to mitigate this issue using implicit score function estimators [Sun et al., 2019] or a linearized BNN $f_{L}$ to obtain a closed-form Gaussian variational measure [Rudner et al., 2022a, b].
+Our proposal in Section 3 follows the linearized BNN approach as it only minimally modifies the BNN, preserving most of its inductive bias [Maddox et al., 2021] while considerably simplifying the problem by turning the pushforward of $q_{\phi}({\bm{w}})$ into a GP.
+More specifically, we consider a Gaussian variational distribution $q_{\phi}\mathopen{}\left({\bm{w}}\right)\mathclose{}=\mathcal{N}\mathopen{}\left({\bm{m}},{\bm{S}}\right)\mathclose{}$ with parameters $\phi=\left\{{\bm{m}},{\bm{S}}\right\}$, and we define a linearized BNN $f_{L}$ by linearizing $f$ as a function of the weights around ${\bm{w}}={\bm{m}}$,
+$f_{L}({\bm{x}};{\bm{w}}):=f({\bm{x}};{\bm{m}})+J({\bm{x}};{\bm{m}})({\bm{w}}-{\bm{m}})$ (2.6)
+with $J({\bm{x}};{\bm{m}})=\nabla_{\!{\bm{w}}}f({{\bm{x}};{\bm{w}}})|_{{\bm{w}}={\bm{m}}}$.
+Thus, ${\bm{w}}\sim q_{\phi}({\bm{w}})$ implies $f_{L}({\bm{x}};{\bm{w}})\sim\mathcal{N}\mathopen{}\left(f({\bm{x}};{\bm{m}}),J({\bm{x}};{\bm{m}}){\bm{S}}J({\bm{x}};{\bm{m}})^{\top}\right)\mathclose{}$ for all ${\bm{x}}\in\mathcal{X}$, and so the function $f_{L}(\,\cdot\,;{\bm{w}})$ is a degenerate GP (as $\operatorname{rank}(J(\,\cdot\,;{\bm{m}}){\bm{S}}J(\,\cdot\,;{\bm{m}})^{\top})\leq p<\infty$),
 $f_{L}\sim\mathcal{GP}\mathopen{}\left(f(\,\cdot\,;{\bm{m}}),J(\,\cdot\,;{\bm{m}}){\bm{S}}J(\,\cdot\,;{\bm{m}})^{\top}\right)\mathclose{}.$ (2.7)
-$D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}$ is infinite in most practically relevant cases.
-Burt et al., (2020) point out an even more severe issue of function-space VI in BNNs: $D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}$ is in fact infinite in many relevant cases, in particular for non-degenerate GP-priors with BNNs.
-The proof in Burt et al., (2020) is somewhat involved, but the fundamental reason for $D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}=\infty$ is that $\mathbb{Q}_{\phi}$ has support on a finite-dimensional submanifold of the infinite-dimensional function space, while the measure $\mathbb{P}$ induced by a (non-degenerate) GP prior has support on the entire function space.
-That such a dimensionality mismatch can lead to infinite KL divergence can be seen in a finite-dimensional example: consider the KL-divergence between two Gaussians in $\mathbb{R}^{n}$ for some $n\geq 2$, one of which has support on the entire $\mathbb{R}^{n}$ (i.e., its covariance matrix ${\bm{\Sigma}}_{1}$ has full rank) while the other one has support only on a proper subspace of $\mathbb{R}^{n}$ (i.e., its covariance matrix ${\bm{\Sigma}}_{2}$ is singular).
-The KL divergence between two multivariate Gaussians has a closed form expression (Eq 2.9 with ${\gamma=0}$) that contains $\log\det\mathopen{}\left({\bm{\Sigma}}_{2}^{-1}{\bm{\Sigma}}_{1}\right)\mathclose{}$, which is infinite for singular ${\bm{\Sigma}}_{2}$.
-We find that the fact that $D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}=\infty$ has severe practical consequences even when the KL divergence is only estimated from samples.
-It naturally explains the stability issues discussed in Appendix D.1 of Sun et al., (2019) (we compare the authors’ solution to this stability issue to our method in Section 3.2).
-Surprisingly, similar complications arise even in the setup by Rudner et al., 2022b , which performs VI in function space with the pushforward of a weight-space prior.
-While the KL divergence is technically finite because prior and variational posterior have the same support, numerical errors lead to mismatching supports and thus to stability issues even there.
-In summary, the ELBO for VI in BNNs is not well-defined for a large class of interesting function-space priors.
+$D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}$ is infinite in most relevant cases.
+Wild et al.
+[2022a] point out an even more severe issue of function-space VI in BNNs: $D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}$ (Eq.
+2.4) is in fact infinite in most relevant cases, in particular for non-degenerate GP-priors.
+Thus, approximating $D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}$ in these settings is futile.
+Their proof is somewhat involved, but the fundamental reason for $D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}=\infty$ is that $\mathbb{Q}_{\phi}$ has support on a finite-dimensional submanifold of the infinite-dimensional function space, while the measure $\mathbb{P}$ induced by a (non-degenerate) GP prior has support on the entire function space.
+That such a dimensionality mismatch can lead to infinite KL divergence can already be seen in a finite-dimensional example: consider the KL-divergence between two Gaussians in $\mathbb{R}^{n}$ for $n\geq 2$, one of which has support on the entire $\mathbb{R}^{n}$ (i.e., its covariance matrix ${\bm{\Sigma}}_{1}$ has full rank) while the other one has support only on a proper subspace of $\mathbb{R}^{n}$ (i.e., its covariance matrix ${\bm{\Sigma}}_{2}$ is singular).
+The KL divergence between multivariate Gaussians has a closed form expression (Eq.
+2.9 with ${\gamma=0}$) that contains $\log\det\mathopen{}\left({\bm{\Sigma}}_{2}^{-1}{\bm{\Sigma}}_{1}\right)\mathclose{}$, which is infinite for singular ${\bm{\Sigma}}_{2}$.
+We find that the fact that $D_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}\,\|\,\mathbb{P}\right)\mathclose{}=\infty$ has severe practical consequences even when the KL divergence is only estimated from finite samples.
+It naturally explains the stability issues discussed in Appendix D.1 of Sun et al.
+[2019] (we compare the authors’ solution to this stability issue to our method in Section 3.2).
+Surprisingly, similar complications arise even in the setup by Rudner et al.
+[2022b], which performs VI in function space with the pushforward of a weight-space prior.
+While this makes the KL divergence technically finite because prior and variational posterior have the same support, numerical errors lead to mismatching supports and thus to stability issues even there.
+In summary, the ELBO for VI in BNNs is not well-defined for most interesting function-space priors.
 In section 3, we propose a solution by using the so-called regularized KL divergence, which we introduce next.
-2.2 Regularized KL divergence for Gaussian measures
+2.2 Regularized KL divergence
 Our solution to the negative infinite function-space ELBO builds on a regularized KL divergence, which is expressed in terms of Gaussian measures for the variational posterior and prior.
-We obtain these Gaussian measures from corresponding Gaussian processes (GP).
+We obtain these Gaussian measures from GPs.
 We first discuss under which conditions a GP induces a Gaussian measure, and then present the regularized KL divergence.
-Gaussian measures and Gaussian processes
+Gaussian measures and Gaussian processes.
 […]
-Definition 2.1 (Gaussian measure, Kerrigan et al., (2023), Definition 1).
+Definition 2.1 (Gaussian measure, Kerrigan et al. [2023], Definition 1).
 Let $(\Omega,\mathcal{B},\mathbb{P})$ be a probability space.
@@
 For every GRE $F$, there exists a unique mean element $m\in L^{2}(\mathcal{X},\rho)$ and a finite trace linear covariance operator $C:L^{2}(\mathcal{X},\rho)\mapsto L^{2}(\mathcal{X},\rho)$ such that $\langle g,F\rangle\sim\mathcal{N}\mathopen{}\left(\langle g,m\rangle,\langle Cg,g\rangle\right)\mathclose{}$ for all $g\in L^{2}(\mathcal{X},\rho)$.
-The pushforward of $\mathbb{P}$ along $F$ denoted $\mathbb{P}^{F}:=F_{\#}\mathbb{P}$ is a Gaussian (probability) measure on $L^{2}(\mathcal{X},\rho)$.
-Gaussian measures generalize Gaussian distributions to infinite-dimensional spaces where measures do not have associated densities.
-Following Wild et al., (2022), we notate the Gaussian measure obtained from the GRE $F$ with mean element $m$ and covariance operator $C$ as $\mathbb{P}^{F}:=\mathcal{N}\mathopen{}\left(m,C\right)\mathclose{}$.
+The pushforward of $\mathbb{P}$ along $F$, denoted $\mathbb{P}^{F}:=F_{\#}\mathbb{P}$, is a Gaussian measure on $L^{2}(\mathcal{X},\rho)$.
 […]
-Definition 2.2 (Regularized KL divergence, Quang, (2022) Definition 5).
+Definition 2.2 (Regularized KL divergence, Quang [2022] Definition 5).
 Let $\nu_{1}=\mathcal{N}\mathopen{}\left(m_{1},C_{1}\right)\mathclose{}$ and $\nu_{2}=\mathcal{N}\mathopen{}\left(m_{2},C_{2}\right)\mathclose{}$ be two Gaussian measures with $m_{1},m_{2}\in L^{2}(\mathcal{X},\rho)$ and $C_{1},C_{2}$ bounded, self-adjoint, positive and trace-class linear operators on $L^{2}(\mathcal{X},\rho)$.
-Let $\gamma\in$, $\gamma>0$ be fixed.
+Let $\gamma\in{}_{>0}$ be fixed.
 The regularized KL divergence is defined as follows,
-$D^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\nu_{1}\,\|\,\nu_{2}\right)\mathclose{}:=$ $\,\frac{1}{2}\langle m_{1}-m_{2},(C_{2}+\gamma\mathbb{I})^{-1}\mathopen{}\left(m_{1}-m_{2}\right)\mathclose{}\rangle+\frac{1}{2}\operatorname{Tr}_{X}\left[\mathopen{}\left(C_{2}+\gamma\mathbb{I}\right)\mathclose{}^{-1}\mathopen{}\left(C_{1}+\gamma\mathbb{I}\right)\mathclose{}-\mathbb{I}\right]$
-$-\frac{1}{2}\log\operatorname{det}_{X}\left[\mathopen{}\left(C_{2}+\gamma\mathbb{I}\right)\mathclose{}^{-1}\mathopen{}\left(C_{1}+\gamma\mathbb{I}\right)\mathclose{}\right].$ (2.8)
-For any $\gamma>0$, the regularized KL divergence is well-defined and finite, even if the Gaussian measures are singular (Quang,, 2019).
-It converges to the conventional KL divergence (if it is well-defined) for $\gamma\to 0$ (Quang,, 2022, Theorem 6).
-Furthermore, and importantly for our application to VI, if two Gaussian measures $\nu_{1}$ and $\nu_{2}$ are induced by GPs $f_{i}\sim\mathcal{GP}(\mu_{i},K_{i})$ for $i=1,2$, then $D^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\nu_{1}\,\|\,\nu_{2}\right)\mathclose{}$ can be consistently estimated (Quang,, 2022) from a finite number $M$ of samples using the estimator
-$\hat{D}^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\nu_{1}\,\|\,\nu_{2}\right)\mathclose{}:=$ $\,\frac{1}{2}\mathopen{}\left({\bm{m}}_{1}-{\bm{m}}_{2}\right)\mathclose{}^{\top}({\bm{\Sigma}}^{(\gamma)}_{2})^{-1}\mathopen{}\left({\bm{m}}_{1}-{\bm{m}}_{2}\right)\mathclose{}+\frac{1}{2}\operatorname{Tr}\big{[}({\bm{\Sigma}}_{2}^{(\gamma)})^{-1}{\bm{\Sigma}}^{(\gamma)}_{1}-\mathbb{I}_{M}\big{]}$
-$-\frac{1}{2}\log\det\big{[}({\bm{\Sigma}}_{2}^{(\gamma)})^{-1}{\bm{\Sigma}}^{(\gamma)}_{1}\big{]}$ (2.9)
+$D^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\nu_{1}\,\|\,\nu_{2}\right)\mathclose{}:=\frac{1}{2}\langle m_{1}-m_{2},(C_{2}+\gamma\mathbb{I})^{-1}\mathopen{}\left(m_{1}-m_{2}\right)\mathclose{}\rangle$
+$\qquad+\frac{1}{2}\operatorname{Tr}_{X}\left[\mathopen{}\left(C_{2}+\gamma\mathbb{I}\right)\mathclose{}^{-1}\mathopen{}\left(C_{1}+\gamma\mathbb{I}\right)\mathclose{}-\mathbb{I}\right]$
+$\qquad-\frac{1}{2}\log\operatorname{det}_{X}\left[\mathopen{}\left(C_{2}+\gamma\mathbb{I}\right)\mathclose{}^{-1}\mathopen{}\left(C_{1}+\gamma\mathbb{I}\right)\mathclose{}\right].$ (2.8)
+Here $\operatorname{Tr}_{X}$ and $\operatorname{det}_{X}$ are the extended trace and extended Fredholm determinant [Quang, 2022].
+For any $\gamma>0$, the regularized KL divergence is well-defined and finite (following Quang [2017, Proposition 1]), even if the Gaussian measures are singular [Quang, 2019].
+It converges to the conventional KL divergence (if it is well-defined) for $\gamma\to 0$ (Quang, 2022, Theorem 6).
+Furthermore, if the Gaussian measures $\nu_{1}$ and $\nu_{2}$ are induced by GPs $\mathcal{GP}(\mu_{i},K_{i})$ for $i=1,2$, respectively, then $D^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\nu_{1}\,\|\,\nu_{2}\right)\mathclose{}$ is consistently estimated [Quang, 2022] by
+$\hat{D}^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\nu_{1}\,\|\,\nu_{2}\right)\mathclose{}:=$ $\,\frac{1}{2}\mathopen{}\left({\bm{m}}_{1}-{\bm{m}}_{2}\right)\mathclose{}^{\top}({\bm{\Sigma}}^{(\gamma)}_{2})^{-1}\mathopen{}\left({\bm{m}}_{1}-{\bm{m}}_{2}\right)\mathclose{}$
+$+$ $\,\frac{1}{2}\operatorname{Tr}\big{[}({\bm{\Sigma}}_{2}^{(\gamma)})^{-1}{\bm{\Sigma}}^{(\gamma)}_{1}-\mathbb{I}_{M}\big{]}$
+-- $\,\frac{1}{2}\log\det\big{[}({\bm{\Sigma}}_{2}^{(\gamma)})^{-1}{\bm{\Sigma}}^{(\gamma)}_{1}\big{]}$ (2.9)
 […]
@@
 This section presents our proposed generalized function-space variational inference (GFSVI) method, which addresses the problem of the infinite KL divergence discussed in section 2.1, which we take for an indication that VI is too restrictive if one wants to use genuine function-space priors.
-We instead consider generalized variational inference (Knoblauch et al.,, 2019), which reinterprets the ELBO in Eq.
+We instead consider generalized variational inference [Knoblauch et al., 2022], which reinterprets the ELBO in Eq.
 2.1 as a regularized expected log-likelihood and explores alternative divergences for the regularizer.
 Specifically, we propose to use the regularized KL divergence.
-The presentation in this section builds heavily on tools introduced in section 2, which turn out to fit together perfectly: the pushforward of a Gaussian variational distribution in weight-space through the linearized neural network (Eq.
+This section builds heavily on tools introduced in section 2, which turn out to fit together perfectly: the pushforward of a Gaussian variational distribution in weight-space through the linearized neural network (Eq.
 2.6) induces a GP variational posterior (Eq.
-2.7) that admits a Gaussian measure on $L^{2}(\mathcal{X},\rho)$ (definition 2.1).
-Further selecting a GP prior which has an associated Gaussian measure on $L^{2}(\mathcal{X},\rho)$ allows us to use the regularized KL divergence (Eq.
+2.7) that admits a Gaussian measure on $L^{2}(\mathcal{X},\rho)$.
+Further, selecting a GP prior which has an associated Gaussian measure on $L^{2}(\mathcal{X},\rho)$ allows us to use the regularized KL divergence (Eq.
 2.8).
-We present the GFSVI method in section 3.1 and discuss connections to prior work in section 3.2.
-3.1 Generalized function-space variational inference
-We present a simple and well-defined objective for function-space inference, and an algorithm for its optimization.
-Objective function
+We present GFSVI in section 3.1 and compare it to prior work in section 3.2.
+3.1 Generalized function-space VI
+We present a well-defined objective for function-space inference, and a simple algorithm for its optimization.
+Objective function.
 We start from the ELBO in Eq.
@@
 2.6).
-The function-space prior may be any GP that has an associated Gaussian measure $\mathbb{P}^{F}$ on $L^{2}(\mathcal{X},\rho)$ (see definition 2.1 and discussion below it).
+The function-space prior may be any GP that has an associated Gaussian measure $\mathbb{P}^{F}$ on $L^{2}(\mathcal{X},\rho)$.
 We now replace the KL divergence in the ELBO with the regularized KL divergence $D_{\text{KL}}^{\gamma}$ (Eq.
 2.8), which is well-defined and finite for any pair of Gaussian measures.
-For a likelihood $p\mathopen{}\left(\mathcal{D}\,|\,{\bm{w}}\right)\mathclose{}=\prod_{i=1}^{N}p\mathopen{}\left(y_{i}\,|\,f_{L}({\bm{x}}_{i};{\bm{w}})\right)\mathclose{}$, we thus obtain
+For a likelihood function $p\mathopen{}\left(\mathcal{D}\,|\,{\bm{w}}\right)\mathclose{}=\prod_{i=1}^{N}p\mathopen{}\left(y_{i}\,|\,f_{L}({\bm{x}}_{i};{\bm{w}})\right)\mathclose{}$, we obtain
 $\mathcal{L}(\phi)\!:=\!\!\sum_{i=1}^{N}{\mathop{{}\mathbb{E}}}_{q_{\phi}({\bm{w}})}\mathopen{}\left[\log p\mathopen{}\left(y_{i}|f_{L}({\bm{x}}_{i};{\bm{w}})\right)\mathclose{}\right]\mathclose{}\!-\!D^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}^{F}\,\|\,\mathbb{P}^{F}\right)\mathclose{}.$ (3.1)
-Input: linearized BNN $f_{L}$ with measure $\mathbb{Q}_{\phi}^{F}$, Gaussian process prior $\mathcal{GP}(\mu,K)$ with measure $\mathbb{P}^{F}$, measurement point distribution $\rho({\bm{x}})$, data $\mathcal{D}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}$, batch size $B$, $\gamma>0$.
-for each minibatch $({\bm{x}}_{\mathcal{B}},y_{\mathcal{B}})\sim\mathcal{D}$ do
-       Calculate ^$\hat{\ell}_{1}=\frac{N}{B}{\mathop{{}\mathbb{E}}}_{q_{\phi}({\bm{w}})}\mathopen{}\left[\log p\mathopen{}\left(y_{\mathcal{B}}\,|\,f_{L}({\bm{x}}_{\mathcal{B}},{\bm{w}})\right)\mathclose{}\right]\mathclose{}$;
-      [-1.5ex] Draw measurement points ${\bm{x}}=\{{\bm{x}}^{(i)}\}_{i=1}^{M}\stackrel{{\scriptstyle\!\!\text{i.i.d.}\!\!}}{{\sim}}\rho({\bm{x}})$;
-      [-1.5ex] Calculate ^$\hat{\ell}_{2}=\hat{D}_{\text{KL}}^{\gamma}\big{(}\mathbb{Q}_{\phi}^{F}\,\big{|}\!\big{|}\,{\mathbb{P}^{F}}\big{)}$ using ${\bm{x}}$ (Eq. 2.9);
-      [-1.5ex] Calculate ^$\hat{\mathcal{L}}(\phi)=\hat{\ell}_{1}-\hat{\ell}_{2}$;
-       Update $\phi$ using a step in the direction $\nabla_{\!\phi}\hat{\mathcal{L}}(\phi)$;
-end for
-Algorithm 1 Generalized function-space variational inference (GFSVI)
-Estimation and optimization
+Estimation and optimization.
 […]
-Technical details ($\gamma$ and $\rho$)
-It turns out that increasing $\gamma$ reduces the influence of the prior on inference (see figure 17 in Appendix).
+Technical details ($\gamma$ and $\rho$).
+It turns out that increasing $\gamma$ reduces the influence of the prior on inference (see figure 20).
 At the same time, $\gamma$ acts as jitter that prevents numerical errors (see section 3.2).
 We recommend setting $\gamma$ large enough to avoid numerical errors but sufficiently small to strongly regularize the objective in Eq.
-3.1 (see figure 15 in Appendix).
-We found that GFSVI is robust to a wide range of values and fixed $\gamma=10^{-10}$.
-Furthermore, the probability measure $\rho$ defined with $L^{2}(\mathcal{X},\rho)$ has to assign non-zero probability to any open set of $\mathcal{X}$ to regularize the BNN on all of its support.
-Following Rudner et al., 2022b , we draw measurement points from a uniform distribution over $\mathcal{X}$ when using tabular data and explore different configurations (such as samples from other data sets) when using high-dimensional image data (see section A.3.4).
+3.1 (see figure 18 in appendix) and setting $M$ to the largest value allowed by the computational budget.
+We found that the estimator $\hat{D}_{\text{KL}}^{\gamma}\big{(}\mathbb{Q}_{\phi}^{F}\,\big{|}\!\big{|}\,{\mathbb{P}^{F}}\big{)}$ converges quickly to a finite value (especially for smooth kernels, see figure 20 in appendix), and that GFSVI is robust to a wide range of values (we fixed $\gamma=10^{-10}$).
+The probability measure $\rho$ for $L^{2}(\mathcal{X},\rho)$ has to assign non-zero probability to any open set of $\mathcal{X}$ to regularize the BNN on all of its support.
+Following Rudner et al.
+[2022b], we draw measurement points from a uniform distribution over $\mathcal{X}$ when using tabular data and explore different configurations (samples from other data sets) for high-dimensional image data (see section C.4).
+Figure 2: Inference on synthetic data (gray circles) using a Matérn-1/2 prior for function-space methods GFSVI and FVI. The proposed GFSVI provides the best approximation of the exact GP posterior.
+Algorithm 1 Generalized function-space variational inference (GFSVI)
+Linearized BNN $f_{L}$with measure $\mathbb{Q}_{\phi}^{F}$, GP prior $\mathcal{GP}(\mu,K)$with measure $\mathbb{P}^{F}$, measurement point distribution $\rho({\bm{x}})$, data $\mathcal{D}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}$, batch size $B$, $\gamma>0$.
+minibatch $({\bm{x}}_{\mathcal{B}},y_{\mathcal{B}})\sim\mathcal{D}$
+1:Calculate ^$\hat{\ell}_{1}=\frac{N}{B}{\mathop{{}\mathbb{E}}}_{q_{\phi}({\bm{w}})}\mathopen{}\left[\log p\mathopen{}\left(y_{\mathcal{B}}\,|\,f_{L}({\bm{x}}_{\mathcal{B}},{\bm{w}})\right)\mathclose{}\right]\mathclose{}$;
+2:Draw measurement points ${\bm{x}}=\{{\bm{x}}^{(i)}\}_{i=1}^{M}\stackrel{{\scriptstyle\!\!\text{i.i.d.}\!\!}}{{\sim}}\rho({\bm{x}})$;
+3:Calculate ^$\hat{\ell}_{2}=\hat{D}_{\text{KL}}^{\gamma}\big{(}\mathbb{Q}_{\phi}^{F}\,\big{|}\!\big{|}\,{\mathbb{P}^{F}}\big{)}$ using ${\bm{x}}$ (Eq. 2.9);
+4:Calculate ^$\hat{\mathcal{L}}(\phi)=\hat{\ell}_{1}-\hat{\ell}_{2}$
+[…]
+[…]
+[…]
 3.2 Connections to prior work
-TFSVI (Rudner et al., 2022b, ) and FVI (Sun et al.,, 2019) solve stability issues by introducing jitter/noise, which has a similar effect as the regularization in Eq.
+TFSVI [Rudner et al., 2022b] and FVI [Sun et al., 2019] solve stability issues by introducing jitter/white noise, which has a similar effect as the regularization in Eq.
 2.8.
-However, as mentioned in Section 2.1, TFSVI introduces jitter only to overcome numerical issues and is fundamentally restricted to prior specification in weight space since its function-space prior is the pushforward of a weight-space prior.
-FVI does not linearize the BNN, and therefore does not have access to an explicit variational measure in function space.
-This severely complicates the estimation of (gradients of) the KL divergence, and the authors resort to implicit score function estimators, which make their method difficult to use in practice (Ma and Hernández-Lobato,, 2021).
-Our proposed GFSVI does not suffer from these difficulties as the variational posterior is an explicit (Gaussian) measure.
-This allows us to estimate the (regularized) KL divergence without sampling any noise or using implicit score function estimators.
+However, TFSVI introduces jitter only to overcome numerical issues and is fundamentally restricted to prior specification in weight space since its function-space prior is the pushforward of a weight-space prior.
+Conversely, FVI adds white noise to prevent the KL divergence (Eq.
+2.4) to blow up as $M$ increases.
+However, FVI does not linearize the BNN, and hence does not have access to an explicit variational measure in function space.
+This severely complicates the estimation of (gradients of) the KL divergence in FVI, and the authors resort to implicit score function estimators, which make their method difficult to use in practice [Ma and Hernández-Lobato, 2021].
+Our proposed GFSVI does not suffer from these difficulties as the variational posterior is an explicit Gaussian measure.
+This allows us to estimate the regularized KL divergence without sampling any noise or using implicit score function estimators.
 4 Experiments
-Figure 2: Inference on synthetic data (gray circles) using an RBF prior for function-space methods GFSVI and FVI. The proposed GFSVI provides the best approximation of the exact GP posterior.
-In this section, we evaluate our generalized function-space variational inference (GFSVI) method qualitatively on synthetic data (Section 4.1) and quantitatively on real-world data (Section 4.2).
-We find that GFSVI accurately captures structural properties specified by the GP prior, that it approximates the exact Gaussian process (GP) posterior more faithfully than BNN baselines, and that it performs competitively on regression, classification and out-of-distribution detection tasks.
+In this section, we evaluate our generalized function-space variational inference (GFSVI) method qualitatively on synthetic data and quantitatively on real-world data.
+GFSVI accurately captures structural properties specified by the GP prior, and that it performs competitively on regression, classification and out-of-distribution detection tasks.
 We also discuss the influence of the BNN’s inductive biases.
-Baselines
-We compare the proposed GFSVI method to two weight-space inference methods (mean-field variational inference (Blundell et al.,, 2015) and linearized Laplace (Immer et al.,, 2021)) and to three function-space inference methods (FVI (Sun et al.,, 2019), TFVSI (Rudner et al., 2022b, ) and VIP (Ma et al.,, 2019), where TFSVI performs inference in function space but with the pushforward of a weight-space prior and VIP using a BNN prior).
+Baselines.
+We compare GFSVI to two weight-space inference methods: mean-field VI (MFVI) [Blundell et al., 2015] and linearized Laplace [Immer et al., 2021]; and to three function-space inference methods: FVI [Sun et al., 2019], TFVSI [Rudner et al., 2022b] and VIP [Ma et al., 2019] (TFSVI performs inference in function space but with the pushforward of a weight-space prior; VIP uses a BNN prior).
 All BNNs have the same architecture and fully-factorized Gaussian approximate posterior.
-We also include results for a sparse GP with a posterior mean parameterized by a neural network (GWI) (Wild et al.,, 2022), as well as for a Gaussian Process (GP) (Williams and Rasmussen,, 2006) (when the size of the dataset allows it), and for a sparse GP (Hensman et al.,, 2013) for regression tasks.
-Ẇe consider the GP, sparse GP and GWI as gold standards as they represent the exact (or near exact) posterior for models with GP priors.
-4.1 Qualitative results on synthetic data
-Regression
-We consider synthetic regression data with a one-dimensional feature-space $\mathcal{X}\subset$, where the values $y_{i}$ are randomly sampled around $\sin(2\pi x_{i})$ (circles in Figures 1-11, see Appendix A.3.1).
-The green lines in Figures 1-8 and 11 show functions sampled from the (approximate) posteriors, and the red lines are the inferred mean functions.
-Figure 2 compares GFSVI with an RBF GP prior to baselines and to the exact posterior.
-We find that GFSVI visually matches the true posterior best (the same holds for a Matern-1/2 prior, see Figure 5 in the Appendix).
-Figure 1 in the Introduction and Figure 6 in the Appendix show that GFSVI notably adapts to varying prior assumptions (varying smoothness and length scale, respectively).
+We also include results for a sparse GP with a posterior mean parameterized by a neural network (GWI) [Wild et al., 2022b], and for a Gaussian Process (GP) [Williams and Rasmussen, 2006] (when the size of the dataset allows it), and for a sparse GP [Hensman et al., 2013] for regression tasks.
+We consider the GP, sparse GP and GWI as gold standards as they represent the exact (or near exact) posterior for models with GP priors.
+Figure 3: Results for the ocean current modeling experiment.
+We report the norm of the mean velocity vectors and the squared errors.
+Unlike TFSVI, we find that GFSVI accurately captures ocean current dynamics.
+Qualitative results on synthetic data.
+We consider a $1$-dimensional regression task where the values $y_{i}$ are sampled around $\sin(2\pi x_{i})$ (circles in Figures 1-8 and 11) and the two moons 2-dimensional binary classification task [Pedregosa et al., 2011] (see Figures 9 and 10).
+For regression, the green lines show functions sampled from the (approximate) posteriors, and the red lines are the inferred mean functions.
+For classification, the first and second row show the inferred mean probability of class 1 (blue dots) and its 2-standard deviations with respect to posterior samples.
+More details in Appendix C.1.
+We find that GFSVI captures the beliefs of the RBF and Matérn-1/2 GP priors better than BNN-baselines in the regression setting (see figures 5 and 2) as well as in classification (see figures 9 and 10), and shows greater uncertainty outside of the support of the data.
+figures 1 and 6 show that GFSVI notably adapts to varying prior assumptions (varying smoothness and length scale, respectively).
 In addition, figures 4 and 8 in the Appendix show that GFSVI provides strong regularization when the data generative process is noisy, and that it can be trained with fewer measurement points $M$ than FVI without significant degradation.
-Classification
-We further consider a binary classification task derived from the 2-dimensional two moons data set (Pedregosa et al.,, 2011), see red and blue dots in Figures 9 and 10 in the Appendix.
-The first and second row in both figures show the inferred mean probability of class 1 (blue dots) and its 2-standard deviations with respect to posterior samples.
-Just like for regression, we find that GFSVI captures the beliefs of the RBF and Matern-1/2 GP priors better than all BNN-baselines, and shows greater uncertainty outside of the support of the data than the BNN baselines.
-Inductive biases
-figure 11 in the Appendix compares GFSVI to the exact GP-posterior across two different priors and three model architectures (details in Appendix A.3.1).
-We find that, with ReLU activations, small models are prone to underfitting for smooth priors (RBF), and to collapsing uncertainty for rough priors (Matern-1/2).
-By contrast, with smooth activations (Tanh), smaller models suffice, and they are compatible with most standard GP priors (the results shown in figure 11 extend to RBF, Matern, and Rational Quadratic in our experiments).
+Inductive biases.
+figure 11 in the Appendix compares GFSVI to the exact GP-posterior across two different priors and three model architectures (details in Appendix C.1).
+We find that, with ReLU activations, small models are prone to underfitting for smooth priors (RBF), and to collapsing uncertainty for rough priors (Matérn-1/2).
+By contrast, with smooth activations (Tanh), smaller models suffice, and they are compatible with most standard GP priors (the results shown in figure 11 extend to RBF, Matérn, and Rational Quadratic in our experiments).
 We also analyzed how the number $M$ of measurement points affects performance.
-figures 7 and 14 in Appendix show that capturing the beliefs of rough GP priors and estimating $D_{\text{KL}}^{\gamma}$ with these priors requires larger $M$.
+figures 7 and 17 in the appendix show that capturing the beliefs of rough GP priors and estimating $D_{\text{KL}}^{\gamma}$ with these priors requires larger $M$.
+4.1 Quantitative results on real-world data
 Table 1: Test expected log-likelihood (higher is better) of evaluated methods on regression datasets. GFSVI performs competitively compared to all BNN baselines and obtains the best mean rank.
@@
 Table 2: Test expected log-likelihood, accuracy, expected calibration error and OOD detection accuracy on MNIST and Fashion MNIST.
-Data Metric Function-space priors Weight-space priors GP-based
+Metric Function-space priors Weight-space priors GP-based
 GFSVI (rnd) GFSVI (kmnist) FVI (rnd) FVI (kmnist) TFSVI (rnd) TFSVI (kmnist) MFVI VIP Laplace GWI
@@
 ($\uparrow$) -0.853 $\pm$ 0.005 -0.997 $\pm$ 0.001 -0.925 $\pm$ 0.005 -0.975 $\pm$ 0.002 -0.802 $\pm$ 0.006 -0.779 $\pm$ 0.010 -0.805 $\pm$ 0.010 -0.790 $\pm$ 0.010 -0.826 $\pm$ 0.006 -0.792 $\pm$ 0.005
-4.2 Quantitative results on real-world data
-We study the fidelity of GFSVI compared to the true GP-posterior on real data, and evaluate GFSVI on regression, classification, and out-of-distribution detection tasks.
+We evaluate GFSVI on regression, classification, and out-of-distribution detection.
 In all tables, we bold the highest score and any score whose error bar (standard error) overlaps with the highest score’s error bar.
-Ocean current modeling
-Figure 3: Results for the ocean current modeling experiment.
-We report the mean velocity vectors and the squared errors of compared methods.
-Unlike TFSVI, we find that GFSVI accurately captures ocean current dynamics.
-We consider the problem of modeling ocean currents in the Gulf of Mexico to measure the extend to which GFSVI can incorporate knowledge specified via a GP prior on real-world data.
-Specifically, we follow the setup by Shalashilin, (2024) and Cinquin et al., (2024), and use the GulfDrifters dataset (Lilly and Pérez-Brunius,, 2021) to estimate ocean currents from $20$ two-dimensional velocity vectors collected from drifter buoys.
-We embed physical properties of fluid motions into the GP prior and to the neural networks by applying the Helmholtz decomposition (Berlinghieri et al.,, 2023; Cinquin et al.,, 2024).
-We compare our GFSVI to a GP and to TFSVI.
-More details can be found in section A.3.2.
-Table 3: Results for the ocean current modeling task. We find that incorporating prior knowledge via a Gaussian process prior in GFSVI improves score over weight-space priors in TFSVI.
-Dataset GFSVI (ours) TFSVI GP
-Log-like. -6.627 $\pm$ 0.753 -22.651 $\pm$ 2.947 -0.507 $\pm$ 0.000
-MSE -0.021 $\pm$ 0.002 -0.034 $\pm$ 0.003 -0.013 $\pm$ 0.000
-Results are shown in table 3 and in figure 3.
-We find that incorporating prior knowledge via an informative GP prior in GFSVI improves expected log-likelihood (Log-like.) and mean squared-error (MSE) over weight-space priors in TFSVI.
-However, for this relatively small dataset, exact GP inference is still possible, and it outperforms both BNN-based predictions.
-This suggests that the physically motivated GP describes the true fluid dynamics well enough that the additional inductive bias introduced by a neural network hurts performance rather than helping it.
-In the following, we consider experiments with larger datasets (making exact GP inference computationally infeasible in many cases), and where structural prior knowledge in function space exists but is not derived from strict laws of nature.
-Regression
-We assess the predictive performance of GFSVI on data sets from the UCI repository (Dua and Graff,, 2017) (described in table 6 in the Appendix).
-Table 1, and Table 7 in the Appendix, show expected log-likelihood and mean squared error, respectively.
-We perform 5-fold cross validation and report means and standard errors across the test folds.
-We also rank the methods for each dataset and report the mean rank of each method across all datasets.
-See section A.3.3 for more details.
-We find that GFSVI performs competitively compared to baselines and obtains the best mean rank for both metrics, matching the top performing methods on nearly all datasets.
-In particular, we find that using GP priors in the linearized BNN with GFSVI yields improvements over the weight-space priors used in TFSVI and that GFSVI performs slightly better than FVI.
+Table 3: Results for the ocean current modeling task.
+Metric GFSVI (ours) TFSVI VIP GP
+Log-like. -6.627 $\pm$ 0.753 -22.651 $\pm$ 2.947 -11.631 $\pm$ 3.171 -0.507 $\pm$ 0.000
+MSE -0.021 $\pm$ 0.002 -0.034 $\pm$ 0.003 -0.026 $\pm$ 0.001 -0.013 $\pm$ 0.000
+Ocean current modeling.
+We measure how well GFSVI can incorporate knowledge specified via a GP prior on real-world data by considering the problem of modeling ocean currents in the Gulf of Mexico.
+We follow the setup by Shalashilin [2024] and use the GulfDrifters dataset [Lilly and Pérez-Brunius, 2021] to estimate ocean currents from $20$ $2$-dimensional velocity vectors collected from drifter buoys.
+We embed physical properties of fluid motions into the GP prior and to the neural networks by applying the Helmholtz decomposition [Berlinghieri et al., 2023, Cinquin et al., 2024].
+We compare our GFSVI to a GP, to TFSVI and to VIP.
+More details can be found in section C.2.
+We find that incorporating knowledge via an informative GP prior in GFSVI improves performance over weight-space priors in TFSVI and VIP (see table 3 and figure 3).
+However, the GP outperforms both BNNs, which suggests that the physically motivated kernel describes the fluid dynamics well enough that the additional inductive bias introduced by a neural network hurts performance rather than helping it.
+In the following, we consider experiments with larger datasets (making exact GP inference computationally infeasible in many cases), and where structural prior knowledge in function space exists but is not derived from laws of nature.
 Table 4: Out-of-distribution accuracy (higher is better) of evaluated methods on regression datasets. GFSVI (ours) performs competitively on OOD detection and obtains the highest mean rank.
@@
 Mean rank 1.455 2.364 1.909 2.909 3.364 2.909 - - -
-Table 5: Average point-wise Wasserstein-2 distance (lower is better) between exact and approximate posterior of reported methods. GFSVI (ours) provides a more accurate approximation than FVI.
-Dataset Boston Concrete Energy Wine Yacht Mean rank
-GFSVI (ours) 0.0259 $\pm$ 0.0040 0.0499 $\pm$ 0.0029 0.0035 $\pm$ 0.0004 0.0571 $\pm$ 0.0097 0.0036 $\pm$ 0.0006 1.0
-FVI 0.0469 $\pm$ 0.0044 0.0652 $\pm$ 0.0037 0.0037 $\pm$ 0.0004 0.1224 $\pm$ 0.0167 0.0052 $\pm$ 0.0013 1.6
-GP sparse 0.0074 $\pm$ 0.0022 0.0125 $\pm$ 0.0016 0.0042 $\pm$ 0.0003 0.0170 $\pm$ 0.0035 0.0035 $\pm$ 0.0008 -
-Classification
-We further evaluate the classification performance of our method on the MNIST (LeCun et al.,, 2010) and FashionMNIST (Xiao et al.,, 2017) image data sets.
+Regression.
+We assess the predictive performance of GFSVI on data sets from the UCI repository [Dua and Graff, 2017].
+Table 1, and Table 6 in the appendix, show expected log-likelihood and mean squared error, respectively.
+We perform 5-fold cross validation and report means and standard errors across the test folds.
+We also rank the methods for each dataset and report the mean rank of each method across all datasets.
+See section C.3 for more details.
+We find that GFSVI performs competitively compared to baselines and obtains the best mean rank for both metrics, matching the top performing methods on nearly all datasets.
+In particular, we find that using GP priors in the linearized BNN with GFSVI yields improvements over the weight-space priors used in TFSVI, and that GFSVI performs slightly better than FVI despite being simpler.
+Further, we find that GFSVI approximates the exact GP-posterior more accurately that FVI (see table 7 and section D.3), and that it converges in slightly more steps than TFSVI (figure 15).
+Classification.
+We further evaluate classification performance of our method on the MNIST [LeCun et al., 2010] and FashionMNIST [Xiao et al., 2017] image data sets.
 We fit the models on a random subset of 90% of the training set, use the remaining 10% as validation data, and evaluate on the provided test split.
-We repeat with 5 different random seeds and report the mean and standard error of the expected log-likelihood, accuracy and expected calibration error (ECE) of the predictive distribution’s mean in table 2.
-For GFSVI, FVI, and TFSVI, we tested measurement points from both a random (rnd) distribution $\rho({\bm{x}})$ and from kmnist (see Algorithm 1 and section A.3.4).
-We find that GFSVI performs competitively on MNIST, exceeding the expected log-likelihood and accuracy of top-scoring baselines (see table 2) and similarly to best baselines on FashionMNIST.
-Additionally, GFSVI yields well-calibrated models with low ECE.
-Out-of-distribution detection
-We further evaluate our method by testing if its epistemic uncertainty is predictive of out-of-distribution (OOD) data.
-We consider two settings: (i) with tabular data and a Gaussian likelihood following Malinin et al., (2020), and (ii) with image data and a categorical likelihood following Osawa et al., (2019).
-We report the accuracy of classifying OOD from in-distribution (ID) data using a (learned) threshold on the predictive uncertainty.
-Additional details are provided in section A.3.5.
-In setting (i) (table 4), we find that GFSVI performs competitively and obtains the highest mean rank.
-Likewise, in setting (ii) (table 8 in the Appendix), GFSVI strongly outperforms all baselines when using the kmnist measurement point distribution $\rho({\bm{x}})$.
-We find that with high-dimensional image data, the choice of measurement point distribution highly influences OOD detection accuracy.
-We provide more details on this point in Appendix A.4.4.
-In both settings, we find that using GP priors with GFSVI rather than weight-space priors with TFSVI is beneficial, and that GFSVI improves over FVI.
-Variational measure evaluation
-table 5 evaluates our inference method by comparing samples drawn from the exact posterior (where computationally feasible) with the approximate posterior obtained with our method (GFSVI).
-More details are provided in section A.3.6.
-We find that GFSVI approximates the exact posterior more accurately that FVI, obtaining a higher mean rank, but worse than the gold standard sparse GP, which demonstrates to be most accurate.
+We repeat with 5 different random seeds and report the mean and standard error of the expected log-likelihood, accuracy, and expected calibration error (ECE) in table 2.
+For GFSVI, FVI, and TFSVI, we tested measurement points from both a uniform random (rnd) distribution $\rho({\bm{x}})$ and from kmnist.
+Details in section C.4.
+We find that GFSVI performs competitively on MNIST, exceeding the expected log-likelihood and accuracy of top-scoring baselines and similarly to best baselines on FashionMNIST.
+GFSVI also yields well-calibrated models with low ECE.
+Out-of-distribution detection.
+We next evaluate our method by testing if its epistemic uncertainty is predictive of out-of-distribution (OOD) data.
+We consider two settings: (i) with tabular data and a Gaussian likelihood [Malinin et al., 2021], and (ii) with image data and a categorical likelihood [Osawa et al., 2019].
+We report the accuracy of classifying OOD vs.
+in-distribution (ID) data using a (learned) threshold on the predictive uncertainty.
+More details in section C.5.
+In setting (i), GFSVI performs competitively and obtains the highest mean rank (table 4).
+Likewise in setting (ii), GFSVI strongly outperforms all baselines when using the kmnist measurement point distribution $\rho({\bm{x}})$ (figure 12, tables 2 and 8).
+We find that with high-dimensional image data, the choice of measurement point distribution highly influences OOD detection accuracy (see Appendix D.5 for a discussion).
+In both settings, using GP priors with GFSVI rather than weight-space priors with TFSVI is beneficial, and GFSVI also improves over FVI.
+GFSVI’s uncertainty is also well-calibrated under distribution shift of the input features (see section D.6).
 5 Related work
-In this section, we review related work on function-space variational inference with neural networks, and on approximating functions-space measures with weight-space priors.
-Function-space inference with neural networks
-Prior work on function-space VI in BNNs has addressed issues (i) and (ii) mentioned in Section 2.1.
-Sun et al., (2019) address (i) (intractable variational posterior in function space) by using implicit score function estimators, and (ii) (intractable KL divergence) by replacing the supremum with an expectation.
-Rudner et al., 2022a ; Rudner et al., 2022b address (i) by using a linearized BNN (Khan et al.,, 2020; Immer et al.,, 2021; Maddox et al.,, 2021), and (ii) by replacing the supremum with a maximum over a finite set of samples.
-Other work abandons approximating the posterior over neural network weights altogether and instead uses a BNN only to specify a prior (Ma et al.,, 2019), or deterministic neural networks to fit basis functions for Bayesian linear regression (Ma and Hernández-Lobato,, 2021) or the mean of a generalized sparse GP with Wasserstein-2 metric (Wild et al.,, 2022).
-Our work combines linearized BNNs with generalized variational inference, but we use the regularized KL divergence (Quang,, 2019), which naturally generalizes the conventional KL divergence and allows for informative GP priors.
-Approximating function-space measures with weight-space priors in BNNs
-Flam-Shepherd et al., (2017); Tran et al., (2022) minimize a divergence between the BNN’s prior predictive and a GP before performing inference on weights, while Wu et al., (2023) directly incorporate the bridging divergence inside the inference objective.
-Alternatively, Pearce et al., (2020) derive BNN architectures mirroring GPs, and Matsubara et al., (2022) use the Ridgelet transform to design weight-spaces priors approximating a GP in function space.
-Another line of work considers weight-space priors which regularize in function space by comparing the model’s predictions to those of a reference model (Nalisnick et al.,, 2020) and using an empirical prior (Rudner et al.,, 2023).
+In this section, we review related work on function-space VI with neural networks, and on approximating functions-space measures with weight-space priors.
+Function-space inference with neural networks.
+Prior work on function-space VI in BNNs has addressed issues (i) intractable variational posterior in function space and (ii) intractable KL divergence discussed in Section 2.1.
+Sun et al.
+[2019] address (i) by using implicit score function estimators, and (ii) by replacing the supremum with an expectation.
+Rudner et al.
+[2022b] address (i) by using a linearized BNN [Khan et al., 2019, Immer et al., 2021, Maddox et al., 2021], and (ii) by replacing the supremum with a maximum over a finite set.
+Other work abandons approximating the neural network’s posterior and instead uses a BNN to specify a prior [Ma et al., 2019], or deterministic neural networks as features for Bayesian linear regression [Ma and Hernández-Lobato, 2021] or the mean of a generalized sparse GP [Wild et al., 2022b].
+Unlike our more expressive GP posterior covariance, Wild et al.
+[2022b] uses a simple stationary sparse GP posterior covariance (table 2) which has higher sampling cost and can lead to model misspecification (figure 14).
+Our work combines linearized BNNs with generalized VI, but we use the regularized KL divergence [Quang, 2019], which naturally generalizes the KL divergence and allows for informative GP priors.
+Approximating function-space measures with weight-space priors.
+Flam-Shepherd et al.
+[2017], Tran et al.
+[2022] minimize a divergence between the BNN’s prior predictive and a GP before performing inference on weights, while Wu et al.
+[2023] directly incorporate the bridging divergence inside the inference objective.
+Alternatively, Pearce et al.
+[2020] derive BNN architectures mirroring GPs, and Matsubara et al.
+[2021] use the Ridgelet transform to design weight-spaces priors approximating a GP in function space.
+Similarly, Rudner et al.
+[2023] and Sam et al.
+[2024] use empirical weight-space priors to regularize in function space and encode domain knowledge specified via a loss function, respectively.
+Yang et al.
+[2020] instead imposes functional constraints directly via the prior.
 6 Discussion
-We proposed a simple inference method with a well-defined variational objective function for Bayesian neural networks with Gaussian process (GP) priors in function-space.
+We proposed a simple inference method with a well-defined variational objective function for BNNs with GP priors in function-space.
 As standard VI with functions-space priors suffers from an infinite KL divergence problem, we propose to follow the generalized VI framework.
 Specifically, we substitute the conventional KL divergence in the ELBO by the regularized KL divergence, which is always finite, and which can be estimated consistently within the linearized BNN approximation.
-We demonstrated that our method allows to incorporate interpretable structural properties via a GP prior, accurately approximates the true GP posterior on synthetic and small real-world data sets, and provides competitive uncertainty estimates for regression, classification and out-of-distribution detection compared to BNN baselines with both function-space and weight-space priors.
-Future work should investigate the use of more expressive variational distributions, such as Gaussian with low-rank plus diagonal covariance proposed by Tomczak et al., (2020), which is compatible with our proposed method.
-Acknowledgments
-This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645.
+We demonstrated that our method allows to incorporate interpretable structural properties via a GP prior, accurately approximates the true GP posterior on synthetic and small real-world data sets, and provides competitive uncertainty estimates for regression, classification and out-of-distribution detection compared to BNNs with both function-space and weight-space priors.
+Future work should investigate the use of more expressive variational distributions, such as Gaussian with low-rank plus diagonal covariance proposed by Tomczak et al.
+[2020].
+Acknowledgements.
+This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645, and by the German Research Foundation (DFG) under project 448588364 of the Emmy Noether Programme.
 Additional support was provided by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039A.
-Robert Bamler acknowledges funding by the German Research Foundation (DFG) for project 448588364 of the Emmy Noether Programme.
 The authors extend their gratitude to the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Tristan Cinquin.
-Finally, we thank Marvin Pförtner for very useful discussions on the theory and experiments, and Vincent Fortuin for his feedback.
-Appendix A Appendix
-A.1 Divergences between Gaussian measures
-A.1.1 The KL divergence is infinite
+Finally, Tristan Cinquin thanks Marvin Pförtner for very useful discussions on the theory and experiments, and Vincent Fortuin for feedback.
+Supplementary Material
+Appendix A Divergences between Gaussian measures
+A.1 The KL divergence is infinite
 In this section, we show that the Kullbach-Liebler (KL) divergence between the Gaussian measures $\mathbb{Q}_{\phi}^{F}\sim\mathcal{N}\mathopen{}\left(m_{Q},C_{Q}\right)\mathclose{}$ and $\mathbb{P}^{F}\sim\mathcal{N}\mathopen{}\left(m_{P},C_{P}\right)\mathclose{}$, respectively induced by the linearized BNN in Eq 2.7 and by a non-degenerate Gaussian process satisfying conditions given in section 2.2, is infinite.
-While this has already been shown by Burt et al., (2020), the proof is easier for Gaussian measures.
+While this has already been shown by Wild et al.
+[2022a], the proof is easier for Gaussian measures.
 We first need the Feldman-Hàjek theorem which tells us when the KL divergence between two Gaussian measures is well-defined.
-Theorem A.1 (Feldman-Hàjek, Quang, (2022) Theorem 2, Simpson, (2022) Theorem 7).
+Theorem A.1 (Feldman-Hàjek, Quang [2022] Theorem 2, Simpson [2022] Theorem 7).
 Consider two Gaussian measures $\nu_{1}=\mathcal{N}\mathopen{}\left(m_{1},C_{1}\right)\mathclose{}$ and $\nu_{2}=\mathcal{N}\mathopen{}\left(m_{2},C_{2}\right)\mathclose{}$ on $L^{2}(\mathcal{X},\rho)$.
@@
 otherwise $\nu_{1}$ and $\nu_{2}$ are singular.
-If $\nu_{1}$ and $\nu_{2}$ are equivalent, then the Radon-Nikodym dervative exists and $D_{\mathrm{KL}}\mathopen{}\left(\nu_{1}\,\|\,\nu_{2}\right)\mathclose{}$ admits an explicit formula.
+If $\nu_{1}$ and $\nu_{2}$ are equivalent, then the Radon-Nikodym derivative exists and $D_{\mathrm{KL}}\mathopen{}\left(\nu_{1}\,\|\,\nu_{2}\right)\mathclose{}$ admits an explicit formula.
 Otherwise, $D_{\mathrm{KL}}\mathopen{}\left(\nu_{1}\,\|\,\nu_{2}\right)\mathclose{}=\infty$.
@@
 ∎
-A.1.2 The regularized KL divergence
-In this section, we provide the bound describing the asymptotic convergence of the regularized KL divergence estimator.
-Theorem A.2 (Convergence of estimator, Quang, (2022) Theorem 45).
+A.2 The regularized KL divergence
+We provide the bound describing the asymptotic convergence of the regularized KL divergence estimator in equation A.1.
+The error results from the fact that taking a finite number M of context points effectively cuts off the spectra of the covariance operators and the estimator $\hat{D}_{\text{KL}}^{\gamma}$ converges to $D_{\text{KL}}^{\gamma}$ as $M\to\infty$ with high probability.
+Theorem A.2 (Convergence of estimator, Quang [2022] Theorem 45).
 Assume the following:
@@
 […]
-$|D_{\mathrm{KL}}\mathopen{}\left(\mathcal{N}\mathopen{}\left(\mu_{1}(\mathbf{x}),K_{1}(\mathbf{x},\mathbf{x})+M\gamma\mathbb{I}_{M}\right)\mathclose{}\,\|\,{\mathcal{N}\mathopen{}\left(\mu_{2}(\mathbf{x}),K_{2}(\mathbf{x},\mathbf{x})+M\gamma\mathbb{I}_{M}\right)\mathclose{}}\right)\mathclose{}\\ -D^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\mathcal{N}\mathopen{}\left(m_{1},C_{1}\right)\mathclose{}\,\|\,\mathcal{N}\mathopen{}\left(m_{2},C_{2}\right)\mathclose{}\right)\mathclose{}|\\ \leq\frac{1}{2\gamma}(B_{1}+B_{2})^{2}[1+\kappa_{2}^{2}/\gamma]\mathopen{}\left(\frac{2\log\frac{48}{\gamma}}{M}+\sqrt{\frac{2\log\frac{48}{\gamma}}{M}}\right)\mathclose{}\\ +\frac{1}{2\gamma^{2}}[\kappa_{1}^{4}+\kappa_{2}^{4}+\kappa_{1}^{2}\kappa_{2}^{2}(2+\kappa_{2}^{2}/\gamma)]\mathopen{}\left(\frac{2\log\frac{12}{\gamma}}{M}+\sqrt{\frac{2\log\frac{12}{\gamma}}{M}}\right)\mathclose{}$ (A.1)
-A.2 Additional details on the GFSVI objective estimator
+$|D_{\mathrm{KL}}\mathopen{}\left(\mathcal{N}\mathopen{}\left(\mu_{1}(\mathbf{x}),K_{1}(\mathbf{x},\mathbf{x})+M\gamma\mathbb{I}_{M}\right)\mathclose{}\,\|\,{\mathcal{N}\mathopen{}\left(\mu_{2}(\mathbf{x}),K_{2}(\mathbf{x},\mathbf{x})+M\gamma\mathbb{I}_{M}\right)\mathclose{}}\right)\mathclose{}-D^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\mathcal{N}\mathopen{}\left(m_{1},C_{1}\right)\mathclose{}\,\|\,\mathcal{N}\mathopen{}\left(m_{2},C_{2}\right)\mathclose{}\right)\mathclose{}|\\ \leq\frac{1}{2\gamma}(B_{1}+B_{2})^{2}[1+\kappa_{2}^{2}/\gamma]^{2}\mathopen{}\left(\frac{2\log\frac{48}{\delta}}{M}+\sqrt{\frac{2\log\frac{48}{\delta}}{M}}\right)\mathclose{}\\ +\frac{1}{2\gamma^{2}}[\kappa_{1}^{4}+\kappa_{2}^{4}+\kappa_{1}^{2}\kappa_{2}^{2}(2+\kappa_{2}^{2}/\gamma)]\mathopen{}\left(\frac{2\log\frac{12}{\delta}}{M}+\sqrt{\frac{2\log\frac{12}{\delta}}{M}}\right)\mathclose{}$ (A.1)
+Note that equation A.1 provides a very general bound on the error that does not make assumptions on the spectral decay, and it may therefore dramatically overestimate the error.
+Indeed, we analyze convergence empirically in figure 20 and observe that the estimator converges quickly except for very rough priors (e.g., Matérn-1/2) with very small $\gamma$.
+Appendix B Additional details on the GFSVI objective estimator
 In this section, we present details on the estimation of the generalized function-space variational inference (GFSVI) objective.
@@
 Assuming a likelihood $p\mathopen{}\left(\mathcal{D}\,|\,{\bm{w}}\right)\mathclose{}=\prod_{i=1}^{N}p\mathopen{}\left(y_{i}\,|\,f({\bm{x}}_{i};{\bm{w}})\right)\mathclose{}$ and a Gaussian variational distribution on model weights $q_{\phi}({\bm{w}})=\mathcal{N}\mathopen{}\left({\bm{w}}\,|\,{\bm{m}},{\bm{S}}\right)\mathclose{}$, the GFSVI objective function is
-$\mathcal{L}(\phi)=\sum_{i=1}^{N}{\mathop{{}\mathbb{E}}}_{q_{\phi}({\bm{w}})}\mathopen{}\left[\log p\mathopen{}\left(y_{i}\,|\,f_{L}({\bm{x}}_{i};{\bm{w}})\right)\mathclose{}\right]\mathclose{}-D^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}^{F}\,\|\,\mathbb{P}^{F}\right)\mathclose{}$ (A.2)
+$\mathcal{L}(\phi)=\sum_{i=1}^{N}{\mathop{{}\mathbb{E}}}_{q_{\phi}({\bm{w}})}\mathopen{}\left[\log p\mathopen{}\left(y_{i}\,|\,f_{L}({\bm{x}}_{i};{\bm{w}})\right)\mathclose{}\right]\mathclose{}-D^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}^{F}\,\|\,\mathbb{P}^{F}\right)\mathclose{}$ (B.1)
 where $\mathbb{Q}_{\phi}^{F}$ and $\mathbb{P}^{F}$ are the Gaussian measures induced by the linearized BNN and a Gaussian process prior respectively.
@@
 When considering a Gaussian likelihood, we use the closed form expression available due to the Gaussian variational measure over functions induced by the linearized BNN
-${\mathop{{}\mathbb{E}}}_{q_{\phi}({\bm{w}})}\mathopen{}\left[\log\mathcal{N}\mathopen{}\left(y_{i}\,|\,f_{L}({\bm{x}}_{i};{\bm{w}}),\sigma_{y}^{2}\right)\mathclose{}\right]\mathclose{}=-\frac{1}{2}\log\mathopen{}\left(2\pi\sigma_{y}^{2}\right)\mathclose{}-\frac{\mathopen{}\left(y_{i}-f\mathopen{}\left({\bm{x}}_{i};{\bm{m}}\right)\mathclose{}\right)\mathclose{}^{2}+J\mathopen{}\left({\bm{x}}_{i};{\bm{m}}\right)\mathclose{}{\bm{S}}J\mathopen{}\left({\bm{x}}_{i};{\bm{m}}\right)\mathclose{}^{\top}}{2\sigma_{y}^{2}}.$ (A.3)
+${\mathop{{}\mathbb{E}}}_{q_{\phi}({\bm{w}})}\mathopen{}\left[\log\mathcal{N}\mathopen{}\left(y_{i}\,|\,f_{L}({\bm{x}}_{i};{\bm{w}}),\sigma_{y}^{2}\right)\mathclose{}\right]\mathclose{}=-\frac{1}{2}\log\mathopen{}\left(2\pi\sigma_{y}^{2}\right)\mathclose{}-\frac{\mathopen{}\left(y_{i}-f\mathopen{}\left({\bm{x}}_{i};{\bm{m}}\right)\mathclose{}\right)\mathclose{}^{2}+J\mathopen{}\left({\bm{x}}_{i};{\bm{m}}\right)\mathclose{}{\bm{S}}J\mathopen{}\left({\bm{x}}_{i};{\bm{m}}\right)\mathclose{}^{\top}}{2\sigma_{y}^{2}}.$ (B.2)
 When considering a Categorical likelihood with $C$ different classes, we estimate the expected log-likelihood term using Monte-Carlo integration as
-$\begin{split}{\mathop{{}\mathbb{E}}}_{q_{\phi}({\bm{w}})}\mathopen{}\left[\log\text{Cat}(y_{i}\,|\,\sigma(f_{L}({\bm{x}}_{i};{\bm{w}})))\right]\mathclose{}&=\frac{1}{K}\sum_{k=1}^{K}\sum_{c=1}^{C}\mathbb{I}[y_{i}=c]\Bigl{[}f^{c}_{L}({\bm{x}}_{i};{\bm{w}}^{(k)})\\ &\qquad-\log\left[\sum_{c^{\prime}=1}^{C}\exp\left(f^{c^{\prime}}_{L}({\bm{x}}_{i};{\bm{w}}^{(k)})\right)\right]\Bigr{]}\end{split}$ (A.4)
+${\mathop{{}\mathbb{E}}}_{q_{\phi}({\bm{w}})}\mathopen{}\left[\log\text{Cat}(y_{i}\,|\,\sigma(f_{L}({\bm{x}}_{i};{\bm{w}})))\right]\mathclose{}=\frac{1}{K}\sum_{k=1}^{K}\sum_{c=1}^{C}\mathbb{I}[y_{i}=c]\left[f^{c}_{L}({\bm{x}}_{i};{\bm{w}}^{(k)})-\log\left[\sum_{c^{\prime}=1}^{C}\exp\left(f^{c^{\prime}}_{L}({\bm{x}}_{i};{\bm{w}}^{(k)})\right)\right]\right]$ (B.3)
 where ${\bm{w}}^{(k)}\sim q_{\phi}({\bm{w}})$ for $k=1,\dots,K$, $\mathbb{I}[\cdot]$ is the indicator function, $\sigma(\cdot)$ is the softmax function and $f^{c}_{L}(\,\cdot\,;{\bm{w}})$ is the logit for class $c$ obtained from $f_{L}$.
@@
 We estimate the regularized KL divergence using its consistent estimator (Eq. 2.9)
-$\hat{D}^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}^{F}\,\|\,\mathbb{P}^{F}\right)\mathclose{}=\frac{1}{2}\mathopen{}\left(f({\bm{x}};{\bm{m}})-\mu({\bm{x}})\right)\mathclose{}^{\top}(K({\bm{x}},{\bm{x}})+\gamma M\mathbb{I}_{M})^{-1}\mathopen{}\left(f({\bm{x}};{\bm{m}})-\mu({\bm{x}})\right)\mathclose{}\\ +\frac{1}{2}\operatorname{Tr}\big{[}(K({\bm{x}},{\bm{x}})+\gamma M\mathbb{I}_{M})^{-1}(J({\bm{x}};{\bm{m}})SJ({\bm{x}};{\bm{m}})^{\top}+\gamma M\mathbb{I}_{M})-\mathbb{I}_{M}\big{]}\\ -\frac{1}{2}\log\det\big{[}(K({\bm{x}},{\bm{x}})+\gamma M\mathbb{I}_{M})^{-1}(J({\bm{x}};{\bm{m}})SJ({\bm{x}};{\bm{m}})^{\top}+\gamma M\mathbb{I}_{M})\big{]}$ (A.5)
+$\hat{D}^{\gamma}_{\mathrm{KL}}\mathopen{}\left(\mathbb{Q}_{\phi}^{F}\,\|\,\mathbb{P}^{F}\right)\mathclose{}=\frac{1}{2}\mathopen{}\left(f({\bm{x}};{\bm{m}})-\mu({\bm{x}})\right)\mathclose{}^{\top}(K({\bm{x}},{\bm{x}})+\gamma M\mathbb{I}_{M})^{-1}\mathopen{}\left(f({\bm{x}};{\bm{m}})-\mu({\bm{x}})\right)\mathclose{}\\ +\frac{1}{2}\operatorname{Tr}\big{[}(K({\bm{x}},{\bm{x}})+\gamma M\mathbb{I}_{M})^{-1}(J({\bm{x}};{\bm{m}})SJ({\bm{x}};{\bm{m}})^{\top}+\gamma M\mathbb{I}_{M})-\mathbb{I}_{M}\big{]}\\ -\frac{1}{2}\log\det\big{[}(K({\bm{x}},{\bm{x}})+\gamma M\mathbb{I}_{M})^{-1}(J({\bm{x}};{\bm{m}})SJ({\bm{x}};{\bm{m}})^{\top}+\gamma M\mathbb{I}_{M})\big{]}$ (B.4)
 […]
-A.3 Additional details on the experimental setup
-A.3.1 Experiments on synthetic data
+Computational complexity
+Evaluating the objective in Eq.
+B.1 has complexity $O(BKC+M^{3})$ for Categorical likelihoods and $O(B+M^{3})$ for Gaussian likelihoods, where $B$ is the batch size, $K$ the number of variational posterior samples, $C$ the number of classes, and $M$ the number of context points.
+The first term corresponds to the expected log-likelihood in our objective and the second term to the regularized KL divergence estimator.
+We note that evaluating the linearized neural network can be efficiently done in about 3x the cost of one forward pass using the Jacobian-vector product computational primitive.
+Appendix C Additional details on the experimental setup
+C.1 Experiments on synthetic data
 Regression
 We consider the following generative model for the toy data
-$y_{i}=\operatorname{sin}(2\pi x_{i})+\epsilon\quad\text{with}\quad\epsilon\sim\mathcal{N}\mathopen{}\left(0,\sigma_{n}^{2}\right)\mathclose{}$ (A.6)
+$y_{i}=\operatorname{sin}(2\pi x_{i})+\epsilon\quad\text{with}\quad\epsilon\sim\mathcal{N}\mathopen{}\left(0,\sigma_{n}^{2}\right)\mathclose{}$ (C.1)
 and draw $x_{i}\sim\mathcal{U}([-1,-0.5]\cup[0.5,1])$.
@@
 All the BNN baselines have the same architecture and fully-factorized Gaussian approximate posterior.
-The prior scale of TFSVI (Rudner et al., 2022b, ) is set to $\sigma_{p}=0.2$ and $\sigma_{p}=0.75$ for MFVI (Blundell et al.,, 2015) and Laplace (Immer et al.,, 2021).
-For the Gaussian process posterior baseline, we fit the prior parameters by maximizing the log-marginal likelihood (Williams and Rasmussen,, 2006).
-Apart from the cases where the parameters of the GP prior used for GFSVI (our method) and FVI (Sun et al.,, 2019) are explicitly stated, we consider a constant zero-mean function and find the parameters of the covariance function by maximizing the log-marginal likelihood from mini-batches (Chen et al.,, 2021).
+The prior scale of TFSVI [Rudner et al., 2022b] is set to $\sigma_{p}=0.2$ and $\sigma_{p}=0.75$ for MFVI [Blundell et al., 2015] and Laplace [Immer et al., 2021].
+For the Gaussian process posterior baseline, we fit the prior parameters by maximizing the log-marginal likelihood [Williams and Rasmussen, 2006].
+Apart from the cases where the parameters of the GP prior used for GFSVI (our method) and FVI [Sun et al., 2019] are explicitly stated, we consider a constant zero-mean function and find the parameters of the covariance function by maximizing the log-marginal likelihood from mini-batches [Chen et al., 2022].
 Except where otherwise stated, we estimate the functional KL divergences with 500 measurement points and use the regularized KL divergence with $\gamma=10^{-10}$.
 Classification
-We sample 100 data points perturbed by Gaussian noise with $\sigma_{n}=0.1$ from the two moons data (Pedregosa et al.,, 2011).
+We sample 100 data points perturbed by Gaussian noise with $\sigma_{n}=0.1$ from the two moons data [Pedregosa et al., 2011].
 On the plots, the data points are shown as red (class 0) and blue (class 1) dots.
@@
 All the BNN baselines have the same architecture and fully-factorized Gaussian approximate posterior.
-The prior scale of MFVI (Blundell et al.,, 2015) is set to $\sigma_{p}=0.8$ and $\sigma_{p}=1.0$ for TFSVI (Rudner et al., 2022b, ) and Laplace (Immer et al.,, 2021).
-For the Gaussian process posterior baseline, we approximate the intractable posterior using the Laplace approximation and find the prior parameters by maximizing the log-marginal likelihood (Williams and Rasmussen,, 2006).
-The GP prior for GFSVI (our method) and FVI (Sun et al.,, 2019) has a constant zero-mean function and we find the parameters of the covariance function by maximizing the log-marginal likelihood from mini-batches (Chen et al.,, 2021) using the method to transform classifications labels into regression targets from Milios et al., (2018).
+The prior scale of MFVI [Blundell et al., 2015] is set to $\sigma_{p}=0.8$ and $\sigma_{p}=1.0$ for TFSVI [Rudner et al., 2022b] and Laplace [Immer et al., 2021].
+For the Gaussian process posterior baseline, we approximate the intractable posterior using the Laplace approximation and find the prior parameters by maximizing the log-marginal likelihood [Williams and Rasmussen, 2006].
+The GP prior for GFSVI (our method) and FVI [Sun et al., 2019] has a constant zero-mean function and we find the parameters of the covariance function by maximizing the log-marginal likelihood from mini-batches [Chen et al., 2022] using the method to transform classifications labels into regression targets from Milios et al.
+[2018].
 We estimate the functional KL divergences with 500 measurement points and use the regularized KL divergence with $\gamma=10^{-10}$.
-A.3.2 Ocean current modeling experiment
-Following Cinquin et al., (2024), we apply the Helmholtz decomposition to the neural network $f$ as
-$f(\cdot,{\bm{w}})=\operatorname{grad}\Phi(\cdot,{\bm{w}}_{1})+\operatorname{rot}\Psi(\cdot,{\bm{w}}_{2})$ (A.7)
+C.2 Ocean current modeling experiment
+Following Cinquin et al. [2024], we apply the Helmholtz decomposition to the neural network $f$ as
+$f(\cdot,{\bm{w}})=\operatorname{grad}\Phi(\cdot,{\bm{w}}_{1})+\operatorname{rot}\Psi(\cdot,{\bm{w}}_{2})$ (C.2)
 where ${\bm{w}}=\left\{{\bm{w}}_{1},{\bm{w}}_{2}\right\}$ and, $\Phi(\cdot,{\bm{w}}_{1})$ and $\Psi(\cdot,{\bm{w}}_{2})$ are 2-layer fully-connected neural networks with $50$ hidden units per layer and hyperbolic tangent activation functions.
@@
 We fit the neural networks on the entire dataset and average the scores with respect to five different random seeds.
-A.3.3 Regression experiments with tabular data
+C.3 Regression experiments with tabular data
 Datasets and pre-processing
-We evaluate the predictive performance of our model on regression datasets from the UCI repository (Dua and Graff,, 2017) described in table 6.
-These datasets are also considered in Sun et al., (2019); Wild et al., (2022) but we include two additional larger ones (Wave and Denmark).
+We evaluate the predictive performance of our model on regression datasets from the UCI repository [Dua and Graff, 2017] described in table 5.
+These datasets are also considered in Sun et al.
+[2019], Wild et al.
+[2022b] but we include two additional larger ones (Wave and Denmark).
 We perform 5-fold cross validation, leave out one fold for testing, consider 10% of the remaining 4 folds as validation data and the rest as training data.
@@
 Baseline specification
-We compare our GFSVI method to two weight-space inference methods (mean-field variational inference (Blundell et al.,, 2015) and linearized Laplace (Immer et al.,, 2021)) and two function-space inference methods (FVI (Sun et al.,, 2019) and TFSVI (Rudner et al., 2022b, )).
+We compare our GFSVI method to two weight-space inference methods (mean-field variational inference [Blundell et al., 2015] and linearized Laplace [Immer et al., 2021]) and two function-space inference methods (FVI [Sun et al., 2019] and TFSVI [Rudner et al., 2022b]).
 While FVI uses GP priors, TFSVI performs inference in function space but with the pushforward to function space of the variational distribution and prior on the weights.
-We compute the function-space (regularized) KL divergence using a set of 500 measurement points sampled from a uniform distribution for GFSVI and TFSVI, and 50 points drawn from a uniform distribution along with 450 samples from the training batch for FVI as specified in Sun et al., (2019).
+We compute the function-space (regularized) KL divergence using a set of 500 measurement points sampled from a uniform distribution for GFSVI and TFSVI, and 50 points drawn from a uniform distribution along with 450 samples from the training batch for FVI as specified in Sun et al.
+[2019].
 All the BNN baselines have the same architecture and fully-factorized Gaussian approximate posterior.
-We also provide results with a GP (Williams and Rasmussen,, 2006) when the size of the dataset allows it, and a sparse GP (Hensman et al.,, 2013).
+We also provide results with a GP [Williams and Rasmussen, 2006] when the size of the dataset allows it, and a sparse GP [Hensman et al., 2013].
 As we restrict our comparison to BNNs, we do not consider the GP and sparse GP as baselines but rather as gold-standards.
-All models have a Gaussian heteroskedastic noise model with a learned scale parameter.
-All the BNNs are fit using the Adam optimizer (Kingma and Ba,, 2017) using a mini-batch size of 2000 samples.
+All models have a Gaussian homoskedastic noise model with a learned scale parameter.
+All the BNNs are fit using the Adam optimizer [Kingma and Ba, 2017] using a mini-batch size of 2000 samples.
 We also perform early stopping when the validation loss stops decreasing.
 Model selection
-Hyper-parameter optimization is conducted using the Bayesian optimization tool provided by Wandb (Biewald,, 2020).
+Hyper-parameter optimization is conducted using the Bayesian optimization tool provided by Wandb [Biewald, 2020].
 BNN parameters are selected to maximize the average validation expected log-likelihood across the 5 cross-validation folds.
 We optimize over prior parameters (kernel and prior scale), learning-rate and activation function.
-We select priors for GFSVI, FVI, sparse GP and GP among the RBF, Matern-1/2, Matern-3/2, Matern-5/2, Linear and Rational Quadratic covariance functions.
-The GP prior parameters used with GFSVI and FVI are selected by maximizing the log-marginal likelihood from batches as proposed by Chen et al., (2021) and done in Sun et al., (2019).
+We select priors for GFSVI, FVI, sparse GP and GP among the RBF, Matérn-1/2, Matérn-3/2, Matérn-5/2, Linear and Rational Quadratic covariance functions.
+The GP prior parameters used with GFSVI and FVI are selected by maximizing the log-marginal likelihood from batches as proposed by Chen et al.
+[2022] and done in Sun et al.
+[2019].
 Hyper-parameters for GPs and sparse GPs (kernel parameters and learning-rate) are selected to maximize the mean log-marginal likelihood of the validation data across the 5 cross-validation folds.
@@
 Number features 13 16 4 9 6 8 8 8 11 49 2
-A.3.4 Classification experiments with image data
+C.4 Classification experiments with image data
 Datasets and pre-processing
-We further evaluate the predictive performance of our model on classification tasks with the MNIST (LeCun et al.,, 2010) and Fashion MNIST (Xiao et al.,, 2017) image data sets.
+We further evaluate the predictive performance of our model on classification tasks with the MNIST [LeCun et al., 2010] and Fashion MNIST [Xiao et al., 2017] image data sets.
 We fit the models on a random subset of 90% of the provided training split, consider the remaining 10% as validation data and evaluate on the provided test split.
@@
 Baseline specification
-We compare our GFSVI method to the same baselines as for the regression experiments (see A.3.3).
+We compare our GFSVI method to the same baselines as for the regression experiments (see C.3).
 All the BNN baselines have the same architecture and fully-factorized Gaussian approximate posterior.
@@
 The random measurement point distribution is sampled from by drawing 50% of the samples from the training data batch and 50% of the samples from a uniform distribution over $[p_{min},p_{max}]^{H\times W\times C}$, where $H$, $W$ and $C$ are respectively the height, width and number of channels of the images, and $p_{min}=v_{min}-0.5\times\Delta$ and $p_{max}=v_{max}+0.5\times\Delta$ where $\Delta=v_{max}-v_{min}$ is the difference between the minimal ($v_{min}$) and maximal ($v_{max}$) pixel values of the data set.
-The random pixel measurement point distribution is taken from Rudner et al., 2022b and is sampled from by randomly choosing each pixel value among the ones available from the training data batch at the same position in the $28\times 28$ pixel grid.
-Finally, the kmnist measurement point distribution is also taken from Rudner et al., 2022b and is drawn from by randomly sampling data points from the Kuzushiji-MNIST (KMNIST) dataset (Clanuwat et al.,, 2018).
+The random pixel measurement point distribution is taken from Rudner et al.
+[2022b] and is sampled from by randomly choosing each pixel value among the ones available from the training data batch at the same position in the $28\times 28$ pixel grid.
+Finally, the kmnist measurement point distribution is also taken from Rudner et al.
+[2022b] and is drawn from by randomly sampling data points from the Kuzushiji-MNIST (KMNIST) dataset [Clanuwat et al., 2018].
 The KMNIST dataset is a collection of 70’000 gray-scale images of size $28\times 28$ which we preprocess by standardizing the images.
-We sample 10 measurement points when using random, 25 measurement points when using random pixel and 20 when using kmnist.
-All the BNNs are trained using the Adam optimizer (Kingma and Ba,, 2017) using a mini-batch size of 100.
+We sample 25 measurement points when using random, 25 measurement points when using random pixel and 20 when using kmnist.
+All the BNNs are trained using the Adam optimizer [Kingma and Ba, 2017] using a mini-batch size of 100.
 We also perform early stopping when the validation loss stops decreasing.
 Model selection
-Hyper-parameter optimization is conducted just like for the regression tasks (see A.3.3).
-The Gaussian process prior parameters used with GFSVI and FVI are selected by maximizing the log-marginal likelihood from batches (Chen et al.,, 2021) using the method to transform classifications labels into regression targets from Milios et al., (2018).
-We optimize the same hyper-parameters as for the regression experiments with the exception of the additional $\alpha_{\epsilon}$ parameter introduced by Milios et al., (2018) for the function-space VI methods with GP priors (FVI and GFSVI).
-A.3.5 OOD detection
+Hyper-parameter optimization is conducted just like for the regression tasks (see C.3).
+The Gaussian process prior parameters used with GFSVI and FVI are selected by maximizing the log-marginal likelihood from batches [Chen et al., 2022] using the method to transform classifications labels into regression targets from Milios et al.
+[2018].
+We optimize the same hyper-parameters as for the regression experiments with the exception of the additional $\alpha_{\epsilon}$ parameter introduced by Milios et al.
+[2018] for the function-space VI methods with GP priors (FVI and GFSVI).
+C.5 OOD detection
 Tabular data with a Gaussian likelihood
-Following the setup from Malinin et al., (2020) we take epistemic uncertainty to be the variance of the mean prediction with respect to samples from the posterior.
-We consider the test data to be in-distribution (ID) data and a subset of the song dataset (Bertin-Mahieux et al.,, 2011) of equal length and with an equal number of features as out-of-distribution (OOD) data.
-We use the same preprocessing as for regression as well as the same baselines with the same hyper-parameters (see section A.3.3).
+Following the setup from Malinin et al.
+[2021] we take epistemic uncertainty to be the variance of the mean prediction with respect to samples from the posterior.
+We consider the test data to be in-distribution (ID) data and a subset of the song dataset [Bertin-Mahieux et al., 2011] of equal length and with an equal number of features as out-of-distribution (OOD) data.
+We use the same preprocessing as for regression as well as the same baselines with the same hyper-parameters (see section C.3).
 We first fit a model, then evaluate the extend by which the epistemic uncertainty under the model is predictive of the ID and OOD data using a single threshold obtained by a depth-1 decision tree fit to minimize the classification loss.
@@
 Image data with a Categorical likelihood
-Following the setup by Osawa et al., (2019), we take the epistemic uncertainty to be the entropy of the mean of the predictive distribution with respect to samples from the posterior.
+Following the setup by Osawa et al.
+[2019], we take the epistemic uncertainty to be the entropy of the mean of the predictive distribution with respect to samples from the posterior.
 We evaluate models trained on MNIST using MNIST’s test split as ID data and a subset of the training set of Fashion MNIST as OOD data.
 Likewise, we evaluate models trained on Fashion MNIST using Fashion MNIST’s test split as ID data and a subset of the training set of MNIST as OOD data.
-We use the same preprocessing as for classification, as well as the same baselines with the same hyper-parameters (see section A.3.4).
+We use the same preprocessing as for classification, as well as the same baselines with the same hyper-parameters (see section C.4).
 We first fit a model, then evaluate the extend by which the epistemic uncertainty under the model is predictive of the ID and OOD data using a single threshold obtained by a depth-1 decision tree fit to minimize the classification loss.
 We estimate mean of the predictive distribution by Monte-Carlo integration using 100 posterior samples.
-We report the mean and standard error of the accuracy of the threshold to classify OOD from ID data based on epistemic uncertainty for the 5 models trained on different random seeds (see section A.3.4).
-A.3.6 Variational measure evaluation
+We report the mean and standard error of the accuracy of the threshold to classify OOD from ID data based on epistemic uncertainty for the 5 models trained on different random seeds (see section C.4).
+C.6 Variational measure evaluation
 We evaluate our inference method by comparing the samples drawn from the exact posterior over functions with the approximate posterior obtained with our method (GFSVI).
-We follow the setup by Wilson et al., (2022) and we compute the average Wasserstein-2 metric between 1000 samples drawn from a GP posterior with a RBF kernel evaluated at the test points, and samples from the approximate posterior of GFSVI, sparse GP and FVI evaluated at the same points and with the same prior.
-We consider the Boston, Concrete, Energy, Wine and Yacht datasets for which the exact GP posterior can be computed and use the same preprocessing as for regression (see section A.3.3).
+We follow the setup by Wilson et al.
+[2022] and we compute the average Wasserstein-2 metric between 1000 samples drawn from a GP posterior with a RBF kernel evaluated at the test points, and samples from the approximate posterior of GFSVI, sparse GP and FVI evaluated at the same points and with the same prior.
+We consider the Boston, Concrete, Energy, Wine and Yacht datasets for which the exact GP posterior can be computed and use the same preprocessing as for regression (see section C.3).
 We report the mean and standard error of the average Wasserstein-2 metric across the 5 folds of cross-validation.
-The Wasserstein-2 metric is computed using the Python Optimal Transport library (Flamary et al.,, 2021).
+The Wasserstein-2 metric is computed using the Python Optimal Transport library [Flamary et al., 2021].
 Baseline specification
@@
 We further consider a sparse GP with 100 inducing points.
-A.3.7 Software
-We use the JAX (Bradbury et al.,, 2018) and DM-Haiku (Hennigan et al.,, 2020) Python libraries to implement our Bayesian neural networks.
+C.7 Software
+We use the JAX [Bradbury et al., 2018] and DM-Haiku [Hennigan et al., 2020] Python libraries to implement our Bayesian neural networks.
 MFVI, linearized Laplace and TFSVI were implemented based on the information in the papers, and code for FVI was adapted to the JAX library from the implementation provided by the authors.
-We further use the GPJAX Python library for experiments involving Gaussian processes (Pinder and Dodd,, 2022).
-A.3.8 Hardware
+We further use the GPJAX Python library for experiments involving Gaussian processes [Pinder and Dodd, 2022].
+C.8 Hardware
 All models were fit using a single NVIDIA RTX 2080Ti GPU with 11GB of memory.
-A.4 Additional experimental results
+Appendix D Additional experimental results
 In this section, we present additional figures for our qualitative uncertainty evaluation as well as further experimental results on regression, out-of-distribution detection and robustness under input distribution shift tasks.
 We also provide plots illustrating the eigenvalue decay of different kernels, and figures showing the influence of $\gamma$ in the regularized KL divergence.
-A.4.1 Qualitative uncertainty evaluation
+D.1 Qualitative uncertainty evaluation
 Figure 4: Our method (GFSVI) effectively regularizes functions generated by the Bayesian neural network (BNN) both in settings where the generative process is very noisy ($\sigma_{n}=1$) or not ($\sigma_{n}=0.1$).
@@
 We further find that our method (GFSVI) provides strong regularization when the data generative process is noisy (see figure 4) and is more robust than FVI to situations where ones computational budget constrains the number of measurement points $M$ to be small (figure 8).
-In contrast to FVI, GFSVI accurately approximates the exact GP posterior under rough (Matern-1/2) GP priors effectively incorporating prior knowledge defined by the GP prior to the inference process (see figure 5).
+In contrast to FVI, GFSVI accurately approximates the exact GP posterior under rough (Matérn-1/2) GP priors effectively incorporating prior knowledge defined by the GP prior to the inference process (see figure 2).
 Likewise, GFSVI adapts to the variability of the functions specified by the kernel (see figure 6).
 We also find that GFSVI requires a larger number of measurement points to capture the behavior of a rougher prior (see figure 7).
-Figure 5: Our method (GFSVI) with a Matern-1/2 Gaussian process (GP) prior accurately approximates the exact GP posterior unlike the function-space prior baseline (FVI).
+Figure 5: Our method (GFSVI) with an RBF Gaussian process (GP) prior accurately approximates the exact GP posterior unlike the function-space prior baseline (FVI).
 Weight-space prior baselines do not provide a straight-forward mechanism to incorporate prior assumptions regarding the functions generated by BNNs and underestimate the epistemic uncertainty (MFVI, Laplace).
@@
 Figure 7: Our method (GFSVI) captures the smooth behavior of a Gaussian process (GP) prior with RBF covariance function even if the number of measurement points is small (M=10).
-However, in that setting GFSVI fails to reproduce the rough effect of a GP prior with a Matern-1/2 covariance function, and requires a larger amount of measurement points to do so (M=100).
+However, in that setting GFSVI fails to reproduce the rough effect of a GP prior with a Matérn-1/2 covariance function, and requires a larger amount of measurement points to do so (M=100).
 Figure 8: Our method (GFSVI) already provides a reasonable approximation to the exact posterior with small numbers of measurement points (M=10) while function-space baseline FVI requires many more (M=100).
 Classification
-We find that GFSVI better captures the beliefs induced by the smooth RBF and rough Matern-1/2 Gaussian process priors compared to FVI (see Figures 9 and 10).
+We find that GFSVI better captures the beliefs induced by the smooth RBF and rough Matérn-1/2 Gaussian process priors compared to FVI (see Figures 9 and 10).
 Moreover, GFVSI both accurately fits the training data and shows greater uncertainty outside of its support relative to BNNs baselines with weight-space and function-space priors.
@@
 Weight-space baselines do not provide a straight-forward mechanism to incorporate prior assumptions regarding the functions generated by BNNs and underestimate the epistemic uncertainty (TFSVI, Laplace) or underfit the data (MFVI).
-Figure 10: Our method (GFSVI) with a Matern-1/2 Gaussian process (GP) prior accurately captures the rough decision boundary unlike the function-space baseline (FVI).
+Figure 10: Our method (GFSVI) with a Matérn-1/2 Gaussian process (GP) prior accurately captures the rough decision boundary unlike the function-space baseline (FVI).
 Weight-space baselines do not provide a straight-forward mechanism to incorporate prior assumptions regarding the functions generated by BNNs and underestimate the epistemic uncertainty (TFSVI, Laplace) or underfit the data (MFVI).
 Inductive biases
-figure 11 compares GFSVI to the exact posterior across two different priors and three model architectures (details in A.3.1).
+figure 11 compares GFSVI to the exact posterior across two different priors and three model architectures (details in C.1).
 We find that the BNN’s ability to incorporate the beliefs introduced by the GP prior depends on its size and activation function.
-When using piece-wise linear activations (ReLU), small models are prone to underfitting for smooth priors (RBF), and to collapsing uncertainty for rough priors (Matern-1/2).
-By contrast, when using smooth activations (Tanh), smaller models suffice, and they are compatible with most standard GP priors (the results shown in figure 11 extend to RBF, Matern family, and Rational Quadratic in our experiments).
+When using piece-wise linear activations (ReLU), small models are prone to underfitting for smooth priors (RBF), and to collapsing uncertainty for rough priors (Matérn-1/2).
+By contrast, when using smooth activations (Tanh), smaller models suffice, and they are compatible with most standard GP priors (the results shown in figure 11 extend to RBF, Matérn family, and Rational Quadratic in our experiments).
 We also analyzed how the number $M$ of measurement points affects performance.
-figures 7 and 14 show that capturing the properties of rough GP priors and estimating $D_{\text{KL}}^{\gamma}$ with these priors requires larger $M$.
+figures 7 and 17 show that capturing the properties of rough GP priors and estimating $D_{\text{KL}}^{\gamma}$ with these priors requires larger $M$.
 Figure 11: Our method requires that the Bayesian neural network (BNN) and Gaussian process (GP) prior share similar inductive biases to provide an accurate approximation to the exact posterior.
-A.4.2 Regression on tabular data
+D.2 Regression on tabular data
 Table 6: Test mean square error (MSE) of evaluated methods on regression datasets.
@@
 Finally, GFSVI compares favorably to the GP and sparse GP.
-A.4.3 Out-of-distribution detection with image data
-We here show an additional plot from our out-of-distribution detection experiment with image data (details in A.3.5).
+D.3 Variational measure evaluation
+table 7 evaluates our inference method by comparing samples drawn from the exact posterior (where computationally feasible) with the approximate posterior obtained with our method (GFSVI).
+We follow the setup by Wilson et al.
+[2022] and we compute the average per-sample Wasserstein-2 metric samples drawn from a GP posterior with RBF kernel evaluated at the test points, and samples from the approximate posterior of GFSVI, sparse GP and FVI evaluated at the same points and with the same prior.
+More details are provided in section C.6.
+We find that GFSVI approximates the exact posterior more accurately that FVI, obtaining a higher mean rank, but worse than the gold standard sparse GP, which demonstrates to be most accurate.
+Table 7: Average point-wise Wasserstein-2 distance (lower is better) between exact and approximate posterior of reported methods. GFSVI (ours) provides a more accurate approximation than FVI.
+Dataset Boston Concrete Energy Wine Yacht Mean rank
+GFSVI (ours) 0.0259 $\pm$ 0.0040 0.0499 $\pm$ 0.0029 0.0035 $\pm$ 0.0004 0.0571 $\pm$ 0.0097 0.0036 $\pm$ 0.0006 1.0
+FVI 0.0469 $\pm$ 0.0044 0.0652 $\pm$ 0.0037 0.0037 $\pm$ 0.0004 0.1224 $\pm$ 0.0167 0.0052 $\pm$ 0.0013 1.6
+GP sparse 0.0074 $\pm$ 0.0022 0.0125 $\pm$ 0.0016 0.0042 $\pm$ 0.0003 0.0170 $\pm$ 0.0035 0.0035 $\pm$ 0.0008 -
+D.4 Out-of-distribution detection with image data
+We here show an additional plot from our out-of-distribution detection experiment with image data (details in C.5).
 figure 12 shows the (normalized) histograms of the entropy of the mean prediction produced by each model on the in-distribution (blue) and out-of-distribution (red) data sets considered in our OOD detection experiment.
@@
 GFSVI (ours) best partitions in-distribution and out-of-distribution data based on the entropy of its mean predictive distribution.
-A.4.4 Influence of measurement point distribution for image data.
+D.5 Influence of measurement point distribution for image data
 We present additional results evaluating the influence of the measurement point distribution $\rho$ on the the performance of function-space inference methods when using high-dimensional image data.
-The measurement point distribution are described in section A.3.4.
-Just like in Rudner et al., 2022b , we find that the choice of measurement point distribution may highly influence the OOD detection accuracy.
+The measurement point distribution are described in section C.4.
+Just like in Rudner et al.
+[2022b], we find that the choice of measurement point distribution may highly influence the OOD detection accuracy.
 While the expected log-likelihood, accuracy and expected calibration error (ECE) of a model generally remains comparable across measurement point distributions, the OOD accuracy of GFSVI is greatly improved by using samples from kmnist to evaluate the (regularized) KL divergence.
@@
 OOD acc. ($\uparrow$) -0.853 $\pm$ 0.005 -0.867 $\pm$ 0.005 -0.997 $\pm$ 0.001 -0.925 $\pm$ 0.005 -0.842 $\pm$ 0.006 -0.975 $\pm$ 0.002 -0.802 $\pm$ 0.006 -0.800 $\pm$ 0.007 -0.779 $\pm$ 0.010
-A.4.5 Input distribution shift with rotated image data
+D.6 Input distribution shift with rotated image data
 We here provide an experiment evaluating our method’s (GFSVI) robustness in detecting input distribution shift.
 We expect the predictive uncertainty of a well-calibrated Bayesian model to be low for in-distribution data and to gradually increase as the input distribution shifts further away from the training data distribution.
-To test this property, we follow the setup by Sensoy et al., (2018); Rudner et al., 2022b and assume like the related work that increasing the rotation angle of images gradually increases the level of input "distribution shift".
-We report the mean and standard-deviation of the average mean predictive entropy of models fit on MNIST (LeCun et al.,, 2010) and Fashion MNIST (Xiao et al.,, 2017) for increasingly large angles of rotation of their respective test data partition.
+To test this property, we follow the setup by Sensoy et al.
+[2018], Rudner et al.
+[2022b] and assume like the related work that increasing the rotation angle of images gradually increases the level of input "distribution shift".
+We report the mean and standard-deviation of the average mean predictive entropy of models fit on MNIST [LeCun et al., 2010] and Fashion MNIST [Xiao et al., 2017] for increasingly large angles of rotation of their respective test data partition.
 We find that GFSVI is confident (low predictive entropy) for images with small rotation angles, and that its predictive entropy increases with the angle.
@@
 We see that our method (GFSVI) exhibits the behavior of a well-calibrated Bayesian model.
-A.4.6 Additional plots for kernel eigenvalue decay
-figure 14 shows a plot demonstrating the decay rate of the eigenvalues of RBF and Matern-1/2 kernels evaluated at points sampled uniformly over $\mathcal{X}$ The rate of decay of covariance operator’s eigenvalues gives important information about the smoothness of stationary kernels (Williams and Rasmussen,, 2006) and that increased smoothness of the kernel leads to faster decay of eigenvalues Santin and Schaback, (2016).
-For instance, RBF covariance operator eigenvalues decay at near exponential rate independent of the underlying measure (Belkin,, 2018) and Matern kernels eigenvalues decay polynomialy (Chen et al.,, 2021).
+D.7 Example of model misspecification with Wild et al. [2022b]
+Figure 14: Example of model misspecification when using a periodic GP prior with baseline GWI [Wild et al., 2022b] that does not occur with our method (GFSVI).
+In GWI, only the posterior covariance is periodic, while the neural network parameterizing the posterior mean results in a function that does not capture the beliefs carried by the (periodic) prior.
+In contrast, our method accurately captures the GP prior’s beliefs and yields a (locally) periodic function.
+figure 14 shows an example of model misspecification when using a periodic GP prior with the baseline GWI [Wild et al., 2022b].
+As can be seen in the left panel of the figure, this problem does not occur with our method (GFSVI).
+While the posterior covariance in GWI reflects the periodicity of the prior, the neural network parametrizing the posterior mean does not result in a periodic function, i.e., the mean does not capture the beliefs specified by the periodic GP prior.
+In contrast, our method accurately captures the GP prior’s beliefs and yields a (locally) periodic function.
+D.8 Convergence speed on UCI data
+Table 9: Training time of our method GFSVI and baselines MFVI [Blundell et al., 2015] and TFSVI [Rudner et al., 2022b] on the boston UCI dataset.
+GFSVI (ours) TFSVI MFVI
+Time (s) 44.15 $\pm$ 1.56 36.36 $\pm$ 0.90 38.38 $\pm$ 10.80
+table 9 shows the training time of our method and baselines MFVI [Blundell et al., 2015] and TFSVI [Rudner et al., 2022b] on the boston dataset using $M=100$ context points averaged over $5$ cross-validation splits, as well as figure 15 showing the convergence of the validation expected log-likelihood on the boston dataset.
+Our method converges in more steps than the TFSVI.
+GFSVI typically takes more time/steps to train that TFSVI as it additionally needs to adapt its features to the beliefs specified by the Gaussian process prior.
+Figure 15: Validation expected log-likelihood of our method (GFSVI) and baselines TFSVI and MFVI. GFSVI (ours) converges on the boston dataset in slightly more steps than TFSVI but in fewer than MFVI.
+Figure 16: Validation expected log-likelihood of our method (GFSVI) and baselines TFSVI and MFVI on MNIST. GFSVI (ours) converges in slightly more steps than TFSVI but in fewer than MFVI.
+D.9 Additional plots for kernel eigenvalue decay
+figure 17 shows a plot demonstrating the decay rate of the eigenvalues of RBF and Matérn-1/2 kernels evaluated at points sampled uniformly over $\mathcal{X}$ The rate of decay of covariance operator’s eigenvalues gives important information about the smoothness of stationary kernels [Williams and Rasmussen, 2006] and that increased smoothness of the kernel leads to faster decay of eigenvalues Santin and Schaback [2016].
+For instance, RBF covariance operator eigenvalues decay at near exponential rate independent of the underlying measure [Belkin, 2018] and Matérn kernels eigenvalues decay polynomialy [Chen et al., 2022].
 We find that the kernel evaluated at points sampled from a uniform distribution over $\mathcal{X}$ share this same behavior (see figure 17).
@@
 The mean eigenvalues are arranged in increasing order.
-The eigenvalues of the Gram matrix associated with the smooth RBF kernel decays much faster than those of the Matern-1/2.
+The eigenvalues of the Gram matrix associated with the smooth RBF kernel decays much faster than those of the Matérn-1/2.
 Furthermore, the eigenvalues decay at a slower rate in high dimensions (D=100).
-A.4.7 Additional plots for choosing $\gamma$ in $D_{\text{KL}}^{\gamma}$
+D.10 Additional plots for choosing $\gamma$ in $D_{\text{KL}}^{\gamma}$
 The $\gamma$ parameter controls the magnitude of the regularized KL divergence (see figure 20) and adjusts the relative weight of the regularized KL divergence and expected log-likelihood term in the training objective (see figure 18).
@@
 Figure 18: The $\gamma$ parameter of the regularized KL divergence controls the magnitude of the regularizer in the objective and should be small enough to provide strong regularization.
-Figure 16: The BNN’s covariance adaptation to the prior’s covariance rank depends on its activation function. BNNs fit with a RBF prior (full) show lower rank than with a Matern-1/2 (dotted).
+Figure 19: The BNN’s covariance adaptation to the prior’s covariance rank depends on its activation function. BNNs fit with a RBF prior (full) show lower rank than with a Matérn-1/2 (dotted).
 Figure 20: $\gamma$ explicitly controls the magnitude of the regularized KL-divergence $D_{\text{KL}}^{\gamma}$.
-Rougher priors (Matern-1/2) require more measurement points to accurately estimate $D_{\text{KL}}^{\gamma}$ than smooth priors (RBF).
+Rougher priors (Matérn-1/2) require more measurement points to accurately estimate $D_{\text{KL}}^{\gamma}$ than smooth priors (RBF).